A Complete Bibliography of Publications of John von Neumann

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 1.220

Abstract This bibliography records publications of John von Neumann (1903–

1957).

Title word cross-reference

1 + 2 [vN51c]. $125 [Lup03]. $19.95 [Kev81]. 2,000 [MRvN50]. $23.00 [MC00]. $25.00 [Jon04]. $29.95 [CK12]. 2^{1/3} [vNT55]. $35.00 [Ano91, Pan92]. $37.50 [Ano91]. $45.00 [Ano91]. e [MRvN50, Rei50]. F_{∞} [vN62]. H [vN29b, von10]. N [vN46b]. ν [vN62]. p [DJ13]. p = Kρ^{4/3} [vNxxb]. π [MRvN50, Rei50, She12]. q [DJ13]. T_{ρ}^{ρ} = 0 [vN35h].

-theorem [von10]. -Theorems [vN29b].

/119.00 [Emc02]. /89.00 [Emc02].

0 [Lup03, MC00, Rec07]. 0-19-286162-X [Twe93]. 0-201-50814-1 [Ano91]. 0-262-01121-2 [Ano91, Bow92]. 0-262-12146-8 [Ano91]. 0-262-16123-0 [Ano91]. 0-521-52094-0 [Jon04]. 0-7923-6812-6 [Emc02, Lup03]. 0-8027-1348-3 [MC00]. 0-8218-3776-1 [Rec07].

1 [GvN47b, GvN47c, GvN47d, Rec07, vNW28a]. 10/18/51 [McC83]. 12th [Var88, Wah96]. 1900s [IM95]. 1927 [Has10a]. 1940s [IM09a]. 1942-1952 [Fit13]. 1946 [vN57b]. 1949 [Ano51]. 1950s [IM09a, Mah11]. 1954 [Réd99]. 1957 [Ano57c, OPP58, Tel57, Wig57, Wig67a, Wig67c, Wig80]. 1981 [Bir83]. 1982 [TWA^{+}87]. 1988 [B^{+}89, GIS90, Var88]. 1989 [B^{+}89, RMK94]. 1989/90 [B^{+}89]. 1990 [CW91]. 1992 [Bir97]. 1997 [Ano97]. 1999 [E^{+}00, Fro00]. 1R [Har67].

2 [GvN47e, GvN47f, GvN47g, vNW28b]. 20th [CW91, Cin85, Cip00, Kőr99]. 20th-Century [Cin85]. 25th [Ano97, Tót93]. 28 [B^{+}89, Tel57]. 28.September [B^{+}89].

3 [MC00, vNW28c]. 31.50 [Bow92]. 37 [Has10b].

4 [B^{+}89, Har67]. 4/67 [Har67]. 4/67-1R [Har67]. 423pp [CK12].

6 [Lup03]. 6.25 [Ran74]. 67-1R [Har67].

80th [Sze83b].

’94 [BR94].

= [Har89a, Har89b].

A. [Fie15]. A.E.C. [Ano57a]. Abbildungen [vN32a]. Abelian [vN40b]. Aberdeen [Ber00]. abstract [Mat03]. Abstracting [Lor19]. abzählbar [vN28c]. Academic [Emc02, Lup03, B^{+}89]. Accomplish [Mir92]. accomplishments [Bre04]. Accounting [Tod74]. accumulation [AFB79]. Accuracy [Neu33]. ACM [Ano91, Ano87]. acting [CvN44a, CvN44b]. Activity [IM09a]. Adaptive [Hol62]. adatok [Nag87b]. Addison [Ano91]. Addison-Wesley [Ano91]. Addition [vN29i]. Additions [vN32i]. address [Réd99]. Adiabatic [vNW29a]. adiabatischen [vNW29a]. Adjoint [vN32e, vN32f, DNvN55]. adjungierte [vN32e, vN32f]. Advanced [Asp87, Mit80, Reg87]. adventurous [ABES12]. Aerodynamical [vN10, vN51d]. aerodynamics [Ano51]. African [CFK^{+}91]. after [Ano93]. age [Mar96, CW91]. Agencies [Ano57a]. ago [Ano95, Bal04, Csá03, Kov03b, Szé78, Sze97, Zád83]. agy [Neu64, Neu72, Neu06]. agyról [Érd85]. agyú [Sze76]. AI [Bou18]. aiming [vN41a]. akademische [B^{+}89]. Akik [Kőr99]. al [CFK^{+}91]. al-Khwarizmi [CFK^{+}91]. Alakja [Bal04]. Alamos [Fit99, Fit13, Lax03, HPR14a, Lax03, Lax05]. Alamos-i [Lax03]. Alan [Lie11, Ano96, Asp80, IM13, Lie11, Müh09, Ran72]. alapelvének [Kov03b]. Algebra [Bac78, Bac87, vN29f, ABS96, Réd59, vNS35]. Algebraic [vN37a, JvNW34, vN36d, vN31a]. algebraical [vN43b]. Algebraically [vN28e]. algebraisch [vN28e]. Algebraische [vN31a]. algebras [Ara90, DK04, Kad58]. Algorithm [Fer99, JML13, DG17, Dya71, LLY22, iP13]. Algorithms [Cip00, Knu98, FR96]. alkalmazó [Sze03a]. alla [Ven02]. Allgemeine [Tur60a, vN51e, vN28a, vN29a]. allgemeinen [vN28h, vN29g, vN29i]. Almost [BvN35a, BvN35b, Tau61b, vN34a, vN34b, vNW40]. általános [Von90b, vN26a]. am [B^{+}89]. America [Jon04, FB69, Tar77]. American [Ame52, Rec07, Bir77, Nor98, Tar77]. Amiss [HSS09]. Amplification [vN53c]. Amplifying [vN59a, vN54b]. AMS [DK04]. Amsterdam [Réd99]. analizu [vN87]. analógia [Érd85]. Analogie [Cso04]. analogue [Dya71]. analogy [Cso04, Érd85]. Analysis [Grc11, Grc12, JML13, Tau63b, Wil71, Asp89c, LY13, LLY22, Mad12, RKH^{+}21, vN44b, vN87]. Analytic [vN29c, vN33, KvN37]. analytischen [vN29c]. analytischer [vN33]. Analyzing [HKLS93]. Ananyo [Kum21]. Andrew [WS92, WS03]. Anecdotes [TWA^{+}87]. Angeles [HFG51]. Anhang [Tur60a]. Animal [vN49a]. Animate [Hon09]. anläßich [B^{+}89]. anläßlich [B^{+}89]. Anniversary [CFK^{+}91, Sze83b, Tót93]. annual [Ano88, JC87, Tar77, Wei92]. Answering [HKLS93]. Appendix [Asp83, vN41a, Tur60a]. Apple [CW91]. Application [Kat60, TWA^{+}87]. Applications [BR94, vN32c, vN52a]. Applied [Ali09, Fro00, Kah97, Día05, Stö04]. Appraisal [Bub10, Asp83, Ste81]. Approach [Mar46, Pos95a, LY13]. Approximative [vN42a]. aranycsinálója [Sze83c]. Arbeit [vN29i, vN31b, vN32i]. Architect [Ano57a, Kov03b]. architecture [Cer00, Roj94]. architectures [RH00]. arising [CvN42, CvN43, CvN44a, CvN44b]. arithmetics [vN37c]. Army [Ber00]. Arthur [All87, Har67]. Article [vN93a]. Artificial [Cso04, Wah96, Mar98]. Ascent [Bro73a, CW91]. Asimov [CFK^{+}91]. Aspect [Pos95b]. Aspects [GvN47a]. Aspray [All87, Ano91, Bow92, Ham92, CFK^{+}91, CW91, Kid92, Mah93, Pan92]. assignment [vN53a]. Association [Ano91, Jon04, Tar77]. assuming [vNxxb]. assured [Byr10]. astronomy [FF91]. Astrophysics [Tau63c]. Asymptotic [LLY22]. Atanasoff [Ano96, Smi10]. Atomic [vN46a, vN57b, vN55b, vN55c]. áttekintés [Döm03]. áttekintése [SI^{+}79]. atyja [Kov84]. Aufbau [vN27c]. August [Ano51, BR94, Fro00, Var88, Wah96]. Ausbildung [B^{+}89]. Ausführungen [vN31b]. Ausnutzung [vN54a]. Australian [CFK^{+}91]. authorship [Fie15]. Automata [Bur70, SM56, Tau63b, Tur60a, vN41c, vNB66, vN96d, Sha58, Tar72, Vol06, vN51g, vN63b, vN71b, Asp90c, CFK^{+}91, Har67]. Automaten [Tur60a]. automates [vN96d]. Automatic [vN47a]. Automaton [Bur61, Hon09]. avenue [Ném93]. Averages [HKLS93]. Averaging [HKLS93]. avtomatov [Tur60a, vN71b]. award [B^{+}89, Ano57a, Ano87, Fie15]. Away [Lor19]. axiom [Ben72]. Axiomatic [vN29d, Haj59, Mit22]. axiomatikus [vN26a]. axiomatischen [vN29d]. Axiomatisierung [vN25b, vN28b]. Axiomatization [vN28b, vN25b, von97]. Axiomatizations [Lac00]. Azarya [GC97].

Babbage [Ano91, CW91]. Back [Asp90a, CW91]. Baleful [Kah97]. Baltimore [DK04]. barotropic [CFvN50]. Barry [CW91]. Basics [Asp90a, CW91]. Basis [vN27a, vN31d]. Bateson [Hei77]. Be [Bac78, Bac87]. Beauregard [Ber69]. Beautiful [Mor02, Nas98a]. became [Ném93]. Become [Kov84]. Bedford [Asp83]. Before [CFK^{+}91, vN46a, vN57b]. Beginning [CLBdB89]. Beginnings [Sma83]. Begründung [vN27a]. Behavior [Ano57a, vNW29a, vNM44, vNM47, vNM53b, vNM61, vNM64, vNM67a, vNM67b, vNM73, vNM74, vNM04, Jef51, vNM53a, vNM70, vNM80, vNM07, Ebe67, Haw45, Hur48, Rel92]. Behaviour [K.44]. Believe [Bec04]. Believes [Gra56]. Bell [Sch91, Ali09, Pos95b, Whi16]. Bells [Sch91]. Bemerkungen [vN28f, vN31b]. Benchmarks [Kah97]. Berlin [Has10b, Has10a, Ano30, Has10a]. Best [Cip00, vN48b]. Bets [HPR14a]. between [DD13, GM04, Rel92]. bev [Kor86]. Beweis [vN29b]. Beweise [vN29e]. Beweistheorie [vN27d]. Beweistheorie“ [vN31b]. Beyond [IM09a]. Bhattacharya [Kum21]. bibliography [KW64, Asp83]. Bicentenary [CFK^{+}91]. bicentennial [Tar77]. Biographies [Ula82b, Wei88]. Biography [Von88, Wig80, Wig96a, CFK^{+}91, Smi10]. biológiája [Ven03]. Biological [Ano97]. biology [HD13, Ven03]. Birkhoff [Ven02, Zun02]. birth [Ano93, Sze83b]. Birthday [Grc05]. bis [vN31a]. bit [DG17]. Bits [Per92]. Blast [BFvN^{+}44, BFH^{+}47, GvN55, PPvNF44]. body [KvN97]. Bohm [Pin77, Pos95b, Stö99]. Bomb [Ano57a, Lan93a, Twe93, Gol84, Pou93, Pou03]. Bombe [CFK^{+}91, Pou03]. Bombs [DK41, Wal16]. Book [All87, Asp83, Ber69, Bir83, Bir97, Bow92, Cor11, Dic03, Dys81, Ebe67, Emc02, God74, Har67, Haw45, HB82, Hod81, Jon04, K.44, Kev81, Lup03, Lus19, MC00, Mah93, Pan92, Ran74, Rec07, Sto83, Tau60, Tro76, Twe93, vN49a, vH67]. Books [CK12, Hal85]. Boolean [vNS35]. Boostered [vN43c, RKH^{+}21]. Borel [Del95, vN53b]. Born [Kov94, Kov03b, Szé78, Zád83, Bac22]. Boston [Emc02]. bottleneck [PJW^{+}22]. boundaries [Fie15]. Boundary [FvN51]. Brain [Ano57a, BH89, Fre13, Neu64, Neu72, Neu06, URR86b, vN58a, vN60b, vN65, vN93a, vNCC00, vN12, Bro73b, Érd85, PA13, PJW^{+}22, Ula58b, vN70, vN80, vN86, vN91a, vN91b, PA13, Tau60]. Brains [IM13, Wal16]. breaking [Sip05, Sze76]. brief [CFK^{+}91]. brilliant [Kov03a]. brother [Neu03, Von87, Von91]. brothers [Fil03]. Brouwer [vN37f]. Brouwerschen [vN37f]. Budapest [RS01b, Var88, Wah96]. Building [Emc97]. Built [Ano57a]. bulk [vN96c]. Burks [Har67, All87]. Burning [CEE^{+}51]. burst [vNR57]. Business [CFK^{+}91, CW91]. Byte [CFK^{+}91].

C [CW91, Sel92]. C. [Bre08, vN52a]. cache [PJW^{+}22]. calculated [MRvN50]. Calculating [Ake07, CFK^{+}91]. Calculation [CW91, GvN55, vN48a, vN48d, vNR50]. calculator [Bau98a, Érd85]. Calculus [vN30]. California [HFG51]. Cambridge [Ano91, ABS96, Bir83, Bow92, CFK^{+}91, Jon04, Kev81]. came [Bal04, Csá03]. Campbell [CFK^{+}91]. Campbell-Kelly [CFK^{+}91]. Can [Bac78, Bac87, Gra56, vN55a, Kow15, Tur60a, TvN99]. cannot [Neu89b]. Capacitance [vN59a, vN54b]. Capital [Wal16]. Card [CW91]. Career [Lan89]. Carlo [HFG51, Ben16, Dys12a, Eck87, HPR14a, Hur85, JML13, KW64]. Case [Asp89a, Dal01, NAH^{+}89, Ven02]. Case-Study [NAH^{+}89]. Caso [Ven02]. Catastrophe [Leo06, Kow15]. Cathedral [CK12, Dys12c]. causality [GL01]. Celebration [Ber00, DK04]. Cells [vN61b]. Cellular [Bur61, Bur70, Vol06]. Centenario [Pel03]. Centenary [Pel03]. Centennial [Grc05, DK04]. Center [Ano88, Joh86]. Centre [E^{+}00]. Century [CW91, Cor11, Bal12, Brü87, Har06b, Her88, Her95, Whi16, Cin85, Cip00, IM01, IM09b, Kőr99, MHR80]. Cerebral [Jef51]. Ceres [Kow15]. Certain [vN39, FvN22, vN36c, vN53a, vN58b]. Chain [JML13]. challenges [Fro00]. Change [CLBdB89]. changed [Har06b, Mar06]. changing [Nat88]. Character [LvN47]. Characteristics [vN29c]. Characterization [Bon04, vN61a, vN35b]. Charakterisierung [vN35b]. Charles [Ano91, CW91]. Charts [Kev81]. Checking [McC83]. chemical [vN55c]. Chemistry [Kah97]. Chess [Leo06, Leo10]. Chips [CLBdB89]. chromodynamics [Fro00]. científico [IM01]. Circle [Lup03]. Circuit [CW91]. Circulation [Sma83]. Clarifying [DD13]. classes [vNS35]. Classic [Lew21, vN93a]. Classical [vN32h, vN32i, vNH41, HvN42, Jam74b]. Claude [TWA^{+}87]. Co [Fie15, MC00, Ano91, Göt03]. Co-authorship [Fie15]. co-operation [Göt03]. co-published [Ano91]. Co. [CFK^{+}91]. Code [HPR14b, PJW^{+}22]. Coding [TWA^{+}87, GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, vNG48a, vNG48b]. Cold [Ake07, Bre04]. COLING [Var88]. Collaboration [Ber10, MJ09, Rel92]. Collapse [Bec04]. Collapsed [vN35f]. Collected [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Wig96c]. Collection [MHR80]. College [B^{+}89, RMK94]. collision [vNS43]. collisions [WvN54]. colloquium [B^{+}89]. Combination [vN47f]. combinations [UvN47b]. come [IM95, IM08]. commemorating [Tar77]. Commemoration [BH89]. Commentary [Gol84]. Comments [Bre08, SW80, vN31b]. Committee [vN46a, vN57b]. Communicate [Pin77]. Communication [vN49a, vN53b]. community [Ins12, Mit80]. commutative [DK04]. como [IM01]. compact [BvN35c]. Comparing [Khr08]. Comparison [Sch91, vN61b, For72a, For72b, Lie11]. compendium [vNBV95]. complementary [KvN97]. Complemented [vNH37]. complete [vN35c, vN35d]. completion [Mit22]. complex [Veb37, VG55]. complexity [DG17]. Components [vN56b, vN63c]. Computation [Ste74]. Computational [Müh09, MJ09, Var88]. Computations [Hur85]. Computer [All87, AB87, Asp89b, BN71, BH89, BR94, CW91, Dan85, Fre13, Fro00, God74, GH93, Gol72, Gol83b, Gra63, Gri03, Kah97, Knu70, Kov89, Kov94, Lap96, Lew21, LR67, Mac92a, Mac92b, Mac94, McC01, Neu64, Neu72, Neu06, Pel99, Ran74, Sla87, SW80, Swa76, Tau60, Tro75, Tro76, vN58a, vN60b, vN65, vN70, vN80, vN86, vN91a, vN91b, vNCC00, vN12, Ano96, Ano97, Asp80, Asp87, Asp89c, Ben16, Bre04, Bro73b, CK12, Cer00, Érd85, Fie15, Gol87b, Gol93, Gol03, Gol87c, HP20, Kov03b, Lie11, Mac99b, Nor98, Rév83, Roj94, Shu84a, Shu85a, Shu96a, Smi10, Sze83a, Sze03a, Sze83b, Sze76, CW91, MC00]. computer-minded [Sze76]. Computer-Research [Mac94]. Computerforschung [Mac94]. Computerforschung-Facetten [Mac94]. Computers [Ano91, CLBdB89, E^{+}00, IM13, Kov84, Lie11, Ran82, Tau63b, URR86b, vN47a, Ake07, PJW^{+}22, Ran72, RH00, Ste81, Ula58b, URR86c, CFK^{+}91, Asp83]. Computing [All87, Ano91, AB87, Asp89b, Asp90b, Ber00, Bow92, BGvN82, CFK^{+}91, CDJ84, E^{+}00, Fro00, GvN47a, GvN63, Lax05, Mah11, MHR80, MCP23, Ste80, Tót93, Ula80, URR86d, vNG46, vN47b, vN54c, vN81a, vN81b, vN88, Ham92, BGvN46, Gal93, GvN45, GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, Lax03, MH11, Tar58, vNG48a, vNG48b, vN51f, Ano91, CFK^{+}91, CW91, Pan92, Kid92, Mah93]. Concavity [GM04]. concentrate [Ano30]. Concept [Asp90a, HPR14c, Réd01, CW91]. Conception [Hon09, Kje01]. Conceptions [Pel99]. concerning [Tar58, vN42b, vN52a]. Condition [Kat60, FvN22]. Conditions [Pin77, KvNS63]. conducted [Veb37, VG55]. Conduction [CN47]. Conference [McC83, Var88, Wah96]. Configurations [vN35f]. Conformal [vN51c]. Congress [Ame52, BR94, Réd99]. Congressional [vN57b]. congruent [vN28c]. Conjecture [HKLS93, Sam92, vNG53]. Connection [For51, NHN89, vN51b, vN63d]. Connections [Lan93b, Lan99]. considerations [Khr08]. Consistency [vN29d]. constant [KvNS63]. Construction [Fra07]. constructivity [Asp80]. Context [Cin85, RKH^{+}21]. Contexts [Kje01]. Continuation [Lac00]. Continued [vNT55]. Continuous [vN36a, vN37b, vN37c, vN52b, vN60a, vNH81, vN98, Hal85, KvN32, vN27e, vN36b, vN37a, vN58b, Tau62]. Contradiction [vN29d]. Contribution [Kap95, RW12]. Contributions [Ano57a, Asp89b, Sha58, vH58]. Control [vN49a]. Convergence [HKLS93, JML13]. conversation [Ano95]. Coordinate [Neu33]. coordinatization [FH55]. correlations [CvN43, CvN44a]. Correspond [Chu21]. cosmical [Ano51]. Costa [Ber69]. Coulomb [KvN97]. countable [vN28c]. covering [KW64]. Craig [Ano91]. Created [Isa14, Lew21]. Creation [Wal16, Leo10]. Creativity [ADS83, Asp89a, NAH^{+}89]. creators [Str11]. Critiche [Ven02]. Criticism [Ven02]. Cross [SvN46a, SvN48a, SvN46b, SvN48b]. Cross-Space [SvN46a, SvN48a, SvN46b, SvN48b]. Crossley [CFK^{+}91]. crossroads [ET07]. csodálatos [Mor02]. cuántica [vN49b, vN91c]. Cybernetics [Kay97, vN49a].

d [Cso04, CEE^{+}51]. Darstellungen [vN27e, vN29c]. data [Nag87b]. daughter [vNW12]. David [Zun02, Pin77]. Day [Ano97, DW16, HHW99]. Days [Gol87a]. Death [CFK^{+}91, Kac82, Ano97, Hei80, Sto83, Dys81, Hod81, Kev81, Ste82, HB82]. Deavours [CFK^{+}91]. Debate [SW80, Roj94]. Debugging [FS12]. Dec [Wig80]. December [Tel57]. Decimal [Rei50, MRvN50]. decisions [Str63a]. Defense [vN55b]. defined [KvN37]. Definition [vN28h]. Degeneracy [vN35f, vNxxb]. degenerate [Khr08]. degree [ABS11]. del [IM95, IM01]. della [IG02, Lup18, Lus19]. Demons [Dys12d]. denken [Tur60a]. Dense [vN25a]. Department [Gol87a]. departments [Ano30]. Description [vN51c]. Design [BGvN82, BGvN46, HP20, Swa76, vNF46, Tau63b]. Desire [Kri95]. Desk [WE14]. destruction [Byr10]. Determination [Rei50, She12, vN48b, vNS35]. determine [vN54d]. Deterministic [vN47f, UvN47b]. Deterrence [Mac92a, Mac92b, Mac99b]. detonation [vN42c, vN44a, vN48c]. Detonations [vN43c, RKH^{+}21]. Detroit [Ame52]. Deutschen [vN49c]. Deutschland [Has10b, Has10a]. Development [Ano57a, Cin85, Dan85, GM04, Mad12, Swa76, vN54a]. developments [vN56a]. devices [vN54b]. Devouring [CW91]. Diagrams [vN32a]. Diákkori [Ném84]. dichte [vN25a]. Did [Bec04, Kov84, Réd96]. Didn’t [Fie14, HP20]. died [Sze97]. dies [Ano57a]. Difference [Ano91, CW91, Har42, LY13, vN41b, vNKBH41, vN42b]. differences [vNK40]. Different [Kje01]. Differential [CN47, Fri80, vN51c, ABS96, BvN35c, BvN50, vNR47a, vN59b]. Diffusion [CW91, RvN85, RUvN47, vNR47b]. Digital [Asp83, CK12, Gra63, Isa14, LR67, Ran82, SW80, Ste80, Tós04, Dys12c, Ran72, Smi10]. Digits [For51, MRvN50, vN51b, vN63d]. Dilemma [Twe93, Pou93, Pou03]. dilemme [Pou03]. Dimensional [BMvN^{+}59, vN44c]. Dimensions [Ano51]. Dirac [Neu33, Mar65, Neu33, TVvN34, vN28f, vN40a]. Dirac-félé [Neu33]. Diracschen [Neu33, vN28f]. direct [vN38]. disciplinary [Fie15]. discontinuities [vN44c]. discoverers [Nat88, CLBdB89]. Discovery [Cin85, Rob12]. Discrete [B^{+}89, vNW29b]. Discussion [Asp89a, BGvN82, NAH^{+}89, vN40a, vN51d, vN52a, vN63a, vN74, vN10, BGvN46]. Disequilibrium [vN53c]. Diskrete [B^{+}89, vNW29b]. distance [CvN44a]. Distributed [CDJ84]. Distribution [Coo57, vN41b, CvN42, CvN43, CvN44a, CvN44b, Dya71, vN42b]. distributions [FR96, For72a, For72b]. Divergences [Del95]. Divisions [Pin77]. DNA [Bre84a]. Do [Pin77]. Documentary [Ano19, NPP66]. documents [Nag87a]. Does [Chu21, LR67, Pin77]. Dokumentumok [Nag87a]. domains [Veb37, VG55]. domestic [Frö93, Lad93]. Dominio [IG02]. Dominion [IG02]. Donald [CFK^{+}91]. Dordrecht [Emc02, Lup03].

Dr. [Ano57a]. Draft [Asp83, vN45a, vN82, vN93b, vN21]. Drafting [SW80]. Drehelektrons [vNW28a, vNW28b, vNW28c, vN28f]. durch [vN28h]. during [Ame52, Ake07]. Dynamical [KvN32, CvN43, RW12]. Dynamics [E^{+}00]. Dyson [CK12]. Dzhon [Dan90].

Early [Ano91, CW91, Gol87a, Hur85, Lac00, Lax05, Sma83, Lax03]. ECAI [Wah96]. eccentricity [Reg87]. Eckert [Asp83, Ano96, Ste81]. Economic [Ano57a, Ebe67, Haw45, Hur48, K.44, vNM44, vNM47, vNM53b, vNM61, vNM64, vNM67a, vNM67b, vNM73, vNM74, vNM04, Rel92, vN45b, vNM53a, vNM70, vNM80, vNM07]. Economics [GM04, Laf89, Mar46, Sam92, vN37f, DCG89, KT58, Zal03, vN56a]. economy [Wei92, vN56a]. Ed [Rec07, Lus19]. Eden [CW91]. Edge [Gal93, Hol18]. Edited [Emc02]. edition [Twe93]. Editor [Har67, Ano91, WE14]. Editors [Cip00]. Eds [Lup03, Dic03]. EDVAC [Asp83, SW80, vN45a, vN82, vN93b, vN21]. Edvard [Ano91, CW91]. Edward [Hor97]. Effect [Kah97, vNR57]. Efficient [PJW^{+}22]. Egyenletesen [vN25a]. egyik [Cze08]. Eigenschaften [vNW28a, vNW28b, vNW28c, vN29c]. Eigenvalue [vN28a, vNW28a, vNW28b, vNW28c, vN29a, vN28d]. Eigenvalues [vNW29b, vNW29a]. Eigenwerte [vNW29b]. Eigenwerten [vNW29a]. Eigenwerteproblem [vN28d]. Eigenwerttheorie [vN28a, vN29a]. Eindeutigkeit [vN31c]. einem [Tur60a]. Einführung [vN23, vN33]. Einige [vN28f, vN32a]. einiger [vNW28a, vNW28b, vNW28c]. Einstein [Hol18, Khr08, Reg87, vN35h]. Einweihung [B^{+}89]. Eisler [CW91]. Ekman [KvNS63]. Electron [Neu33, vN28f, vNW28a, vNW28b, vNW28c]. Electronic [Ano57a, BGvN82, GvN47a, LR67, SW80, Ste80, vN47a, Ben16, BGvN46, GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, Tar58, vNG48a, vNG48b]. elefántcsonttoronyban [Neu89b]. elektron [Neu33]. Elektrons [Neu33]. Elements [Fit99, Fit13, vNS35]. elért [SI^{+}79]. élete [Kov03c, Sze76, SI^{+}79]. életébol [MS73]. életének [Szé88, Szé03b]. életútjáról [Döm03]. élhet [Neu89b]. Elimination [GvN47h, Grc11, Grc12]. Elizabeth [Kor86]. elme [Mor02]. elméletében [Neu33]. els [Szé88]. elso [Sze03a, Szé03b]. ember [Sze76]. emberi [Von90b]. embodies [Cze08]. Emergence [Kay97, MJ09]. Emerging [MCP23]. Emerson [Ano91]. Eminent [Ano57a]. Emlékeim [Neu03]. emlékek [Ném84]. emlékezete [Kov03b, Pal87]. Emlekkepek [Kov87a]. Emulator [Bru07, FS11]. Encyclopedia [CFK^{+}91, Cra98]. End [CW91]. Energy [Seg85, vN57b, vN55c, vN46a]. Engineering [Gra63, HPR14b]. engineers [Ake07, Ano96]. Engines [Ano91, CW91, Shu84a, Shu85a, Shu96a]. Eniac [Asp83, Ber00, HPR14a, HPR14b, McC01, MRvN50, Rei50, She12, Ste81, MC00]. Enough [CFK^{+}91]. Ensembles [Pos95a]. Entire [vNxxb]. entropies [DD13]. Entropy [Chu21, Hen03, Pet01, ABS11]. Entwicklung [vN54a]. epic [Rob12]. Epistemological [GRZ99]. Epistemology [Dal01]. EPR [Pos95b]. Equation [vN37f, vN40a, vN51c, CFvN50, LY13, TVvN34, vN59b]. Equations [CN47, Fri80, vN35h, vN10, ABS96, BvN35c, BvN50, vNR47a, vN51d]. equilibrium [vN45b]. equivalent [vN53a]. Era [MCP23]. eredményeinek [SI^{+}79]. erfand [Bau98a]. Erfinder [Lie11]. Ergodensatzes [vN29b]. Ergodic [Tau61b, vN29b, vNU45, Hal58, Zun02, vN32c, vN32d, von10]. Erhard [Bau98b]. Erklärung [vNW28a, vNW28b, vNW28c]. Erratum [Has10b]. Error [Hon09, Wil71, vNK40, Kow15]. Errors [DK41]. Erzsébet [Kor86]. escape [Mar06]. Essays [ADS83, Bur70, MHR80, GC97, Sel92, War08, Wig67b]. Establishment [Cob72]. Estimates [GvN47h]. Estimation [Kat60, vNK40]. Euclidean [LvN47]. Eugene [Hor97, WS92, Wig96c, WS03]. Euler [Jon04, Jam02]. EUR [Emc02]. Europe [FB69]. European [Wah96]. Evading [Sil23]. Evaluation [CN47]. éve [Ano93, Bal04, Csá03, Kov03b, Szé78, Sze97, Zád83]. évei [Lax03]. évek [Lax03]. Event [Fie14]. Ever [Fra05]. Everett [Byr10]. éves [Sip05]. évfordulóján [Sze83b]. evolution [Shu96a, Von90a, War08]. évvel [Ano95]. Exact [iP13]. example [Asp89c]. Examples [BN71, vN36b]. excursions [Hol18]. Executive [Döm03]. Exiles [Fra05]. Existence [vN10, vN51d]. Expansion [vNT55]. expected [ABS11, DG17]. Expelled [MP01]. Experimental [HSS09, GRZ99]. experiments [vNT45]. explain [Kow15]. Explanation [vNW28a, vNW28b, vNW28c]. exponential [FR96, Fer99, Wal96]. Exposition [Ame52]. Extension [KvNS63]. Extensions [Mon79]. Extracting [Per92]. extraordinary [Hal90]. Extremely [Hon09]. ezelott [Ano95].

Facets [Mac94]. Facetten [Mac94]. Factors [vN61a]. faculty [Mit80]. Failing [Pin77]. Failure [Ano91, CW91]. family [Byr10]. famous [Ném93]. Fasori [Ném93]. Fast [FR96, Wal96]. Faster [Gol83a, Bou18]. Father [Bre04, CFK^{+}91, Kov84, von05, Mar98]. Fauvel [CFK^{+}91]. Feb [Wig80]. February [E^{+}00, Tel57]. fegyverek [Jék03a, Jék03b]. Fejér [Szé88, Szé03b, Szé03b]. félé [Neu33]. felel [Neu90]. felépítése [vN26a]. félideje [Szé88, Szé03b]. Fermi [Ano57a, Kow15]. Festkolloquiums [B^{+}89]. Few [CLBdB89]. fiatal [Frö93, Lad93]. fiction [Cas03]. Field [Gol83b, CvN42, CvN43, CvN44a, CvN44b, Sze83a, Sze87, vN35h]. Fields [Ano57a]. Fifteenth [CFK^{+}91]. fifty [Tar77]. fifty-nineth [Tar77]. Finite [vN42a, CvN44a, LY13]. First [Asp83, Knu70, MC00, McC01, SW80, vN45a, vN48a, vN82, vN93b, vN21, Ano87, Ben16, MRvN50, PA13, RH00, Sze03a, Szé88, Szé03b]. Five [vN52b, Har06b]. Fixed [vN37f, B^{+}89]. Fixed-Point [vN37f]. Fixpunktsatzes [vN37f]. fizika [Ano93]. fizikus [Sze03a]. Flagellar [Lig76]. Flat [Hon09]. fled [Mar06]. Floating [Ste74]. Flow [vN48a, vN48d, CF48]. fluctuations [CvN42, CvN43]. Fluid [vN35h]. Flying [vNT45]. fon [Dan90]. fondamenti [Lup18, Lus19]. fondements [vN47c]. fontos [Kov03b]. Footnote [Gol60]. footsteps [Érd85]. force [CvN44b]. forces [CvN44a]. Foreword [Kal13]. Formal [vN31d]. formalism [JvNW34, Réd96, vN36d]. formalistische [vN31d]. Formative [Von89, Von90a]. Forms [vN29e]. FORmula [Lor19]. Formulae [Coo57]. Formulation [Mar65]. Forschung [B^{+}89]. Forschungsinstitut [B^{+}89]. Forschungskollegs [B^{+}89]. Fortran [Lor19]. forty [Sze97]. foundation [Mat03]. Foundational [For10]. Foundations [Dic03, Emc97, Emc02, GL23, HvNN27, Mur04, vN32b, vN47c, vN49c, vN55d, vN64, vN67, vN68, vN71a, vN96b, vN18, Ben03, DJ13, Lup18, Lup03, RS01b, RS01a, vN49b, vN91c, vN96a]. Founders [ADS83]. founding [Mar98]. Four [Lan93b, Lan99, Lan11]. Fourier [vNS39, vNS41]. Fraction [vNT55]. Fragen [vN28h]. France [CW91]. Free [Sil23]. Freedom [Ano57a]. Frege [vH67]. French [Ano96, Pou03, vN28g, vN47c, vN96d]. Frenkel [CFK^{+}91]. friction [CvN43, KvN97]. Friedrich [Has10b, Has10a, ABES12, Has10a]. Friedrich-Wilhelms-Universität [Has10b, Has10a]. Frobenius [DH55]. Fuchs [Ber10]. Fuller [CFK^{+}91]. Functional [Bac78, Bac87, vN28a, vN29a, vN29f, vN31e, vN32e, vN32f, vN35a, vN51a, vN87]. Functionaloperatoren [vN28d]. Functions [Hei77, Tau61b, vN31a, vN31e, vN34a, BvN35a, BvN35b, vN34b]. Fundamentos [vN49b, vN91c]. Fungus [CW91]. Funktionaloperatoren [vN29f, vN28a, vN29a, vN31e, vN32e, vN32f]. Funktionen [vN31a, vN31e]. funktsional’nomu [vN87]. further [vN42b]. Fusion [MJ09]. Future [Bha22, Lew21, vN47b, vN51f, Kum21].

G [Ano91]. Gallery [Pai00]. Game [Cor11, DH55, EMN89, Eri15, IM09b, Kal13, Leo06, Mac92a, Mac92b, Sim03, Twe93, vN48b, Del95, IM95, IM01, IM08, Leo10, Mac99b, Pou93, Pou03, Wei92, vN53a, Sze83b]. Games [Bon04, Ebe67, Haw45, Hur48, K.44, Kap95, Tau63c, vN28i, vN28g, vNM44, vN46b, vNM47, vNM53b, vNM61, vNM64, vNM67a, vNM67b, vNM73, vNM74, vNM04, Bar59, BvN50, KT58, Rel92, vNM53a, vNM70, vNM80, vNM07]. Garden [Dys13]. Gas [vN35f]. Gas-Degeneracy [vN35f]. Gaseous [Ano51]. gauge [FKC86]. Gaussian [Grc11, Grc12]. GBP [Emc02]. Gebiet [Sze83a]. Geeks [Isa14]. Gehirn [vN60b, vN65, vN70, vN80, vN86, vN91b]. Genauigkeit [Neu33]. General [Sma83, vN28h, vN29g, vN29i, vN41c, vN46b, vN51e, vN66, vN96d, PJW^{+}22, vN45b, vN51g, vN63b, Tur60a, vN28a, vN29a]. générale [vN96d]. Generalization [vN37f, JvNW34, vN36d]. Generalized [FKC86]. generating [Mon79]. generation [LLY22, iP13]. Generator [JK99]. Generators [Bre08, Gal93, Wal96]. Genese [CW91]. Genesis [Leo06, Del95]. Geneva [Yor87]. Genies [Mac94]. Genius [Fra07, Kov03d, Mac92a, Mac92b, Mac94, Pai00, Vei19, Ano03, Bal12, Cze08, Kar04, Mac99b, Reg87]. Geniuses [Isa14, Lan93a, Mac72, Mac86]. Geometries [vN37b, vN36b, vN37a, vNH81]. Geometry [CEE^{+}51, Tau62, Veb37, VG55, vNS39, vN52b, DK04, Hal85, vN36a, vNS41, vN51a, vN60a, vN98]. Georg [ABES12]. George [CK12, Zun02]. Gerdes [CFK^{+}91]. German [Bau98a, Bau82, B^{+}89, Cso04, FvN22, Has10a, HvNN27, LS83, Lie11, Mac94, Neu33, Rév83, Sch91, Sze83a, Tur60a, vN23, vN25b, vN25a, vN26b, vN27d, vN27a, vN27e, vN27b, vN27c, vN28a, vN28b, vN28h, vN28f, vNW28a, vNW28b, vNW28c, vN28d, vN28e, vN28i, vN28c, vN29c, vN29a, vN29f, vN29g, vN29b, vN29e, vNW29b, vN29h, vNW29a, vN29d, vN29i, vN30, vN31a, vN31b, vN31c, vN31e, vN31d, vN32e, vN32f, vN32a, vN32b, vN32h, vN32g, vN32i, vN33, vN34c, vN35b, vN37f, vN49c, vN51e, vN54a, vN60b, vNM61, vN65, vN67, vNM67a, vN68, vN70, vN71a, vNM73, vN80, vN86, vN91b, vN96a, vN96b]. Germany [BR94, E^{+}00, Has10a]. Gesamtheiten [vN27b]. Gesellschaftsspiele [vN28i]. Gesture [Man17]. gewisser [FvN22]. Giants [FS12]. Gift [MP01]. gifted [Kov03a]. Gimnáziumtól [Ném93]. gioco [IM95, IM08]. given [Ano91]. Gleichmässig [vN25a]. Gleichungssystem [vN37f]. glimpse [Stö07]. Glory [CW91, Ano91]. Go [Fra05]. Gödel [Bre99, CR03, Har89a, Har89b, Hol18, vH67, vN51h]. Going [HSS09]. golden [Mar96]. Goldstine [Ran74, God74, Tro75, Tro76, TN85]. Good [CLBdB89]. got [Reg87]. Government [Ano57a, CW91]. gradient [KvNS63]. Grain [vN52a]. Grammar [Ném93]. grands [Ano96]. Graphics [SvNG89]. Grass [vN13]. Grasshopper [vN13]. gravitational [CvN42, CvN43, CvN44a, CvN44b]. Great [Bla57, Lap96, Mar06, Ano96]. Gregory [Hei77]. Grenzen [Neu33]. Ground [Ber00, Sip05, Sze76]. ground-breaking [Sip05, Sze76]. Group [Isa14, Tau61b, UvN47a, vN34a, vN34b]. Groups [vN29c, vN33, vN39, vN40b, BvN35a, BvN35b, SvN50, vN27e, vN34c, vNW40]. Grundlagen [Ber69, HvNN27, vN32b, vN49c, vN64, vN67, vN68, vN71a, vN96a, vN96b, Kri95]. Grundlegung [vN31d]. Gruppen [vN29c, vN27e, vN33, vN34c].

H [Ano91, DK04, God74, Ran74, Tro75, Tro76, vN32g, Ano57a]. H-Bomb [Ano57a]. Haar [vN34c, vN36g]. Haarschen [vN34c]. Habilitation [Has10b, Has10a, Has10a]. Hackers [Isa14]. hagyatékából [Szé88, Szé03b]. hagytak [Kőr99]. half [Her88, Her95, Szé88, Szé03b]. half-century [Her88, Her95]. Hall [CW91]. halmazelmélet [vN26a]. halt [Sze97]. Hamburg [BR94]. Handbook [FBD^{+}22]. hardcover [CK12, Emc02, MC00]. Három [Frö93, Lad93]. Harsanyi [Sel92]. Hart [Kow15]. határai [Neu33]. Hatvan [Sip05]. haz.hu [Tós04]. hazai [Frö93, Lad93]. Heat [CN47]. Heat-Conduction [CN47]. heaven [Cas03]. height [vNR57]. Heims [Bir83, Dys81, Kev81, Hod81, Ste82]. Heisenberg [Bac22]. Held [HFG51, Ame52, Ano51, GIS90]. Help [DW16]. helped [Bre04]. Hempstead [GIS90]. Henry [CFK^{+}91]. Heppenheimer [CW91]. Heredity [Kay97]. Heritage [Emc97, Pet05]. Herman [Ran74, Tro75, Tro76, TN85]. Hermitescher [vN29a]. Hermitian [vN29a]. Herrn [vN31b, vN32g]. heterogeneous [ABS11]. Hidden [Bel66, Bub10, Jam74a, GL01, Mit22]. Hidden-Variable [Jam74a]. Hide [DH55]. High [BMvN46, CW91, GvN47h, GvN51, Kal13, vN42a, vNG47, vN54c, vN81a, vN51f]. High-Speed [vN54c, vN81a, vN51f]. High-tech [CW91]. Hilbert [vN31b, Fri80, Lac00, Mat03, Mit22, Mur04, Réd96, vN27d, vN39]. Hilbert-style [Mit22]. Hilbertian [Lup18]. hilbertiana [Lup18, Lus19]. Hilbertschen [vN27d, vN31b]. Hilfssatz [vN30]. him [HP20]. Hiroshima [Yor87]. Historia [Has10b]. historian [Cer00]. Historical [GL23, HP20, Jam74c, Seg85, Car95, RKH^{+}21]. Histories [MH11]. History [Ano91, CFK^{+}91, Dal01, FBD^{+}22, Lor19, MHR80, Rec07, Ven02, Bre84a, Del95, ET07, Phi88, RW12, RH00, RMK94, Shu84a, Shu85a, War08, Wei92, CFK^{+}91, Wei92]. Hitler [Mar06, MP01]. Hixon [Jef51]. Hofstra [GIS90]. Hogyan [Kov84]. Homeomorphisms [UvN47a]. Homer [MS18]. honor [Sel92]. honoring [DK04]. honour [GC97]. horizontal [KvNS63]. House [Bir97]. Houtermans [ABES12]. Howell [CLBdB89]. Hugh [Byr10]. Human [Wal16, Von90b]. hundred [Bal04, Csá03]. Hungarian [Ano93, Ano95, Ano03, Bal04, Ben03, Brü87, Csá03, Cze08, Döm03, Érd85, Fil03, Fra07, Frö93, Gol87b, GW03b, GW03a, Gol03, Gol87c, Göt03, Gyö75, Kov87a, Hal77, Hal03, Has10a, Jék03a, Jék03b, Kar04, Kor86, Kőr99, Kov84, Kov87b, Kov89, Kov94, Kov99, Kov03b, Kov03c, Kov03e, Kov03d, Lad93, Lan93a, Lax03, Mar96, MS73, Mor02, Nag87b, Nag87a, Nag03, Ném84, Ném93, Neu33, Neu64, Neu72, Neu89a, Neu89b, Neu90, Neu03, Neu06, Pal03a, Pal87, Pal03b, Pet05, Rut03, Sim03, Sip05, Sza03b, Sza03a, Szá03c, Szé78, Sze87, Sze03a, Szé88, Szé03b, Sze83b, Sze83c, Sze97, Sze76, SI^{+}79, Var94, Ven03, Von90b, Wig67c, Zád83, Zal03, vN25a]. Hungarians [Lan11]. Hungary [Wah96, Mac72, Mac86, Tót93]. Hurrah [SvNG89]. hydrodynamic [vNR50]. Hydrodynamical [BMvN^{+}59, vN44b]. Hydrodynamics [BFvN^{+}44, Lig76, Tau63c, vN45d, PPvNF44, vN44c]. hypothesis [vN32c, vN32d].

IAS [FS11]. IAS/von [FS11]. IASSim [FS11]. IBM [Ano91]. ICL [CFK^{+}91]. Idea [Wol16]. ideal [vN26b]. idealen [vN26b]. Ideas [CW91, Lew21, Hor97, IM08, Wol16]. idee [IM08]. Ideology [Kri95]. IEEE [DCH05]. If [Pin77]. IFIP [BR94]. Igniting [Fit99, Fit13]. ihrer [vN29c]. II [Wig96c, BvN35a, BvN35b, CvN43, GvN51, HvN42, Kad58, Kap95, MvN37, SvN46a, SvN46b, Tau61b, vNH41, vNG48a, vNG48b]. III [Byr10, CvN44a, SvN48a, SvN48b, Tau63a, vN40c, vNG48b, vN61a]. Illinois [Har67]. illustrations [Kev81, Pan92]. im [Has10b, Has10a, Lie11]. Image [CFK^{+}91]. images [GC97]. Impact [Dan85, vN55c, vN56a]. Impacts [BR94]. imperfectly [vN41a]. Implementation [Pes95]. Implementing [HPR14b]. important [Kov03b]. impossibility [Car95, Die17]. impossible [GL01]. impressions [Ins12]. In-Memory [MCP23]. inauguration [B^{+}89]. inception [Von90a]. Incompleteness [For10, CR03]. Incompressible [FvN51, FvN55]. Independence [vN62, FKC86]. Independent [vN28e, vN51c]. index [Kev81]. Inductance [vN59a, vN54b]. Induction [vN28h, KvN37]. Induktion [vN28h]. Inequalities [vN37e]. inequality [Mir75]. infinite [vN38]. Influence [Ste80]. Information [CLBdB89, CW91, Kay97, Laf89, Tót93, War08, BR94]. informatique [CW91]. ingénieurs [Ano96]. inner [JvN35]. Innocence [CW91]. innovation [CW91, HD13]. Innovators [Isa14]. insights [Rel92]. inspiration [Ano97, Hal90]. Inspiring [Chu14]. Instability [FvN51, FvN55, vN53d]. instead [Réd96]. Institute [ABS96, B^{+}89, E^{+}00, Fro00, GIS90, Lup03, Asp87, Mit80, Reg87]. Instrument [BGvN82, GvN47a, BGvN46, GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, vNG48a, vNG48b]. Integral [vN35b]. Integraloperators [vN35b]. Integrals [vNS39, vNS41, vN51a]. Integration [vN51c, CFvN50]. integrity [Whi16]. Intel [JK99]. intellectual [FB69]. Intelligence [Müh09, MJ09, Wah96, Luc11]. Interaction [Hei77, Ula80, URR86d, Sel92]. Interdisciplinary [Hei77, Fro00]. interfaces [Ame52]. intermediate [vN44a]. International [BH89, Réd99, Var88]. Internet [Kőr99, Rut03]. interneten [Rut03]. Internetig [Kőr99]. Interpretation [Stö99, Bac22]. Interpretations [FBD^{+}22, Jam74c]. intersection [vN45c]. interval [vN28c]. Intervalles [vN28c]. Interview [TWA^{+}87, TN85]. Introduction [TWA^{+}87, vN33, vN23]. Introductory [vN44a]. intuitionism [CR03]. Invariant [vN36f, vN99]. inventé [Ano96]. Invented [Sla87, SW80, Ano96, Bau98a, Roj94, Smi10]. Inventing [Bre04]. Invention [Pri10]. inventor [Lie11]. Inventors [CLBdB89, Nat88]. Inverting [GvN47h, GvN51, vNG47]. Investigation [Cso04]. Ioan [Jon04]. Irma [Gri03]. Irrisolto [Ven02]. Irving [TWA^{+}87]. ISBN [Ano91, Bow92, Emc02, Jon04, Lup03, MC00, Rec07, Twe93]. isomorphism [RW12, vN58b]. Issues [GC97]. Italian [IM95, IG02, IM08, Lup18, Ven02]. Iterated [Kat60]. Iterating [Per92]. IV [CvN44b, MvN43, Tau62]. ivory [Neu89b]. ix [Emc02]. Izbrannye [vN87].

J [CW91, Dys81, Hod81, Kev81, Réd99, Sch91, Ste82, Tur60a]. J. [Ano57e, DH55, Ebe67, FH55, Haj59, PC06, Sch91, SN57, Tar58, Tur60a]. Jack [vN03]. Jacobi [GMvN59]. Jahr [B^{+}89, Has10b, Has10a]. James [Jon04]. János [Gol87c, MS73, Nag87a, Sza03a, Wig67c, Ano57d, Ano93, Ano03, Bal04, Ben03, Csá03, Cze08, Döm03, Dys12b, Érd85, Fil03, GW03b, GW03a, Göt03, Gyö75, Jék03a, Jék03b, Kar04, Kor86, Kov87b, Kov99, Kov03b, Kov03c, Kov03e, Lax03, Mor02, Nag87b, Nag03, Neu90, Pal03a, Pal87, Pal03b, Pet05, Rut03, Sim03, Sip05, Sza03b, Szé78, Sze87, Sze03a, Szé88, Szé03b, Sze83c, Sze97, Sze76, SI^{+}79, Var94, Ven03, Von90b, Zád83, Zal03]. Jánosból [Kov84]. Jánosrol [Kov87a, Ném84, Neu89a, Sze83b]. Jánossal [Ano95]. Jánostól [Kőr99]. January [DK04, vN57b]. játékelmélet [Sim03]. játékelmélett [Sze83b]. Jelenség [Pal87]. Jennings [CW91]. jeux [Pou03, vN28g]. Jewish [Has10a, Mac72, Mac86]. Jews [Mar06]. Johann [Ano91, CW91, Ber69]. John [All87, Ano91, Ano97, ABS96, Asp83, Bow92, B^{+}89, Cor11, Cso04, DCH05, DK04, E^{+}00, Ham92, Har67, Has10b, Has10a, Haw45, K.44, LS83, Lie11, Nag87b, Tau60, Tós04, Twe93, Joh86, Ádá83, Ano57b, Ano57c, Ano75, Ano88, Ano91, Ano95, Ano96, Ano03, Ano19, Ara90, Asp80, Asp87, AB87, Asp89a, Asp89b, Asp89c, Asp90b, Asp90c, Bal12, Bal04, Bau82, Bau98b, Ben03, Ber10, Bha22, Boc58, Bre84a, Bre99, Bre04, Bri89, BH89, BS04, Bro73b, B^{+}89, Bur61, Car95, CR03, Csá03, Cso04, Cze08, Día05, Dic03, Döm03, DCG89, Eck87, Érd85, Fie15, Fil03, For51, Fre13, Fro00, Gal93, GL01, GIS90, GW57, Gol83b, GW96, GW03b, GW03a, Gol87c, Göt03, Grc05, Grc11, Grc12, Gri03, Gyö75, Kov87a, Hal73, Hal91, Hal04, Hal85]. John [Hal87, Hal90, Har42, Has10a, Hei80, Hon09, Hor97, How88, Hur48, IM95, IM01, IM08, IM09b, IM13, Jék03a, Jék03b, Kac82, Kah97, Kar04, Kje01, Kor86, Kőr99, Kov84, Kov87b, Kov94, Kov99, Kov03a, Kov03b, Kov03c, Kov03e, KT58, Kum21, Lan08, Lax90, Lax03, Lax05, LS83, Lie11, Lig76, Luc11, Lup18, Mac92a, Mac92b, Mac94, Mac99a, Mac99b, Mad12, Mar98, Mat03, MS73, Mir75, Mon06, Mor02, Müh09, Mur04, Nag87a, NAH^{+}89, Nag03, Nas98b, Ném84, Neu89a, Neu90, NPP66, OPP58, OPP88, Pal03a, Pal87, Pal03b, Pan92, Pel03, Pet05, Pin77, Pou93, Pou03, Ras07, Réd59, Réd96, RS01b, RS01a, Réd05, Rel92, Rév83, RKH^{+}21, Rut03, Sel92, Sim03, Sip05, Smi10, Smi59, Ste80, Stö99, Stö04, Stö07, Sto83, Sza03b, Sza03a, Szá11, Szé78, Sze83a]. John [Sze87, Sze03a, Szé88, Szé03b, Sze83b, Sze83c, Sze97, Sze76, SI^{+}79, Tar72, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Tel70, Tel57, Tho89, Tod74, Tom57, Tót93, Ula58a, Ula58b, UKTS68, UKTS69, Ula82b, Ula82a, URR86a, URR86b, Val08, Var94, Veb37, VG55, Vei19, Ven03, Vol06, Von87, Von88, Von89, Von90a, Von91, Von90b, Whi16, vNW90, Wig57, Wig67a, Wig67c, Wig80, Wig96a, Wig96b, WG96, Yee82, Zád83, Zal03, Zun02, vN96c, vN05, Bir83, Bir97, Dys81, Emc02, HB82, Hod81, Kev81, Kid92, Lup03, Lus19, Mah93, Rec07, Ste82]. John-von-Neumann-Professuren [B^{+}89]. Johnny [Dys13]. Johnson [Ano91]. joint [Fro00]. Jordan [DJ13, Lac00]. Journey [Kje01]. jöv [Brü87]. Judgments [Has10a]. jüdischen [Has10b, Has10a]. juego [IM01]. Juli [B^{+}89]. Jülich [E^{+}00, Fro00]. July [Ano97, B^{+}89, HFG51]. June [ABS96, GIS90, HFG51, RMK94].

Kann [Tur60a]. Kármán [Hor97]. Kelly [CFK^{+}91]. Kenéz [Gol87c, Gol87c]. Keyboards [CFK^{+}91]. Khwarizmi [CFK^{+}91]. Ki [Kov99]. Kis [Nag87b, Nag87b]. klassischen [vN32h, vN32i]. Klaus [Ber10]. Kluwer [Emc02, Lup03]. Knowledge [Lea55, MWW77, Wal16, Cas03]. Knuth [CFK^{+}91]. kongruente [vN28c]. König [DH55]. Konrad [Lie11]. kontinuierlicher [vN27e]. Koopman [Sil23]. koordináta [Neu33]. koordináta-mérés [Neu33]. Koordinatenmessungs [Neu33]. Koordinatenmessungs-Genauigkeit [Neu33]. korai [Lax03]. korszakalkotó [Sip05]. korunk [Ven03]. közgazdaságtan [Zal03]. Közr [Kor86]. közremuködése [Göt03]. Kruh [CFK^{+}91]. Kuhn [CW91]. különböz [SI^{+}79]. Kummer [vNG53]. Kurt [Bre99, vN51h]. kvantumelmélet [Nag03]. kvantummechanika [Ben03].

l [Érd85, Sze83b]. Laboratory [Lan93b, Stö07, Lan99]. Lage [FvN22].

Landed [Lan93b, Lan99]. Landmark [GGC^{+}05]. Language [Lor19]. Languages [Kah97]. Large [GvN63, vNG46, vN48b, vN81b, vN88, GvN45]. Large-Scale [GvN63, vN81b, vN88]. Last [SvNG89]. László [Kov03a]. late [Ula58b]. Later [FS12, For10]. Lattice [vN40b, Bir58, Fro00]. Lattices [vNH37]. Lausanne [Ano97]. Leaders [Bir77]. Leading [Pin77, Lax03]. leaves [Gyö75]. Leben [Bau82, LS83]. Lebensbegriff [Cso04]. Lebesgue [vN36f]. Lecture [Kah97, For51, Lig76, TWA^{+}87]. lecturers [Ano30]. Lectures [BFvN^{+}44, B^{+}89, Ano87]. left [Kőr99]. Legacy [GIS90, Müh09, Ara90, Pal03b, Szé88, Szé03b, Von90a]. Legend [Hal73, Hal91, Hal77, Hal03, Hal04]. legenda [Hal77, Hal03]. Lemma [vN30]. Leo [Fra05, Hor97, Lan89]. Leopold [Szé03b]. Lesniewski [vN31b, vN31b]. lett [Kov84]. Letter [TWA^{+}87, vN03]. Letters [Szé03b, Lan08, Réd05, vN05, Rec07]. levelei [Gyö75]. Levelek [Szé88, Szé03b]. li [Tur60a, Tur60b, TvN99]. li.vi.5 [CFK^{+}91]. Liberated [Bac78, Bac87]. Library [CFK^{+}91, Tós04]. Licklider [CW91]. Lie [SvN50]. Life [Bau82, Bha22, CW91, Chu14, Cso04, Dys81, HB82, Hod81, Hon09, Kac82, Kay97, Kev81, Kov03c, LS83, MS73, Ste82, von05, ABES12, Döm03, Hei80, IM08, Kum21, Mar98, PC06, Sto83, Szé03b, SI^{+}79, Yee82]. lifetime [Szé88]. Light [Fit99, Fit13, Nag87a, Von90a]. like [Réd96]. liked [Réd96]. limitations [FKC86]. Limits [Neu33, Cas03]. Lindgren [Ano91, CW91]. Linear [BMvN46, Dan85, JML13, KvN97, SvN46a, SvN48a, vN29c, vN29e, vN39, vN59a, JvN35, SvN46b, SvN48b, vN54b, vN59b]. linearer [vN29c]. Linearformen [vN29e]. Linguistics [Var88]. link [DD13]. Lipot [Szé88, Szé88, Szé03b]. Liquids [FvN51, FvN55]. list [Ano57e]. live [Neu89b]. lives [Wol16]. Living [Hon09]. located [vN41a]. Loewner [WvN54]. Logic [Pri10, Ven02, Ara90, Asp80, BvN36, Réd01, vH67, Tau61a]. Logica [Ven02]. Logical [BGvN82, GvN47a, Hol62, Tur60a, vN41c, vN96d, BGvN46, vN51g, vN63b]. logiceskaja [Tur60a]. Logics [vN37d, vN56b, vN63c]. logique [vN96d]. logische [Tur60a]. Logistics [vN37d]. London [Bir83, Bow92, Emc02, Ran74, Rec07]. looks [Cer00]. Lüders [Khr08]. Lundstrom [CLBdB89]. Lupacchini [Lus19]. Lyle [Ano91].

M [Fie15, vN32g]. M.I.T [Bir83]. MA [Asp83]. Mach [vNR57]. Machine [Kem55, Lor19, Pes95, HP20, Her88, Her95, Tur60a, TvN99, FS11, Tod74, vN49a]. Machinery [Ano91, CW91]. Machines [CFK^{+}91, GvN63, Pri10, vNG46, vN54a, vN81b, vN88, GvN45, Tar58]. Macrae [Bir97]. MADDIDA [TWA^{+}87]. Made [Eri15]. Maestro [CFK^{+}91]. Magyar [Nag87a, Ano03, Kov03e]. Magyarok [Kov89]. Mahoney [CK13]. mainframes [Shu96a]. makers [Wol16]. Making [Yor87]. Man [Bha22, Bro73a, CFK^{+}91, Gra56, Kem55, Kum21, Mon06, Smi10]. Managing [CW91]. Manhattan [Ben16, HHW99, Lan93b, Lan99, Lan11]. Maniac [Lab23, Bre04]. manuscript [RKH^{+}21]. Many [Ano57a, CFK^{+}91, Bre04, Byr10, vN28c]. Mapping [vN51c]. Mappings [vNH37, vNH40]. mark [Kőr99, Stö04]. marking [Tót93]. Markov [JML13]. Marling [CFK^{+}91]. Married [Gra56]. Marshall [DK04]. marslakó [Har06a, Kov03b]. Martian [Kov03b, vNW12]. Martians [Har06b, Lan93b, Lan99, Mar96]. Maryland [DK04]. Maschine [Tur60a]. Maschinen [vN54a]. mashina [Tur60b, TvN99]. masina [Tur60a]. Mass [Bir83, Bow92, vN31a, vN34c]. Massachusetts [ABS96, ABS96]. Maße [vN31a]. Masses [Ano51, vN29g]. Maßes“ [vN29i]. Matematica [IG02, Lus19]. matematico [IM95, IM08, IM01]. matemáticos [vN49b, vN91c]. matematikus [Szá03c]. Material [HKLS93]. Mathematica [Has10b]. Mathematical [Cor11, GvN47a, IM09b, Jon04, Kje01, Rec07, vN27a, vN47c, vN49c, vN54a, vN55d, vN18, Asp80, Asp87, Bal12, IM95, IM01, IM08, KT58, RKH^{+}21, Stö04, Val08, vH67, Mar65, Tar77, vN32b, vN49b, vN64, vN67, vN68, vN71a, vN91c, vN96a, vN96b]. Mathematician [Ano57a, Has10a, Bre04, Mon06, Szá03c, Szá11, vN47d]. Mathematicians [Hei77, Jon04, Réd99, Jam02]. Mathematics [Bir77, Die17, Dys81, GM04, HB82, Hod81, IG02, IM09a, Kac82, Kah97, Kev81, Mur04, Rec07, Ste82, Ula80, URR86d, vN31d, CR03, Del95, Día05, FF91, GGC^{+}05, Hei80, Mah11, Phi88, Réd99, RMK94, Stö04, Sto83, Tar77, URR86c, vN54e, B^{+}89, Mac94]. Mathematik [B^{+}89, Mac94, vN31d]. Mathematiker [Has10b, Has10a]. mathématiques [vN47c]. Mathematische [Ber69, vN27a, vN32b, vN49c, vN64, vN67, vN68, vN71a, vN96a, vN96b]. mathematischer [vN54a]. Matrices [GvN47h, GvN51, GMvN59, Kat60, vN42a, vNG47]. Matrix [vN37e]. Matrix-Inequalities [vN37e]. Matrix-Space [vN37e]. Matrizen [vN29h]. Mauchly [Asp83, Ano96, SW80, Ste81]. maximization [vN47e]. maximum [vN63a, vN74]. May [GIS90]. McCartney [MC00]. McKay [Ano91]. McNeil [CFK^{+}91]. mean [Har42, vN41b, vNKBH41, vN42b]. meaning [Hor97]. Measurable [vN32a]. Measure [vN29g, vN29i, vN35a, Hal58, vN34c, vN36g]. Measurement [Jam74b, Neu33, Ara90, Bub01, FKC86, Mol72]. measurements [Khr08]. Measures [vN36f, vN99, vN51a]. mecánica [vN49b, vN91c]. Mécanique [Ber69, vN47c]. meccanica [Lup18, Lus19]. mechanical [Cso04, JvNW34, vN27b, vN36d]. Mechanics [Bel66, Cob72, Emc97, GL23, HvNN27, Jam74c, Lac00, Mar65, Pos95a, Sil23, Tau61a, vN27a, vNW28a, vNW28b, vNW28c, vN29b, vN32b, vN32h, vN32i, vN35g, vNH41, vN47c, vN49c, vN55d, vN64, vN67, vN68, vN71a, vN96b, vN18, Ben03, Ben72, BvN36, Bre99, ET07, HvN42, Lup18, Mit22, Réd96, vN27c, vN49b, vN91c, vN96a, von10]. Mechanik [vN29b, vN32h]. Mechanik“ [vN32i]. mechanischer [Cso04]. mechanisms [Jef51]. Medal [Ano57a, DCH05, Fie15]. Meet [Ano30]. meeting [PA13, Tar77]. Meetings [Hur85]. meets [Stö04]. Még [Neu89a, Sze97]. megalapozása [Ben03]. megalkotója [Kov03b]. megtestesítoje [Cze08]. meine [vN31b]. meltdown [Byr10]. members [Mit80]. memoir [vNW12]. memorabilia [Hal87]. Memorandum [vN45d]. Memoriam [CW91]. memories [Ném84, Kov87a, Neu03]. Memory [Bon04, MCP23, vN59a, Bou18, Kov03b, Pal87, vN54b]. Men [CLBdB89, Str63a]. Menge [vN31a]. Mengenlehre [vN25b, vN28b, vN28h, vN29d]. mérés [Neu33]. merkwürdige [vNW29b]. message [Neu90]. messbare [vN32a]. metal [Ame52, Ame52]. metallurgical [vN52a]. Metamatematica [Lup18]. metamathematics [Lup18]. Metaphysical [Pin77]. meteorológiában [Göt03]. Meteorology [Dal01, Tau63c, vN47b, Göt03]. Method [CN47, Eck87, GvN47h, GMvN59, HKLS93, Ras07, vN44c, vN48b, vN51c, vN55e, Ano93, Ben16, For72a, For72b, HFG51, Mon79, vN44b, vNR50, vN54d]. Methoden [Sze83a]. Methodology [For10]. Methods [RvN85, vN32h, vN32i, vNH41, vN45d, HvN42, KW64, RUvN47, Sze83a, Sze87, vNR47b]. Metric [vNS39, JvN35, vNS41]. Metrization [vN37e]. Michael [Ano91, Emc02, Lup03, CK13]. microprocessors [Shu96a]. Microsystems [CW91]. migration [FB69]. Miklos [Lup03, Emc02, Rec07, Dic03]. Mind [Bla57, DJ13, Gra56, Mor02, Nas98a, Hey47, Shu84a, Shu85a, Shu96a]. minded [Sze76]. Minimally [vNW40]. Minimax [BEMD11, Kje01]. minimum [FvN22]. Minimumpolynome [FvN22]. Minkowski [vN29e]. Minkowskischen [vN29e]. mint [Cze08]. Miracle [HPR14b]. mix [Bou18]. Mixtures [CEE^{+}51]. Model [vNxxa, Cze08, vN45b]. Modeling [Sma83]. modell [Cze08]. Models [Dal01]. Modern [ADS83, Asp90b, Bow92, CW91, DW16, Grc11, Grc12, HPR14b, Ham92, Kid92, Mac92a, Mac92b, Mah93, Pan92, Wil71, Ano91, Asp87, DCG89, Kov03b, Mac72, Mac86, Mac99b, Mad12, RMK94]. Modern-Day [DW16]. modified [LY13]. módszere [Ano93]. módszerek [Sze87]. Modular [vNH37]. Molecular [E^{+}00]. mondo [IM95, IM08]. monográfia [Neu89a]. monograph [Neu89a]. Monte [JML13, Ben16, Dys12a, Eck87, HPR14a, HFG51, Hur85, KW64]. Morgenstern [Ebe67, Hur48, Haw45, K.44, Leo10, Mar46, Mir92, Rel92]. Motion [Ano51, KvN97]. Mounier [CW91]. Mounier-Kuhn [CW91]. Move [Gra56]. Mozet [Tur60a]. Mozhet [Tur60b, TvN99]. Mr. [vN31b, vN32g]. MR2609069 [Has10b]. Ms [CFK^{+}91]. MSRC [Ber00]. Much [Mac92a, Mac92b, Mac99b]. Müller [Ano91, CW91]. Multiple [Fra05, Byr10]. Multiplicity [vN10, vN51d]. mundo [IM01]. munkássága [GW03b, GW03a, Kov03c, Sze87, Sze76, SI^{+}79]. mutual [Byr10]. My [CW91, Neu03, vN31b]. myslit [Tur60a, Tur60b, TvN99]. mystery [Bal04]. myth [Mar96].

Name [Cip00]. named [Ano30]. Nancy [Asp83]. Nash [Ano91]. nation [Tar77]. National [Ame52, Ano88, Lan93b, Joh86, Lan99, Tod74]. natural [Ake07, Cso04]. naturalized [Sch88]. nature [CR03]. Naturphilosophie [Cso04]. Nazi [MP01]. near [KvNS63]. near-neutral [KvNS63]. neat. [HP20]. Negyven [Sze97]. Neiman [Dan90]. nem [Neu89b]. network [PJW^{+}22]. networks [ABS11]. neuen [vN29b]. neuerer [vN54a]. Neumann [All87, Ano57c, Ano88, Ano91, Ano97, ABS96, Asp83, Bau98a, Bau82, Ber69, Bir83, Bow92, B^{+}89, CLBdB89, CFK^{+}91, CW91, Cor11, Cso04, DCH05, Dic03, DK04, Dys81, Ebe67, Emc02, E^{+}00, FS11, Fie15, Fro00, GIS90, Ham92, Har67, Has10b, Has10a, Haw45, HB82, HKLS93, Hod81, Hur48, Jon04, K.44, Kev81, Kid92, LS83, Lig76, Lup03, Lus19, Mac94, Mah93, MS73, NHN89, Nas98b, OPP58, Pan92, Pou03, Ran74, Rec07, Réd99, Rév83, Sch91, Ste82, Sze83a, Tós04, Tót93, Tur60a, Twe93, Wig57, Wig67a, Wig80, Wig96b, Joh86, Acu21, Ádá83, Ali09, ABS11, Ano57a, Ano57b, Ano57d, Ano57e, Ano75, Ano93, Ano95, Ano96, Ano03, Ano19, Ara90, Asp80, Asp87, AB87, Asp89a, Asp89b, Asp89c, Asp90b, Asp90c, Bac22, Bac78, Bac87, Bal12]. Neumann [Bal04, Bar82, Bau98a, Bau82, Bau98b, Bec04, BEMD11, Ben03, Ben72, Ber10, Bha22, Bir58, Bit97, Boc58, Bon04, Bre84a, Bre84b, Bre99, Bre04, Bri89, BH89, BS04, Bro73b, Bub01, Bub10, Bur61, CK12, Car95, Cer00, Chu21, Cin85, CR03, Csá03, Cso04, Cze08, Del95, DD13, DG17, Día05, Die17, Döm03, DCG89, DJ13, Dya71, Dys12b, Dys13, Eck87, Emc97, Érd85, FS12, FKC86, Fie14, Fil03, For10, For51, For72a, For72b, Fre13, Fri80, Frö93, FH55, Gal93, GL01, GIS90, GH93, GW57, Gol72, Gol83a, Gol83b, Gol87b, Gol93, GW96, GW03b, GW03a, Gol03, Gol87c, Göt03, Grc05, Grc11, Grc12, Gri03, Grz89, Gyö75, Kov87a, HP20, Haj59, Hal58, Hal73, Hal77, Hal91, Hal03, Hal04, Hal85, Hal87, Hal90, Har42]. Neumann [Har89a, Har89b, Has10a, Hei80, Hen03, Hep90, HKLS93, Hon09, Hor97, How88, IM95, IM01, IG02, IM08, IM09a, IM09b, IM13, Jam02, Jék03a, Jék03b, JML13, Kac82, Kah97, Kap95, Kar04, Kat60, KS84, Khr08, Kje01, Knu70, Kor86, Kőr99, Kov84, Kov87b, Kov94, Kov99, Kov03a, Kov03b, Kov03c, Kov03e, Kow15, Kri95, KT58, Kum21, Lac00, Lad93, Lan08, Lan93b, Lan99, Lax90, Lax03, Lax05, LS83, Leo06, Leo10, LY13, Lie11, Luc11, Lup18, Mac92a, Mac92b, Mac94, Mac99a, Mac99b, Man17, Mar98, Mar65, Mar46, Mat03, McC83, MS18, MS73, Mir92, Mir75, Mit22, Mol72, Mon79, Mon06, MCP23, Mor02, Müh09, Mur04, Nag87b, Nag87a, NAH^{+}89, Nag03, Ném84, Neu89a, Neu90, NPP66, OPP88, Pal03a, Pal87, Pal03b, Pel03, Per92, Pes95, Pet01, Pet05]. Neumann [Pin77, Pos95b, Pou93, Pou03, PJW^{+}22, Ras07, Réd59, Réd96, RS01b, RS01a, Réd01, Réd05, RW12, Rel92, Rév83, RKH^{+}21, Rut03, Sam92, Sch88, Sha58, Shu84b, Shu85b, Shu96b, Sil23, Sim03, Sip05, Smi59, SW80, Ste80, Stö99, Stö04, Stö07, Sto83, Str63b, Sza03b, Sza03a, Szá11, Szé78, Sze83a, Sze87, Sze03a, Szé88, Szé03b, Sze83b, Sze83c, Sze97, Sze76, SI^{+}79, SN57, Tar58, Tar72, Tau60, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Tel70, Tel57, Tho89, Tod74, Tom57, TWA^{+}87, Tur60a, Ula58a, Ula58b, UKTS68, UKTS69, Ula80, Ula82b, Ula82a, URR86a, URR86b, URR86d, Val08, Var94, Veb37, VG55, Vei19, Ven03, Ven02, Vol06, Von87, Von88, Von89, Von90a, Von91, Von90b, vNW90, Wig67c, Wig96a, WG96, Wil86, Yee82, Zád83, Zal03, Zun02]. Neumann [vH58, vNBV95, vN96c, vN03, vN05, Bir97, God74, Tro75, Tro76].

Neumann-legenda [Hal77, Hal03]. Neumannig [Gol87b, Gol03]. Neumannról [Frö93, Lad93, Nag87b]. Neumanns [Sch91, DH55]. neural [PJW^{+}22]. neutral [KvNS63]. Neutron [RvN85, RUvN47, vNR47b]. Never [DJ13]. News [All54, Axe52, Bou18]. next [Brü87]. NIC [E^{+}00]. nine [Mar06]. nineth [Tar77]. No [Bub10, CR03, GL01, Kow15, Neu89a]. nobles [Mac72, Mac86]. Nodded [MS18]. Non [vN54b, vN59a, DK04, vN58b, vN59b]. non-commutative [DK04]. non-isomorphism [vN58b]. Non-Linear [vN59a, vN54b, vN59b]. nondegenerate [Khr08]. nonlinear [ABS96]. Norberg [CW91]. Norbert [Bir83, Dys81, HB82, Hod81, Kev81, Ste82, Hei80, Kac82, Sto83, vN49a]. NORC [vN54c, vN81a]. Normal [vN29f, Dya71, FR96, For72a, For72b, Wal96, vN35e]. normalen [vN29f]. Norman [Bir97]. Northrop [vN03]. notable [Wol16]. Note [Hur85, Har42]. Notes [Ano57c, Axe52, Bur61, vN52b, vN53c, Kev81, vN53b]. Nothing [Cer00]. Novecento [IM95]. November [Ber00]. NP [Har89a, Har89b]. Nuclear [HPR14a, Mac92a, Mac92b, Seg85, Wig96c, Byr10, Jék03a, Jék03b, Mac99b]. nukleáris [Jék03a, Jék03b]. Null“ [vN31a]. Nullstellen [FvN22]. Number [Bre08, Gal93, JK99, vN25a, LLY22, iP13, Réd59]. Numbers [vN28e, vN48b, FR96, Fer99, vN23, vN26b]. Numerical [BMvN^{+}59, CFvN50, CN47, Fro00, GvN47h, GvN51, Grc11, Grc12, Sma83, Tau63b, vNG47, vN48a, vN48b, vN48d, Asp89c, Mad12, Sze83a, Sze87, vN44b, vNR47a, vNR50, vNG53, vN54d]. numerikus [Sze87]. numerischen [Sze83a]. nyomában [Érd85]. nyomot [Kőr99].

O [Ber69, Ebe67]. O. [vN45d]. Obituary [Ano57d, Ano57c]. Oblique [vN43a, vNS43]. Obscaja [Tur60a]. occasion [B^{+}89]. Occur [Fie14]. October [Ame52]. odyssey [Yor87]. office [Reg87]. Oh [Wei88]. ökonomisches [vN37f]. old [Sip05]. One [vNF46, Bre04, Cas03, vN44c, Bal04, Csá03]. one-dimensional [vN44c]. Online [Cra98]. only [Bre04]. ontology [Stö04]. operation [Göt03]. operational [BvN35c]. operational-differential [BvN35c]. Operations [Pri10]. Operator [HvN42, vN32h, vN32i, vNH41, ABS96, DK04, vN43b, DK04, Kad58]. Operatoren [vN51e, vN29f, vN31c]. Operatorenmethode [vN32h, vN32i]. Operators [Tau63a, vN28a, vN28d, vN29a, vN29f, vN31c, vN31e, vN32e, vN32f, vN35b, vNM36, vN39, vN51a, vN51e, DNvN55, Kad58, MvN36, MvN37, MvN43, Mur58, SN57, vN35e, vN36c, vN40c, vN49d, Tau61b]. opinions [Frö93, Lad93]. Oppenheimer [PC06]. opportunism [Stö04]. optimal [vN53a]. optimism [Stö04]. Optimum [DK41, vN41a, vN54d]. Order [BMvN46, GvN47h, GvN51, vN42a, vNG47]. ordinateur [Ano96]. Organisms [vN56b, vN63c]. Organs [vN59a]. Origin [Gal93]. Origins [Ano91, Asp90b, Asp90c, Bow92, CK12, CFK^{+}91, CW91, Grc11, Grc12, GM04, Ham92, Kid92, Mah93, Pan92, Ran82, Asp80, Dys12c, Ran72]. öröksége [Pal03b, Pet05]. orthogonal [vN51a]. Ortiz [vN49c]. Ortvay [Gyö75, Gyö75, NHN89]. Oskar [Haw45, Hur48, K.44, Rel92]. összefoglaló [SI^{+}79, Döm03]. öt [Har06a]. Other [Tau62, For72a, For72b, vN52a]. our [Nat88, Ven03]. Outline [Hol62]. output [iP13]. Outsider [HD13]. Oversight [MS18]. overview [LS83]. Oxford [Ran74, Twe93, FBD^{+}22].

P [Ano96, Wig67b, WS92, WS03, Har89a, Har89b]. pages [Emc02, Pan92]. Palmer [Ano91]. Pantheon [CK12]. Paper [Gol60, vN52a]. paperback [Jon04, Twe93]. Papers [Ano75, AB87, Lap96, Lew21, Ran82, vN96c, Swa76, Wig96c, Tar77, All87]. papucsban [Kor86]. parabolic [vNR47a]. Paradigm [HPR14b]. paradox [Kow15]. Parallel [E^{+}00]. Parameter [vN33]. Parameters [vN33]. Paris [Ano51]. Part [GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, Kad58, Mur58, Wig96c, vNG48a, vNG48b]. Partial [CN47, Fri80, vN51c, ABS96, vNR47a]. Pasadena [Jef51]. Pascal [God74, Ran74, Tro75, Tro76, Gol72, Gol87b, Gol93, Gol03]. Pascaltól [Gol87b, Gol03]. Passages [MS73]. Passing [Bla57]. path [Döm03]. Paul [Hor97]. peace [Yor87]. Pearcey [CFK^{+}91]. Penzias [CW91]. people [Sze76, URR86c, Wol16]. Peres [LLY22, iP13]. Perfect [vN35h]. period [KW64]. Periodic [Tau61b, vN34a, BvN35a, BvN35b, vN34b, vNW40]. permutability [DNvN55]. Perpendicularity [LvN47]. Perrolle [CLBdB89]. Person [vN46b, vN48b, vN53a]. personal [vNW90, Wol16]. Perspective [Jam74c, Seg85, vNH37, Car95, vNH40]. perspectives [ET07, GRZ99, Wol16]. Petre [CFK^{+}91]. phenomenon [Pal87]. philosophical [GC97, Von90a]. Philosophy [Cra98, Jam74c, Cso04, ET07, GC97]. Photon [vN53c]. Photon-Disequilibrium-Amplification [vN53c]. Physical [Bec04, GL23, Wal16, vN32c, FKC86, vN55c, vN55e]. physicist [ABES12, Sze03a, Yor87]. physicists [Har06b]. Physics [Cin85, Kah97, Pin77, Ano93, Ara90, ET07, FF91, GRZ99, Jam74b, RS01b, RS01a, Str11, Whi16, Wig96c, Dic03, Emc02, Lup03]. Pioneer [Kov94, Smi10]. Pioneered [Mac92a, Mac92b, Mac99b]. Pioneers [Kov89, Wei88, Nor98]. Places [Rei50]. Planned [GH93]. Planning [GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, vNG48a, vNG48b]. platonic [Cas03]. Podolsky [Khr08]. Point [Ste74, vN37f, vNxxa, vNxxb, vN41d, vN57a]. Point-Source [vNxxa]. points [CvN44a]. poker [GMvN53]. polihisztor [Kov03d]. polihisztora [Brü87]. Polikarov [GC97]. political [Wei92]. Politics [Leo06]. polymath [Brü87, Kov03d]. polynomials [FvN22]. pontosságának [Neu33]. Popper [Pos95b, Ven02]. populations [vN27b]. Portrait [Pai00]. Position [SW80]. Post [MCP23]. Post-von [MCP23]. postulate [FKC86]. postulates [Khr08]. Poughkeepsie [RMK94]. Poundstone [Twe93]. Power [CLBdB89]. Pp [Bow92, Jon04, Ran74, Twe93, Ano91, Asp83, Kev81, Lup03, MC00].

Practical [CN47, Luc11]. Précis [Ber69]. Prediction [Sma83]. Preliminary [BGvN46, BGvN82]. Presence [DK41]. presented [Tar77]. President [Ano57a]. Press [Ano91, Asp83, Bir83, Bow92, Jon04, Kev81, Pan92, Ran74, Twe93, Dys81, Har67]. price [Ano91]. prices [AFB79]. Priest [Kal13]. Princeton [Nas98b, Ran74, Ano30, FS11, Mah11]. principal [Swa76]. principle [Kov03b]. Principles [GvN63, vNG46, vN81b, vN88, GvN45]. Printed [CW91]. priority [Zun02]. Prisoner [Pou93, Twe93, Pou03]. prisonnier [Pou03]. Probabilistic [vN56b, vN63c, vN27c]. probabilities [Har42]. Probability [Coo57, vN37b, vN37d, Réd01, vNH81]. Probability-Logics [vN37d]. probable [vNK40]. probe [Kow15]. Problem [Bel66, BMvN^{+}59, Har89a, Har89b, vN53d, RW12, vN47e, vN53a, vN63a, vN74]. Problematic [Sam92]. Problems [vN28d, vN48a, vN48d, vN54c, vN81a, GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, KvN97, Réd99, vN44b, vNG48a, vNG48b, Ano51]. Procedure [Per92]. Procedures [Ben72]. Proceedings [Ano51, HFG51, BR94, PA13, RMK94, Var88, GIS90, Wah96]. Processes [vNW29a, vN47f, UvN47b]. processing [Bou18, BR94]. Prodigy [CFK^{+}91]. product [vN44a]. Production [Coo57, MWW77]. products [JvN35, vN38]. Professors [Veb37, VG55]. Professorship [B^{+}89]. Professuren [B^{+}89]. Program [Asp90a, CW91, HPR14c, Knu70, Lac00, Pel99, Sip05]. programja [Sip05]. Programmable [FS11]. Programming [Bac78, Bac87, Dan85, Lor19, Pri10]. Programs [Bac78, Bac87, FS12]. project [Sch88, Ben16, Fit99, Fit13, HHW99, Lan93b, Lan99, Lan11]. projection [Ben72]. projective [TVvN34]. Prologue [BH89]. Prominent [Chu14]. Proof [BEMD11, Bub10, Pin77, vN27d, vN29e, vN32d, von10, Car95, Die17, vN31b, vN29b]. Properties [vN42a, vN43b]. Property [CLBdB89]. Proposal [vN44b]. Proposals [Cin85]. proposed [vNF46]. Prove [Pin77]. Providence [Rec07]. Proving [Ber00]. proximity [Fie15]. Prozessen [vNW29a]. Pruefer [vN26b]. Prüferschen [vN26b]. Pseudo [Gal93]. Pseudo-Random [Gal93]. pseudorandom [Wal96]. Psychology [Leo06]. published [Ano91]. Publishers [Emc02, Lup03]. Pugh [Ano91]. Punched [CW91]. pupil [Kov03a]. pure [Día05]. purpose [PJW^{+}22]. Puzzle [Twe93, Pou93, Pou03].

Quantenmechanik [vNW28a, vNW28b, vNW28c, vN49c, Ber69, HvNN27, vN27a, vN27c, vN32b, vN64, vN67, vN68, vN71a, vN96a, vN96b]. quantenmechanischer [vN27b]. Quantique [Ber69, vN47c]. Quantistica [Ven02, Lup18, Lus19]. Quantity [vN31a]. Quantization [ABS96, DK04, ABS96]. Quantum [Ano57a, Bel66, Cob72, Emc97, Emc02, ET07, FBD^{+}22, GL23, HvNN27, Jam74c, Lac00, Mar65, Pin77, Pos95a, Sil23, Tau61a, Ven02, vN27a, vNW28a, vNW28b, vNW28c, vN32b, vN35g, vN37d, vN47c, vN49c, vN55d, vN64, vN67, vN68, vN71a, vN96b, vN18, Ara90, Ben03, Ben72, BvN36, Bre99, Bub01, DJ13, Fro00, GRZ99, Jam74b, JvNW34, Lup18, Lup03, Mit22, Nag03, Réd96, RS01b, RS01a, Réd01, Str11, Val08, WvN54, vH58, vN27b, vN27c, vN36d, vN49b, vN91c, vN96a, von10]. quantum-mechanical [vN27b]. Quantum-Mechanics-Free [Sil23]. Quantumness [Ali09]. Quasi [GM04, vN32d]. Quasi-Concavity [GM04]. quasi-ergodic [vN32d]. Queries [SW80]. Question [vN29d, Zun02]. Questions [vN28h]. Qui [Ano96].

R [Ano91, CW91, vN49c]. R. [vN49c]. Radio [Ano95]. Rádióbeszélgetés [Ano95]. RAND [War08]. Random [Bir97, Bre08, For51, Gal93, JK99, Per92, vNU45, ABS11, CvN42, CvN43, CvN44a, CvN44b, FR96, Fer99, For72a, For72b, LLY22, Mon79, iP13, vN51b, vN63d]. Randomness [Bar82]. Rank [Bar82]. Rapid [vN47b]. rate [iP13]. Ratio [Bar82, Har42, vN41b, vN42b]. Rational [Coo57, Sel92]. Rátz [Kov03a]. Raumes [vN51e]. Ray [vN35g]. Ray-Representations [vN35g]. Re [Bub10]. Re-Appraisal [Bub10]. Readings [BN71]. Ready [Fra05]. Real [GMvN59]. Reality [IG02, Hor97]. Realization [Sil23]. Realtà [IG02]. Reason [KMM95, CR03]. Received [Ano57a]. receives [DCH05]. reception [Asp87]. Rechenmaschine [vN60b, vN65, vN70, vN80, vN86, vN91b]. Rechentechniker“ [Sze83a]. Rechner [Bau98a, Rév83]. Recollections [Sma83, WS92, WS03]. Reconsidered [Bri89]. Reconsidering [HPR14c]. Record [Ber00]. Redei [Lup03, Dic03, Emc02, Rec07]. Reduction [vN49d]. Reed [TWA^{+}87]. Reexamined [Grz89]. Reflection [vN43a, vNS43, vN45c]. Reflections [For10, Gol87a, HP20, Wig67b]. Refraction [vN45c]. Refugee [Seg85]. regard [RW12]. Regime [MP01]. regular [Tel70, vN36e]. Reid [CLBdB89]. reinterpretation [Mol72]. reject [CR03]. Rejection [Grz89, DG17]. rejtély [Bal04]. Related [TWA^{+}87, vN28h]. Relation [LvN47]. Relativiste [Ber69]. Relativistic [vN35f, vNxxb]. relativity [TVvN34]. Reliable [vN56b, vN63c]. remark [vN42b, vN52a]. Remarkable [Jam02, Jon04]. Remarks [vN28f, vN44a, vN31b]. Remembering [BS04, CK13, Lax90]. replication [Bre84a]. Reply [Hen03]. Report [Asp83, Ebe67, McC83, vN45a, vN48a, vN48d, vN82, vN93b, vN21, Ano88, JC87, SW80]. Reporter [Gol87c, Nag87b]. Repräsentanten [vN31a]. Representation [vN39]. Representations [Kay97, vN29c, vN31a, vN35g, SvN50, vN27e]. representative [vNS35]. Reproducing [Bur61, Har67, Pes95, vNB66, Vol06, vN71b]. Reprogramming [Pos95a]. Research [E^{+}00, GIS90, HSS09, Mac94, Ake07, Ano88, B^{+}89, JC87, B^{+}89]. Researchers [Chu14]. residual [vNS35]. responsibility [Neu90]. Review [All87, Asp83, Ber69, Bir83, Bir97, Bow92, CK12, Cor11, Dic03, Dys81, Emc02, God74, Hal85, Ham92, Har67, Haw45, HB82, Hod81, Hur48, Jon04, K.44, Kev81, Lup03, Lus19, MC00, Mah93, Pan92, Ran74, Rec07, Sto83, Tau60, Tro76, Twe93, vN49a, Kid92, Kum21, Ste82, Tro75]. Reviews [Ano91, CLBdB89, CFK^{+}91, CW91, Kac82]. Revisited [Acu21]. Revolution [Isa14]. Rhetoric [Kri95, KMM95]. rhetorics [Die17]. RI [Rec07]. Richtmyer [HKLS93]. Riemann [vN44c]. Riesz [Fil03]. right [Bre84b]. rigid [KvN97]. Rings [Tau63a, vNM36, MvN36, MvN37, MvN43, Tel70, vN36c, vN36e, vN37c, vN40c, vN43b, vN49d, vN58b]. Riporter [Gol87c, Nag87b]. rise [Ake07]. Road [Chu14, Lax05, Lax03]. Robots [CFK^{+}91]. róla [Neu89a]. Role [Gol83b, SW80, Wal16, vN47b, Khr08, vN54e]. Rombauer [Gri03]. Rose [CW91]. Rosen [CW91, Khr08]. Rossella [Lus19]. Roundup [All54]. routes [HD13]. Routing [PJW^{+}22]. Routledge [Cra98]. Rudolfhoz [Gyö75]. Rudolt [Gyö75]. Rule [vN66]. Russian [TvN99, vN64, vN71b, vN87].

’s [DJ13, Bre08, CK13, Gri03, Sch91, TWA^{+}87, vN52a, BMvN46]. samovosproizvodyashchihsya [vN71b]. sampling [For72a, For72b]. Satz [vN32g]. Sätze [vN32a]. Sätzes [vN29e]. Scale [GvN63, vNG46, vN81b, vN88, GvN45]. Schelling [Fie14]. Scheme [HKLS93, vN53c]. schemes [LY13]. Scheutz [Ano91, CW91]. Schmidt [Bau98b]. scholars [Ins12, Mit80]. School [Mur04, Ném93]. Schrödinger [Bre84b, vN31c]. Schrödingerschen [vN31c]. Science [Ano88, ADS83, Asp89b, JC87, For10, Fro00, IM09b, KMM95, Lap96, Lew21, Pai00, Pri10, URR86c, Wal16, Ano97, Asp80, Fie15, GC97, Leo10, Mar96, Pal87, Sze83b, vN56a, Har06b, Cor11]. sciences [vN54e, vN55c, vN55e, Tót93]. Scientific [Ano91, ADS83, Asp89a, Chu14, Cob72, GW96, Hur85, IM09a, Mac92a, Mac92b, MWW77, NAH^{+}89, Ras07, Wal16, WG96, Cas03, Cze08, GC97, GW57, Mac99b, Pet05, Wig67b, Wig96c, GW03a]. scientist [IM95, IM01, Neu89b]. Scientists [MP01, Seg85, Ake07, HD13, Neu90, Rob12, Str11]. scienziato [IM95]. Scott [MC00]. Scriptural [Kay97]. Sec. [vN44a]. Second [vN48d]. Secret [Nag87a]. Seek [DH55]. seen [Von87, Von91]. sein [Sze83a]. Selected [Ran82, Rec07, Lan08, Réd05, vN05, MS73, vN87]. Self [Bur61, Har67, Pes95, vNB66, Bre84a, DNvN55, Vol06, vN71b]. self-adjoint [DNvN55]. self-replication [Bre84a]. Self-Reproducing [Bur61, Har67, Pes95, vNB66, Vol06, vN71b]. Semi [vNxxb]. Semi-Relativistic [vNxxb]. seminar [Ame52, Veb37, VG55]. Seminumerical [Knu98]. semisimple [SvN50]. Senate [vN46a]. separated [CvN44a]. September [B^{+}89, BR94, Réd99, Jef51]. Sequence [vN62]. Sequences [vN25a]. Series [Jon04]. service [Die17]. session [DK04]. Set [vN28b, vN28h, vN29d, Haj59, vN25b, von97]. set-theory [Haj59]. Sets [Tau61a, KvN37]. Shannon [TWA^{+}87, ABS11]. shape [Bal04]. shapes [vN52a]. Shenker [Hen03]. Shock [BFvN^{+}44, PPvNF44, vN44c, CF48, RKH^{+}21, vNS43, vN43d, vN44b, vN45c, vN43c]. Shocks [vN43a, vNR50]. Shots [DK41]. Shoulders [FS12]. Showed [CW91, Hep90]. siglo [IM01]. Significance [WvN54]. SILLIAC [Gre10]. simple [HP20]. Simpler [BEMD11]. simulating [Dya71]. simulation [KvN97, PJW^{+}22]. Simulations [HPR14a]. since [Cer00]. sincs [Neu89a]. Single [Pos95a, Mur58]. Sitter [vN40a]. sixty [Sip05]. sixty-year-old [Sip05].

Slave [CFK^{+}91]. slippers [Kor86]. SMIL [Bru07]. SMILemu [Bru07]. Smith [vN52a]. Social [CLBdB89, Fra07, MWW77, Pin77, Leo10, Neu90]. Societal [Hei77]. Society [Rec07, Tót93, vN54e]. Solution [BMvN46, BMvN^{+}59, vN59b, vNxxb, vN41d, vNR47a, vN57a]. Solutions [BvN50, CN47, vN35h, vN46b, vN10, BvN35c, vN51d]. Solvers [JML13]. Some [Ara90, Bir77, Bre08, GvN47h, vNW28a, vNW28b, vNW28c, vN37e, vN46b, KvN37, Wol16, vN43b, vN28f, vN32a]. Son [CFK^{+}91]. Source [vNxxa, vNxxb, vH67, vN41d, vN57a]. Soviet [Gol84]. Space [Fri80, SvN46a, SvN48a, vN37e, vN39, vN40a, SvN46b, SvN48b, Réd96]. Spaces [vN35a, vN51e, JvN35, Mat03, vN35c, vN35d, vN51a]. Spacing [DK41]. Spake [CFK^{+}91]. Spanish [IM01, Pel03, vN49b, vN49c, vN91c]. spatial [CvN43]. special [DK04, RW12, vN46a]. Spectra [vNW28a, vNW28b, vNW28c, vN35b, Khr08, KvN32]. Spectral [vN51e]. Spectrum [Jon04]. Speed [vN54c, vN81a, CvN42, CvN43, vN51f]. Speeded [Ano57a]. Spektraltheorie [vN51e]. Spektren [vNW28a, vNW28b, vNW28c]. Spektrums [vN35b]. Sphere [UvN47a]. Spherical [CEE^{+}51]. Spherically [Coo57]. Spieltheorie [vNM61, vNM67a, vNM73]. spiking [PJW^{+}22]. Spin [vN28f, vNW28a, vNW28b, vNW28c]. spinning [vN44a]. Spinoza [Cso04]. Spinozas [Cso04, Cso04]. spiral [KvNS63]. spirit [Sze83c]. Splitting [vN28c]. sponsored [Ame52]. Springs [ADS83]. square [Har42, vN41b, vNKBH41, vN42b]. ssége [Neu90]. St [vN31b]. Stability [GvN47h, KvNS63]. Stan [Eck87]. Star [vNxxb, CvN44b]. Stars [vN35f, CvN42, CvN43, CvN44a, CvN44b]. Started [LR67]. Statement [vN46a]. statements [Bre99]. States [vN66, Cob72]. Static [Mar46, vN35h, KvNS63]. stationary [vN48c]. Statistical [MRvN50, RvN85, RUvN47, vNR47b, Bac22]. Statistics [Gal93, CvN42, CvN43, CvN44a, CvN44b]. Stencils [vN29h]. Stephen [Ano91]. Stern [Asp83]. Steve [Bir83, Dys81, Hod81, Kev81, Ste82]. Stewart [Whi16]. Stochastic [vN47f, CvN44b, UvN47b]. Stoltzner [Lup03, Emc02]. Stölzner [Dic03]. Stone [DK04, vN32g]. Stonebraker [DCH05]. Stored [Asp90a, CW91, HPR14c, Pel99]. Stored-Program [HPR14c, Pel99]. Storico [Ven02]. Stories [Chu14]. Story [MP01]. Strategies [vN48b]. Strategy [vN28i, Bar59, vN54d]. Strict [vN37d]. Strict- [vN37d]. structure [vN27c]. Structures [BN71]. Student [Ném84]. Studies [Phi88, SM56]. Study [Asp89a, BMvN^{+}59, Dal01, NAH^{+}89, Pin77, vNG53, Asp87, Mit80, Reg87]. Style [Bac78, Bac87, Mit22]. subject [KvN97]. subsets [vN28c]. Subsystem [Sil23]. Success [Chu14]. successive [Har42, vNK40, vN41b, vNKBH41, vN42b]. successor [KS84]. Sudarshan [Sil23]. Sum [vN48b, vN53a]. Summary [For51, Döm03, LS83]. Summer [GIS90]. Summit [CFK^{+}91]. Sun [CW91, HHW99]. Sunburst [CW91]. Super [vNF46]. Supercomputer [Ano88, Joh86]. Supersonic [CF48]. Superworkstations [SvNG89]. Supplement [Ano57e, Wei92]. Surface [UvN47a]. Surprising [MS18]. sürü [vN25a]. survey [Hal85, Her88, Her95]. survive [vN55a]. Switching [vN59a, vN54b]. Switzerland [Ano97]. Sydney [Gre10]. Symmetric [Coo57, GMvN59, vN28a, vN28d, vN46b]. Symmetries [Wig67b]. symmetrischer [vN28a, vN28d]. Symposium [Ano51, ABS96, Ber00, BH89, HFG51, RMK94, Jef51]. Synthesis [vN56b, vN63c]. System [vN28e, vN37f, vN66]. Systematic [DK41]. Systems [BMvN46, Hol62, Hon09, Pos95a, vN51c, KvN32, KvN97, RW12, Ano91]. szakvélemény [Frö93, Lad93]. szamitastechnikaban [Kov89]. számítástechnikához [Lax03]. számítógép [Gol87b, Gol03, Gol87c, Kov03b, Sze83b, Sze76]. számítógép-agyú [Sze76]. számítógép-tudományig [Sze83b]. ’számítógépek [Kov84]. számítógépet [Sze03a]. szamologep [Neu64, Neu72, Neu06]. számológépr [Érd85]. számsorozatok [vN25a]. Szanton [WS92, WS03]. Száz [Csá03, Bal04]. század [Brü87]. századon [Kőr99]. Szeged [Sza03a]. szellem [Sze83c]. Szemelvények [MS73]. Szilard [Fra05, Hor97, Lan89, Lan93b, Lan99]. születésének [Sze83b]. született [Ano93, Bal04, Csá03, Kov03b, Szé78, Zád83].

T [CEE^{+}51]. tables [Kev81]. Tabulation [Har42]. Tales [CW91]. talking [Yor87]. Tall [vN13]. tanárai [Kov03e]. target [vN41a]. társadalmi [Neu90]. Taylor [CW91, FvN51, FvN55, vN53d]. teacher [Kov03a]. teachers [Kov03e]. tech [CW91]. Technical [CFK^{+}91]. technician [Sze83a]. Technique [Grz89]. Techniques [For51, vN51b, vN63d]. technologie [CW91]. Technologies [Dys81, HB82, Hod81, Kac82, Kev81, Ste82, Hei80, Sto83]. Technology [ABS96, CW91, Kov89, Tót93, vN55a, CFK^{+}91, Ran74]. Teilmengen [vN28c]. tell [HP20]. Teller [Hor97, Lan93b, Lan99]. temperature [KvNS63]. teorija [Tur60a]. Teoriya [vN71b]. Terms [Bon04]. területén [Sze87]. Test [Bar82]. Testimony [vN57b]. Tests [Ali09]. testvérek [Fil03]. testvéremrol [Neu03]. testvéröccs [Neu89a]. text [CFK^{+}91]. Their [HHW99, IM13, vN29c, Hor97, Kőr99, Lan93a, vN37c]. Theodore [Hor97]. Theorem [Acu21, BEMD11, Bit97, DH55, Sch91, vN29e, vN32g, FH55, GL01, SvN50, WvN54, von10, Kje01, Sch91, vN37f]. Theorems [Ali09, For10, vN29b, vN32a, vNU45, CR03, Zun02]. theoretical [Luc11]. Theorie [Neu33, Tur60a, vN26b, vN27e, vN28f, vN28i, vN29f, vN29g, vN29h, vN29i, vN96d, Pou03, vN28g]. Theories [Jam74a, vN29b, vN49e]. Theorists [Eri15]. Theory [AFB79, Ano57a, Ara90, AB87, CFK^{+}91, Cso04, EMN89, Ebe67, Fri80, Hol62, Hur48, Kad58, Kal13, Kap95, Leo06, Mac92a, Mac92b, Mar65, Mur58, Neu33, Sim03, Tau61a, Tau61b, Tau63b, Tur60a, Twe93, vN27d, vN28a, vN28b, vN28h, vN28f, vN28i, vN28g, vN29a, vN29f, vN29g, vN29h, vN29d, vN29i, vN41c, vN42c, vN43d, vN44a, vNM44, vNM47, vN51e, vNM53a, vNM53b, vNM64, vNB66, vNM67b, vNM70, vNM74, vNM80, vN96d, vNM04, vNM07, Bar59, Bir58, Bub01, Del95, DJ13, Haj59, Hal58, KT58, Leo10, Mac99b, Mat03, Mol72, Nag03, PA13, Pou93, Pou03, Réd59, Sha58, Sze83b, SN57, Tar72, Val08, Wei92, WvN54, vH58, vN25b, vN26b, vN27e, vN31b, vN37a, vN48c, vN49d, vN51g]. theory [vN63b, von97, Asp90c, Har67, Tau63c, vNM61, vNM67a, vN71b, vNM73, PA13, All87, Haw45, K.44, Rel92]. there [Gol84, Kow15, CR03]. Thermodynamic [Chu21, DD13]. Thermodynamics [Sam92, vN27b]. Thermodynamik [vN27b]. Thermonuclear [Fit99, Fit13]. think [Tur60a, TvN99]. third [Ame52]. Thirty [Ame52]. Thirty-third [Ame52]. Thomas [CFK^{+}91]. Those [Kőr99]. Thought [Stö99, HP20, Hol18, Stö07]. Three [Ano57a, Frö93, Lad93]. Thus [CFK^{+}91]. Tier [Pos95a]. time [Szé03b, Ven03]. Titok [Nag87a]. több [Kov03b]. Todd [Mad12]. told [WS92, WS03]. Top [Cip00]. Topic [MCP23]. Topics [Tau62, TWA^{+}87]. Topological [vN33, vN34c, vN35c, vN35d]. topologischen [vN33, vN34c]. topology [vN36c, vN52a]. Touching [Man17]. tower [Neu89b]. trace [Mir75]. traffic [PJW^{+}22]. Tragedies [MC00, McC01]. Training [B^{+}89]. traits [Von90b]. transfert [CW91]. transfinite [KvN37, vN28h]. transfiniten [vN23]. transformation [Asp89c]. Transformationen [vN29c]. Transformations [SvN46a, SvN48a, vN29c, SvN46b, SvN48b]. Transforms [TWA^{+}87]. transinite [vN23]. Transition [vN37b, vN66, vNH81]. Transitivity [vNH37, vNH40]. Translated [vN49c, Ano91]. Translation [CFK^{+}91, Lor19]. Transmission [Cob72]. treasury [FF91]. treatment [MRvN50, vN44b]. tree [URR86c]. Tribute [Ano57a, vN51h, Tar77]. Trieste [PA13]. Triumphs [MC00, McC01]. trudy [vN87]. True [MP01, Cas03]. Trustees [Ano30]. Trying [Mir92]. tudományban [Pal87]. tudományig [Sze83b]. tudományos [GW03a, Pet05]. tudományterületeken [SI^{+}79]. tudós [Neu89b]. tudósgéniusz [Cze08]. tudósgéniusz-modell [Cze08]. tudósok [Neu90]. Tükrében [Nag87a]. tunnel [vNT45]. Turbulence [vN49e]. Turing [Ano87, CFK^{+}91, Fie15, Lie11, Ano96, Asp80, DW16, Dys12c, HP20, Her88, Her95, IM13, Lie11, Müh09, Ran72, Sch88, CK12]. Twentieth [Cor11, Bal12, Har06b, Whi16, IM01, IM09b, MHR80]. twenty [Ano87]. Two [BMvN^{+}59, FvN51, GMvN53, Pel99, Pos95a, vN48b, CvN44a, vN44c, vN53a]. Two-Dimensional [BMvN^{+}59, vN44c]. Two-Person [vN48b, vN53a]. Two-Tier [Pos95a]. Type [CN47, vN61a, vNxxb, vNR47a].

u [Cso04]. U.S. [Ake07, Bre04]. überblick [LS83]. übersetzt [vN49c]. Új [Nag87b]. Ulam [Dys12d, Eck87, JML13]. ultimate [Hor97]. unabhängiger [vN28e]. unbeschränkten [vN29h]. Uncertainty [Laf89]. undecidable [Bre99]. unfinished [Sch88]. ungarisch [Has10b, Has10a]. ungarisch-jüdischen [Has10b, Has10a]. Unified [Mar65]. Uniformly [vN25a]. Uniqueness [vN31c, vN36f, vN10, vN36g, vN51d]. unitären [vN51e]. Unitary [vN39, vN51e, SvN50]. United [Cob72]. Unity [Lea55]. Univac [Asp83, Ste81, CLBdB89]. universal [HP20, Kar04, Von90b, Her88, Her95]. universe [Dys12c, CK12]. universes [Byr10]. Universität [Has10b, Has10a]. University [CFK^{+}91, Fro00, GIS90, Gre10, Har67, Has10a, Jon04, Ran74, Twe93]. univerzális [Kar04]. Unlikely [Ber10]. Unparalleled [Vei19]. unpublished [RKH^{+}21]. Unreliable [vN56b, vN63c]. Unrestricted [vN29h]. Unsolved [Réd99, Ven02]. Untersuchung [Cso04]. Unusual [vNW29b]. upon [Kah97]. Urteile [Has10b, Has10a]. usage [PJW^{+}22]. USD [Emc02]. Use [vN45d, vN54a, Sze03a, vN63d]. Used [For51, vN51b]. Using [JML13]. út [Lax03]. Utopia [Val08]. úttör [Sze76]. üzenete [Neu90].

v [Cso04, Lie11, Sch91, Tau63b]. Validation [Cin85]. Validity [Bit97]. Value [vN48b, AFB79]. values [MRvN50]. Variable [Bub10, Jam74a]. Variables [Bel66, vN51c, GL01, Mit22, Mon79]. variance [Har42, vN41b, vN42b]. variants [GMvN53]. variates [Wal96]. variation [CvN44b]. Variational [vN45d]. Variations [vN30]. Variationsrechnung [vN30]. Varied [Lan89]. Various [For51, vN63d, vN51b]. Vassar [RMK94]. Veblen [Veb37, VG55, vN45d]. Verallgemeinerung [vN37f]. Vergleich [Lie11, Sch91]. Verhalten [vNW29a, vNM61, vNM67a, vNM73]. Verleihung [B^{+}89]. Version [Bar82, Bru07]. versus [DJ13]. verwandte [vN28h]. Very [vN13]. Vezer [Kor86, Kor86]. vezeto [Lax03]. VI [Tau63c]. via [HKLS93]. viele [vN28c]. Vienna [Lup03]. view [Cso04, vNW90]. Viewed [Kem55]. viewpoint [Roj94]. vignettes [War08]. világformáló [Har06a]. világhírig [Ném93]. vision [Luc11, Whi16]. Visionary [Bha22]. vita [IM08]. Vol [GvN47b, GvN47e, GvN47c, GvN47f, GvN47d, GvN47g, Wig96c, vNG48a, vNG48b, Rec07, vN51a]. volt [Kov99, Pal87]. Volume [BR94, Wei92, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. vom [vN31a]. von-Neumann-Rechner [Bau98a]. vonásai [Von90b]. vorticity [CFvN50]. Vorträge [B^{+}89].

W [Ano91, Ham92, Har67]. Wahrscheinlichkeitstheoretischer [vN27c]. Walk [Dys13]. walked [Hol18]. Walker [MC00]. Wallace [Bre08]. Wanted [DW16]. war [vN55b, Ake07, Bre04]. Was [CFK^{+}91, Bre04, HP20, Kov99, Pal87, Szé78, Zád83]. Watson [CFK^{+}91]. Wave [ABS96, BFH^{+}47, GvN55]. Waves [BFvN^{+}44, vN43c, CF48, PPvNF44, RKH^{+}21, vN42c, vNS43, vN43d, vN44c, vN45c, vN48c]. Way [Hep90, CW91]. Weapon [Fit99, Fit13]. weapons [Jék03a, Jék03b, Yor87]. Weather [Sma83]. Weird [Str11]. Wer [Bau98a]. Were [Mir92]. Werk [LS83, Sze83a, Bau82]. Wesley [Ano91]. West [CW91]. Western [GGC^{+}05]. Where [Lan93b, Lan99]. Who [Ano57a, Gra56, Mac92a, Mac92b, Reg87, Roj94, Sla87, SW80, Har06b, Kőr99, Mac99b, Mar06, Smi10, Ano96, Bau98a, Kov99]. Wichita [Gol84]. Widerspruchfreiheitsfrage [vN29d]. Wiener [Dys81, HB82, Hod81, Kev81, Ste82, vN49a, Cin85, Hei80, Kac82, Sto83, Bir83]. Wigner [Frö93, Hor97, Lad93, Lan93b, Lan99, Wig67b, WS92, Wig96c, WS03].

Wignerrol [Frö93, Lad93]. Wilhelm [Has10a]. Wilhelms [Has10b, Has10a]. will [Ano30]. William [All87, Kid92, Mah93, Ano91, Bow92, Pan92, Twe93]. Wilson [Stö04]. win [Bre04]. wind [vNT45]. winners [Fie15]. wirtschaftliches [vNM61, vNM67a, vNM73]. without [Ben72]. women [HHW99]. Work [Bau82, GW96, LS83, WG96, vN29i, vN32i, Asp89c, GW57, GW03b, GW03a, Haj59, KT58, Réd59, Sze83a, Sze87, Sze76, SN57, Tar58, Yee82, vN31b]. Works [Kov03c, LR67, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Ano57e, Hey47, SI^{+}79, Wig96c, vN87]. Workshop [E^{+}00, Fro00, E^{+}00]. World [BR94, Cor11, Eri15, Gra56, IM09b, MC00, McC01, Ake07, FF91, IM95, IM01, IM08, Mar06, Nat88, Ném93, CW91]. worlds [Byr10]. writings [GGC^{+}05]. wrong [Bre84b]. Wrote [Ano57a]. Wuppertal [Fro00]. www.neumann [Tós04]. www.neumann-haz.hu [Tós04].

X [Twe93, Asp83]. xi [Twe93]. xii [Ran74, vN44a]. XIII [vN44a]. xiv [Jon04]. xix [Kev81]. xvii [Bow92]. XX [IM01].

year [B^{+}89, Sip05]. yearbook [Lup03]. Years [Ber00, FS12, Lax05, Von89, Ano87, Ano93, Ano95, Ano97, Bal04, Csá03, Kov03b, Lax03, Szé78, Sze97, Von90a, Zád83]. York [Ano91, Bir97, GIS90, MC00, RMK94]. young [Frö93, Lad93].

Zahlen [vN23, vN26b, vN28e]. Zahlenfolgen [vN25a]. Zerlegung [vN28c]. Zero [vN31a, vN48b, vN53a]. Zero-Sum [vN48b, vN53a]. zeros [FvN22]. Zseni [Kov03d, Ano03, Kar04]. Zum [vN29e, vN34c]. Zur [vN31b, B^{+}89, vN28f, vN29i, vN32i, vN23, vN26b, vN27d, vN27e, vNW28a, vNW28b, vNW28c, vN28i, vN29f, vN29g, vN29h, vN32h]. zusammenfassender [LS83]. Zusatz [vN29i]. Zusätze [vN32i]. Zuse [Lie11, Lie11].

References

Aspray:1987:PJN

[AB87] William Aspray and Arthur Burks, editors. Papers of John von Neumann on Computing and Computer Theory, volume 12 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-22030-X. xviii + 624 pp. LCCN QA76.5 .P31451 1987. See [Knu70].

Amaldi:2012:ALF

[ABES12] Edoardo Amaldi, S. (Saverio) Braccini, Antonio Ereditato, and Paola Scampoli, editors. The adventurous life of Friedrich Georg Houtermans, physicist (1903–1966). Springerbriefs in Physics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32854-7 (paperback), 3-642-32855-5 (e-book). ISSN 2191-5423. xiv + 152 pp. LCCN Q124.6-127.2; QC16.H688 A43 2012. URL http://www.sif.it/

attivita/saggiatore/recensioni/e_amaldi.

Arveson:1996:QNP

[ABS96] William Arveson, Thomas Branson, and Irving Ezra Segal, editors. Quantization, nonlinear partial differential equations, and operator algebra: 1994 John von Neumann Symposium on Quantization and Nonlinear Wave Equations June 7–11 1994, Massachusetts Institute of Technology, Cambridge, Massachusetts, volume 59 of Proceedings of symposia in pure mathematics, 0082-0717. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0381-6. LCCN QC174.17.G46 J67 1994.

Anand:2011:SNE

[ABS11] Kartik Anand, Ginestra Bianconi, and Simone Severini. Shannon and von Neumann entropy of random networks with heterogeneous expected degree. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 83(3 Pt 2): 036109, March 2011. CODEN PLEEE8. ISSN 1550-2376.

Acuna:2021:NTR

[Acu21] Pablo Acuña. von Neumann’s theorem revisited. Foundations of Physics, 51(3):??, June 2021. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.springer.

com/article/10.1007/s10701-021-00474-5.

Adam:1983:JN

[Ádá83] András Ádám. John von Neumann. In Legendi and Szentivanyi [LS83], pages 11–43. ISBN 3-411-01639-6. LCCN QA76.2.V6 N414 1983g. German translation by Rozsa Nienhaus of Hungarian original, Neumann János élete és munkássága.

Aris:1983:SSC

[ADS83] Rutherford Aris, H. Ted (Howard Ted) Davis, and Roger H. Stuewer, editors. Springs of Scientific Creativity: Essays on Founders of Modern Science. University of Minnesota Press, Minneapolis, MN, USA, 1983. ISBN 0-8166-1087-8. viii + 342 pp. LCCN QC7 .S77 1983.

Abraham-Frois:1979:TVP

[AFB79] Gilbert Abraham-Frois and Edmond Berrebi. Theory of value, prices and accumulation. Cambridge University Press, Cambridge, UK, 1979. ISBN 0-521-22385-7. xvii + 260 pp. LCCN HB97.5 .A2713. A mathematical integration of Marx, von Neumann and Sraffa, A revised version translated from the French by M. P. Kregel-Jaraux.

Akera:2007:CNW

[Ake07] Atsushi Akera. Calculating a natural world: scientists, engineers, and computers during the rise of U.S. Cold War research. Inside technology. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-01231-6. ix + 427 pp. LCCN QA76.17 .A42 2007.

Alicki:2009:NBT

[Ali09] Robert Alicki. On von Neumann and Bell theorems applied to quantumness tests. Foundations of Physics, 39(4):352–360, April 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-009-9285-x.

Allison:1954:NR

[All54] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 10(10):395–397, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1987:BRBc

[All87] David K. Allison. Book review: Papers of John von Neumann on Computing and Computer Theory by John von Neumann; William Aspray; Arthur Burks. Isis, 78(4):603, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231933.

ASM:1952:MIS

[Ame52] American Society for Metals, editor. Metal interfaces; a seminar on metal interfaces held during the Thirty-third National Metal Congress and Exposition, Detroit, October 13 to 19, 1951; sponsored by the American Society for Metals. American Society for Metals, Cleveland, OH, USA, 1952. LCCN QD171 .A62.

Anonymous:1930:PTM

[Ano30] Anonymous. Princeton Trustees meet: will concentrate departments — Berlin lecturers named. New York Times, ??(??):8, January 10, 1930. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/99086947/fulltextPDF/13EDE5EACA87A3ADEAE/

1. This article appears to be the first mention of John von Neumann in this newspaper, noting his appointment, with Eugene Wigner, as lecturers in mathematical physics.

Anonymous:1951:PCA

[Ano51] Anonymous, editor. Problems of cosmical aerodynamics: Proceedings of the Symposium on the Motion of Gaseous Masses of Cosmical Dimensions held at Paris, August 16–19, 1949. Central Air Documents Office, Dayton, OH, USA, 1951. LCCN QB790 .S9 1949. Symposium on the Motion of Gaseous Masses of Cosmical Dimensions (1st: 1949: Paris, France).

Anonymous:1957:DNC

[Ano57a] Anonymous. Dr. von Neumann of A.E.C., 53, dies: Mathematician who speeded development of H-bomb; built electronic ‘brain’; received Fermi Award; architect of contributions to quantum theory; wrote on economic behavior; eminent in three fields; received Medal of Freedom; on many government agencies; tribute by President. New York Times, ??(??):19, February 9, 1957. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/114144499/13EDE5EACA87A3ADEAE/76;   http://www.
nobelprize.org/nobel_prizes/physics/laureates/1954/;
https://www.nytimes.com/1957/02/09/archives/dr-von-neumann-
of-aec-53-dies-mathematician-who-speeded-development.html.

Anonymous:1957:JN

[Ano57b] Anonymous. John von Neumann. Physics Today, 10(4):58–59, April 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/

phtoad/v10/i4/p58_s3. The same page carries an obituary for Walter Bothe (8 January 1891–8 February 1957): see entry

[Ano57a] for comments.

Anonymous:1957:NJN

[Ano57c] Anonymous. Notes: John von Neumann 1903–1957 [obituary]. Mathematical Tables and Other Aids to Computation, 11(58):127, April 1957. CODEN MTTCAS. ISSN 0891-6837. URL http://

www.jstor.org/stable/2002157.

Anonymous:1957:ONJ

[Ano57d] Anonymous. Obituary: Neumann János. Mat. Lapok, 8:1–7, 1957. ISSN 0025-519X.

Anonymous:1957:SLW

[Ano57e] Anonymous. Supplement to the list of the works of J. von Neumann. Mat. Lapok, 8:210, 1957. ISSN 0025-519X.

Anonymous:1975:PJN

[Ano75] Anonymous. Papers of John von Neumann. Historia Mathematica, 2(1):73, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900440.

Anonymous:1987:ATA

[Ano87] Anonymous. ACM Turing Award lectures: the first twenty years, 1966–1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Anonymous:1988:SJN

[Ano88] Science at the John von Neumann National Supercomputer Center: annual research report, 1988. ISSN 1046-0632. Consortium for Scientific Computing, Princeton, NJ, USA.

Anonymous:1991:HCC

[Ano91] Anonymous. History of computers and computing: Reviews of William Aspray, John von Neumann and the Origins of modern Computing, (Cambridge: MIT Press, 1990 — $35.00, ISBN 0-262-01121-2, 376 pp.); Stephen G. Nash, editor, A History of Scientific Computing (New York: Association for Computing Machinery [ACM] Press; co-published by Addison-Wesley, 1990 — price not given, ISBN 0-201-50814-1, 359 pp.); Michael Lindgren, Glory and Failure: The Difference Engines of Johann Müller, Charles Babbage and Edvard Scheutz [translated by Craig G. McKay] (Cambridge: MIT Press, 1990 — $45.00, ISBN 0-262-12146-8, 414 pp.); Emerson W. Pugh, Lyle R. Johnson, and John H. Palmer, IBM’s 360 and Early 370 Systems (Cambridge: MIT Press, 1991 — $37.50, ISBN 0-262-16123-0, 819 pp.) . Communication Booknotes Quarterly, 22(4):78, ???? 1991. ISSN 1094-8007 (print), 1532-

6896 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/10948009109487995.

Anonymous:1993:FME

[Ano93] Anonymous. A fizika módszere: 90 éve született Neumann János. (Hungarian) [The method of physics: 90 years after von Neumann’s birth]. Fizikai Szemle (Budapest), 46(12):473–475, 1993. CODEN FISZA6. ISSN ????

Anonymous:1995:RNJ

[Ano95] Anonymous. Rádióbeszélgetés Neumann Jánossal, 40 évvel ezelőtt. (Hungarian) [Radio conversation with John von Neumann, 40 years ago]. Fizikai Szemle (Budapest), 45(9):298–299, 1995. CODEN FISZA6. ISSN ????

Anonymous:1996:QIO

[Ano96] Anonymous. Qui a inventé l’ordinateur?: grands ingénieurs: Alan Turing, John Mauchly, John P. Eckert, John Atanasoff, John von Neumann. (French) [Who invented the computer? Great engineers: Alan Turing, John Mauchly, John P. Eckert, John Atanasoff, John von Neumann], volume 36 of Les Cahiers de Science et vie (Paris). Excelsior, Paris, France, 1996. ISSN 1157-4887. 96 pp. US$.

Anonymous:1997:NDB

[Ano97] Anonymous, editor. Von Neumann Day: Biological inspiration in computer science: 40 years from the death of John von Neumann Lausanne (Switzerland), July 25th, 1997. ????, ????,

1997. URL http://lslwww.epfl.ch/pages/events/neumann97/
neumann.html.

Anonymous:2003:MZN

[Ano03] Anonymous. Egy magyar zseni, Neumann János. (Hungarian) [A Hungarian genius, John von Neumann]. eVilág, 2(10):1–3, 2003. CODEN ???? ISSN ????

Anonymous:2019:JND

[Ano19] Anonymous. John von Neumann Documentary. Mathematical Association of America, ????, 2019. URL https://archive.org/

details/JohnVonNeumannY2jiQXI6nrE.

Araki:1990:SLJ

[Ara90] Huzihiro Araki. Some of the legacy of John von Neumann in physics: Theory of measurement, quantum logic, and von Neumann algebras in physics. In Glimm et al. [GIS90], pages 119–136. ISBN 0-8218-1487-7. LCCN QA29 .V66 L44 1990.

Aspray:1980:MCC

[Asp80] William F. Aspray. From mathematical constructivity to computer science: Alan Turing, John von Neumann, and the origins of computer science in mathematical logic. Thesis (Ph.D.), University of Wisconsin–Madison, Madison, WI, USA, 1980. v + 443 pp.

Aspray:1983:BRB

[Asp83] William Aspray. Book review: From Eniac to Univac: an appraisal of the Eckert–Mauchly computers: By Nancy Stern. Bedford, MA (Digital Press). 1981. x + 286 pp. Appendix: “First Draft of a Report on the EDVAC” by John von Neumann. Bibliography. Historia Mathematica, 10(2):228–233, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608390071X.

Aspray:1987:MRM

[Asp87] William Aspray. The mathematical reception of the modern computer: John von Neumann and the Institute for Advanced Study computer. In Phillips [Phi88], pages 166–194. ISBN 0-88385-128-8. LCCN ????

Aspray:1989:DJN

[Asp89a] William Aspray. Discussion: John von Neumann — a case study of scientific creativity. Annals of the History of Computing, 11 (3):165–169, July/September 1989. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1989/
pdf/a3165.pdf;   http://www.computer.org/annals/an1989/
a3165abs.htm.

Aspray:1989:JNC

[Asp89b] William Aspray. John von Neumann’s contributions to computing and computer science. Annals of the History of Computing, 11(3):189–195, July/September 1989. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1989/
pdf/a3189.pdf;   http://www.computer.org/annals/an1989/
a3189abs.htm.

Aspray:1989:TNA

[Asp89c] William Aspray. The transformation of numerical analysis by the computer: an example from the work of John von Neumann. In Rowe et al. [RMK94], pages 307–322. ISBN 0-12-599661-6 (vol. 1), 0-12-599662-4 (vol. 2), 0-12-599663-2 (vol. 3).

LCCN QA21 .S98 1989. URL http://www.loc.gov/catdir/
description/els032/89017766.html;   http://www.loc.gov/
catdir/toc/els031/89017766.html.

Aspray:1990:BBS

[Asp90a] William Aspray. Back to basics: The stored program concept. IEEE Spectrum, 27(9):51, September 1990. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aspray:1990:JNO

[Asp90b] William Aspray. John von Neumann and the Origins of Modern Computing. The MIT Press Series in the History of Computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-01121-2. xvii + 376 pp. LCCN QA76.5 .A77446 1990.

Aspray:1990:OJN

[Asp90c] William Aspray. The origins of John von Neumann’s Theory of Automata. In Glimm et al. [GIS90], pages 289–309. ISBN 0-8218-1487-7. LCCN QA29 .V66 L44 1990.

Axel:1952:NNc

[Axe52] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8 (3):94–95, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brunn:1989:AFV

[B^{+}89] Anke Brunn et al., editors. Ausbildung zur Forschung: Vorträge anläßlich des Festkolloquiums zur Einweihung des Forschungskollegs am Forschungsinstitut für Diskrete Mathematik am 4. Juli 1988 und Vorträge anläßich der Verleihung der John-von-Neumann-Professuren für das akademische Jahr 1989/90 am 28.September 1989. (German) [Training as research: Lectures on the occasion of the fixed colloquium for the inauguration of the Research College at the Research Institute for Discrete Mathematics on 4 July 1988 and lectures on the occasion of the award of the John of Neumann Professorship for the academic year 1989/90 on 28 September 1989], volume 71 of Bonner akademische Reden. Bouvier, Bonn, Germany, 1989. ISBN 3-416-09074-8. 51 pp. LCCN ????

Backus:1978:CPL

[Bac78] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. Communications of the ACM, 21(8):613–641, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL http://www.stanford.edu/class/cs242/

readings/backus.pdf. Reproduced in Selected Reprints on Dataflow and Reduction Architectures, ed. S. S. Thakkar, IEEE, 1987, pp. 215–243, and in ACM Turing Award Lectures: The First Twenty Years, ACM Press, 1987, pp. 63–130.

Backus:1987:CPL

[Bac87] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. In ACM Turing Award lectures: the first twenty years, 1966–1985 [Ano87], pages 63–130. ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987.

Bacciagaluppi:2022:SIB

[Bac22] Guido Bacciagaluppi. The statistical interpretation: Born, Heisenberg, and von Neumann, 1926–27. In Freire et al. [FBD^{+}22], page ?? ISBN 0-19-884449-2 (hardcover). LCCN

QC173.98 .O94 2022. URL https://global.oup.com/academic/
product/the-oxford-handbook-of-the-history-of-quantum-
interpretations-9780198844495.

Balogh:2004:ARS

[Bal04] Balogh Vilmos Szilárd. “Alakja rejtély”. Száz éve született Neumann János. (Hungarian) [“The shape of mystery.” One hundred years ago came John von Neumann]. Mérleg, 40(1):91–106, 2004. CODEN ???? ISSN ????

Balachandran:2012:JNM

[Bal12] V. K. Balachandran. John von Neumann—the mathematical genius of the twentieth century. Bull. Kerala Math. Assoc., 9(1): 149–164, 2012. ISSN 0973-2721.

Bargmann:1959:TGS

[Bar59] S. Bargmann. On the theory of games of strategy. Annals of Mathematics Studies, 40(??):13–42, ???? 1959. CODEN ANMAAH. ISSN 0066-2313. English translation of [vN28i].

Bartels:1982:RVN

[Bar82] Robert Bartels. The rank version of von Neumann’s ratio test for randomness. Journal of the American Statistical Association, 77(377):40–46, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287767.

Baumgartel:1982:JNL

[Bau82] Hellmut Baumgärtel. John von Neumann, aus Leben und Werk. (German) [John von Neumann, from life and work]. Mitteilungen der Mathematischen Gesellschaft der Deutschen Demokratischen Republik, 3–4:87–98, 1982. CODEN ???? ISSN ????

Bauer:1998:WEN

[Bau98a] Friedrich L. Bauer. Wer erfand den von-Neumann-Rechner?. (German) [Who invented the von Neumann calculator?]. Informatik Spektrum, 21(2):??, ???? 1998. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://link.springer.

com/article/10.1007/s002870050091.

Baumgartel:1998:ESJ

[Bau98b] Hellmut Baumgärtel. Erhard Schmidt, John von Neumann. In Mathematics in Berlin, pages 97–104. Birkhäuser, Berlin, 1998.

Becker:2004:NDB

[Bec04] Lon Becker. That von Neumann did not believe in a physical collapse. British Journal for the Philosophy of Science, 55(1): 121–135, March 2004. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.org/
content/55/1/121.full.pdf+html;   http://www.jstor.org/
stable/3541837.

Bell:1966:PHV

[Bel66] John S. Bell. On the problem of hidden variables in quantum mechanics. Reviews of Modern Physics, 38(3):447–452, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.38.447;   http://rmp.aps.org/abstract/
RMP/v38/i3/p447_1.

Ben-El-Mechaiekh:2011:SPN

[BEMD11] Hichem Ben-El-Mechaiekh and Robert W. Dimand. A simpler proof of the von Neumann minimax theorem. American Mathematical Monthly, 118(7):636–641, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
07.636.pdf.

Benioff:1972:PQM

[Ben72] P. A. Benioff. Procedures in quantum mechanics without von Neumann’s projection axiom. Journal of Mathematical Physics, 13(9): 1347–1355, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bencze:2003:NJK

[Ben03] Bencze Gyula. Neumann János és a kvantummechanika megalapozása. (Hungarian) [John von Neumann and foundations of quantum mechanics]. Fizikai Szemle (Budapest), 53(12):420–422, 2003. CO-DEN FISZA6. ISSN ????

Benov:2016:MPF

[Ben16] Dobriyan M. Benov. The Manhattan Project, the first electronic computer and the Monte Carlo method. Monte Carlo Methods and Applications, 22(1):73–??, March 2016. CO-DEN MCMAC6. ISSN 0929-9629 (print), 1569-3961 (elec-

tronic). URL https://www.degruyter.com/view/j/mcma.2016.
22.issue-1/mcma-2016-0102/mcma-2016-0102.xml.

Bergmann:1969:BRJ

[Ber69] Peter G. Bergmann. Book review: Johann von Neumann, Mathematische Grundlagen der Quantenmechanik and O. Costa de Beauregard, Précis de Mécanique Quantique Relativiste. Physics Today, 22(5):85, 87, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3035594.

Bergin:2000:YAC

[Ber00] Thomas J. Bergin, editor. 50 Years of Army Computing: From ENIAC to MSRC: A Record of a Symposium and Celebration November 13 and 14, 1996 Aberdeen Proving Ground, volume ARL-SR-93. U.S. Army Research Laboratory, Aberdeen, MD, USA, September 2000. ISBN 0-9702316-1-X. LCCN QA76.5 .S95

1996. URL https://permanent.fdlp.gov/lps58495/lps58495.
pdf.

Bernstein:2010:JNK

[Ber10] Jeremy Bernstein. John von Neumann and Klaus Fuchs: an unlikely collaboration. Physics in Perspective (PIP), 12(1):36–50, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0001-1;   http://www.springerlink.com/
content/y54804274h88v546/.

Bethe:1947:BW

[BFH^{+}47] Hans A. Bethe, Klaus Fuchs, Joseph O. Hirschfelder, John L. Magee, Rudolf E. Peierls, and John von Neumann. Blast wave. U.S. Government Document LA-2000 (TID-4500, 13th edition, supplement), Los Alamos Scientific Laboratory, Los Alamos, NM, USA,

August 1947. 301 pp. URL https://fas.org/sgp/othergov/

doe/lanl/docs1/00320773.pdf. First distributed 27 March 1958.

Bethe:1944:LSH

[BFvN^{+}44] Hans A. Bethe, Klaus Fuchs, John von Neumann, Rudolf Peierls, and W. G. Penney. Lectures in Shock Hydrodynamics and Blast Waves. ????, October 28, 1944. Notes by J. O. Hirschfelder.

Burks:1946:PDL

[BGvN46] Arthur W. Burks, Herman H. Goldstine, and John von Neumann. Preliminary discussion of the logical design of an electronic computing instrument. Technical report, Institute for Advanced Study, Princeton, NJ, USA, June 28, 1946. 42 pp. URL https:

//deepblue.lib.umich.edu/handle/2027.42/3972;   https:/
/grch.com.ar/docs/p1/Apuntes/eng/Logical%20Design%20of%
20an%20Electronic%20Computing%20Instrument.pdf. Report to

the U.S. Army Ordnance Department under contract W-36-034-OKD-7481. Reprinted in [Tau63b, Paper 2], [BN71], [Swa76, pp. 221–259] and [AB87, pp. 97–146].

Burks:1982:PDL

[BGvN82] Arthur W. Burks, Herman H. Goldstine, and John von Neumann. Preliminary discussion of the logical design of an electronic computing instrument (1946). In Randell [Ran82], pages 399–413. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

Brink:1989:PCB

[BH89] Jean R. Brink and C. Roland Haden. Prologue: The computer and the brain: An international symposium in commemoration of John von Neumann (1903–1957). Annals of the History of Computing, 11(3):161–163, July/September 1989. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1989/
pdf/a3161.pdf;   http://www.computer.org/annals/an1989/
a3161abs.htm.

Bhattacharya:2022:MFV

[Bha22] Ananyo Bhattacharya. Man from the Future: the Visionary Life of John von Neumann. W. W. Norton & Co., New York, NY, USA, 2022. ISBN 1-324-00399-5. xiv + 353 pp. LCCN QA29.V66 B43 2022.

Birkhoff:1958:NLT

[Bir58] Garrett Birkhoff. Von Neumann and lattice theory. Bulletin of the American Mathematical Society, 64(3 (part 2)):50–56, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X

(electronic). URL http://projecteuclid.org/euclid.bams/
1183522370.

Birkhoff:1977:SLA

[Bir77] Garrett Birkhoff. Some leaders in American mathematics. In Tarwater [Tar77], pages 25–78. URL https://maa.org/

sites/default/files/pdf/ABOUTMAA/History/Bicentennial_
Tribute.pdf.

Birkhoff:1983:BRB

[Bir83] Garrett Birkhoff. Book review: John von Neumann and Norbert Wiener: By Steve Heims. Cambridge, Mass., and London (M.I.T. Press). 1981. Historia Mathematica, 10(2):243–248, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900745.

Birkhoff:1997:BRB

[Bir97] Garrett Birkhoff. Book review: John von Neumann. By Norman Macrae. New York (Random House). 1992. Historia Mathematica, 24(1):98–100, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086097921376.

Bitsakis:1997:VNT

[Bit97] Eftichios Bitsakis. On the validity of von Neumann’s theorem. In Ginev and Cohen [GC97], pages 47–62. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346. LCCN Q174 .B67

vol. 192. URL https://link.springer.com/book/10.1007/978-
94-011-5788-9.

Blair:1957:PGM

[Bla57] Clay Blair, Jr. Passing of a great mind. Fortune, ??(??):89, February 25, 1957. CODEN FORTAP. ISSN 0015-8259.

Bargmann:1946:SLS

[BMvN46] V. Bargmann, D. Montgomery, and J. von Neumann. Solution[s] of linear systems of high order. Report, Navy Bureau of Ordinance, Washington, DC, USA, 1946. 85 pp. Report prepared for Navy Bureau of Ordnance under contract Nord-9596. Reprinted in [Tau62, vol. 4, pp. 421–478] and [Tau63b, Paper 13].

Blair:1959:SNS

[BMvN^{+}59] A. Blair, N. Metropolis, J. von Neumann, A. H. Taub, and M. Tsingou. A study of a numerical solution to a two-dimensional hydrodynamical problem. Mathematical Tables and Other Aids to Computation, 13(67):145–184, July 1959. CODEN MTTCAS. ISSN 0891-6837. Condensation of Los Alamos Scientific Laboratory Report LA-2165. Reprinted in [Tau63b, Paper 17].

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell, editors. Computer Structures: Readings and Examples. McGraw-Hill, New York, 1971. ISBN 0-07-004357-4. xix + 668 pp. LCCN TK7888.3 .B4.

Bochner:1958:JN

[Boc58] Salomon Bochner. John von Neumann: 1903–1957. Biographical memoirs — National Academy of Sciences of the United States of America, pages 436–457, 1958. CODEN BMNSAC.

ISSN 0077-2933. URL http://books.nap.edu/html/biomems/
jvonneumann.pdf.

Bonanno:2004:CNG

[Bon04] Giacomo Bonanno. A characterization of von Neumann games in terms of memory. Synthese, 139(2):281–295, March 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1023/B%3ASYNT.0000024905.25386.3d.pdf.

Bourzac:2018:FAM

[Bou18] Katherine Bourzac. For faster AI, mix memory and processing [news]. IEEE Spectrum, 55(4):12–13, April 2018. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bowker:1992:BRW

[Bow92] Geof Bowker. Book review: William Aspray. John von Neumann and the Origins of Modern Computing. Cambridge, Mass, and London: MIT Press, 1991. Pp. xvii + 376. ISBN 0-262-01121-2. £31.50. British Journal for the History of Science, 25(3):386–387, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027288.

Brunnstein:1994:IPP

[BR94] Klaus Brunnstein and Eckart Raubold, editors. Information processing ’94: proceedings of the IFIP World Computer Congress, Hamburg, Germany, 28 August–2 September, 1994. Volume 2: Applications and Impacts, volume A52–A53 of IFIP transactions. A, Computer science and technology. North-Holland, Amsterdam, The Netherlands, 1994. ISBN 0-444-81990-8 (set), 0-444-81987-8 (vol. 2), 0-444-81988-6 (vol. 3). ISSN 0926-5473. LCCN QA75.5 .I38 1994.

Brenner:1984:JNH

[Bre84a] Sydney Brenner. John von Neumann and the history of DNA and self-replication. Web interview (7m11s), 1984. URL https:

//www.webofstories.com/play/13258?o=MS;   https://www.
webofstories.com/play/sydney.brenner/45.

Brenner:1984:SWN

[Bre84b] Sydney Brenner. Schrödinger wrong, von Neumann right. Web

interview (2m5s), 1984. URL https://www.webofstories.com/
play/13259?o=MS;   https://www.webofstories.com/play/sydney.
brenner/46.

Breuer:1999:JNM

[Bre99] Thomas Breuer. John von Neumann met Kurt Gödel: undecidable statements in quantum mechanics. In Language, quantum, music (Florence, 1995), volume 281 of Synthese Lib., pages 159–

170. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999.

Brewster:2004:JNM

[Bre04] Mike Brewster. John von Neumann: MANIAC’s father. Inventing the computer that helped the U.S. win the Cold War was only one of this mathematician’s many accomplishments. BusinessWeek online, April 8, 2004. URL http://www.businessweek.

com/bwdaily/dnflash/apr2004/nf2004048_9158_db078.htm.

Brent:2008:SCC

[Bre08] Richard P. Brent. Some comments on C. S. Wallace’s random number generators. The Computer Journal, 51(5):579–584, February 2008. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). See [Wal96].

Brink:1989:JNR

[Bri89] Jean R. Brink. John von Neumann reconsidered. Annals of the History of Computing, 11(3):179–181, July/September 1989. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1989/pdf/a3179.pdf;   http://www.computer.org/
annals/an1989/a3179abs.htm.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:BCJ

[Bro73b] Jacob Bronowski. The brain and the computer: John von Neumann. In The Ascent of Man [Bro73a], pages 432–436. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Bruckner:1987:JSP

[Brü87] Brückner Huba. A jöv század polihisztora. (Hungarian) [The next century’s polymath]. Magyar nemzet, 50(45):7, 1987. CODEN ???? ISSN ????

Brunschen:2007:SSE

[Bru07] Christian Brunschen. SMILemu: The SMIL emulator: Version 1.2. Web site, August 30, 2007. URL http://en.wikipedia.

org/wiki/IAS_computer;   http://en.wikipedia.org/wiki/
SMIL_%28computer%29;   http://video.ldc.lu.se/smil-50.
htm; http://www.smilemu.org/. SMIL (Siffermaskinen i Lund

= Number Machine in Lund) was an early Swedish computer introduced in 1956, and in operation until 1970, that was based on John von Neumann’s 1952 IAS machine that was designed from 1945 to 1951, and was operational until 1958.

Bromberg:2004:RJN

[BS04] Shirley Bromberg and Patricia Saavedra. Remembering John von Neumann. Miscelánea Mat., 40(??):95–114, 2004.

Bub:2001:NTQ

[Bub01] Jeffrey Bub. Von Neumann’s theory of quantum measurement. In Rédei and Stöltzner [RS01b], pages 63–74. ISBN 0-7923-6812-6. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN ????

Bub:2010:NNH

[Bub10] Jeffrey Bub. Von Neumann’s ‘no hidden variables’ proof: a reappraisal. Foundations of Physics, 40(9–10):1333–1340, October 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-010-9480-9.

Burks:1961:NJN

[Bur61] Arthur W. Burks. Notes on John von Neumann’s cellular selfreproducing automaton. Technical Report 108, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, Illinois, 1961. 58 pp.

Burks:1970:ECA

[Bur70] Arthur W. Burks. Essays on Cellular Automata. University of Illinois Press, Urbana, IL, USA, 1970. xxvi + 375 pp. LCCN QA267.5.S4 B87. A classic collection of papers on cellular automata intended as a companion to von Neumann’s Theory of Self Reproducing Automata [vNB66]. Includes papers by Burks, Thatcher, Moore, Myhill, Ulam, and Holland. See [Hol62].

Bochner:1935:APFa

[BvN35a] S. Bochner and J. von Neumann. Almost periodic functions in groups. II. Bulletin of the American Mathematical Society, 41(??): 35–??, ???? 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Bochner:1935:APFb

[BvN35b] S. Bochner and J. von Neumann. Almost periodic functions in groups. II. Transactions of the American Mathematical Society, 37 (1):21–50, ???? 1935. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). Reprinted in [Tau61b, Paper 26]. This twopart paper won the American Mathematical Society Bôcher Memorial Prize in 1938.

Bochner:1935:CSO

[BvN35c] S. Bochner and J. von Neumann. On compact solutions of operational-differential equations. I. Annals of Mathematics (2), 36(1):255–291, ???? 1935. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 1].

Birkhoff:1936:LQM

[BvN36] Garrett Birkhoff and John von Neumann. The logic of quantum mechanics. Annals of Mathematics, 37(??):823–843, ???? 1936. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 7].

Brown:1950:SGD

[BvN50] G. W. Brown and J. von Neumann. Solutions of games by differential equations. In Contributions to the Theory of Games, volume 24 of Annals of Mathematics Studies, pages 73–79. Princeton University Press, Princeton, NJ, 1950. Reprinted in [Tau63c, Paper 4].

Byrne:2010:MWH

[Byr10] Peter Byrne. The many worlds of Hugh Everett III: multiple universes, mutual assured destruction, and the meltdown of a nuclear family. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-955227-4 (hardcover), 0-19-965924-9 (paperback). xiii + 436 pp. LCCN QC774.E94 .B97

2012. URL https://global.oup.com/academic/product/the-
many-worlds-of-hugh-everett-iii-9780199552276.

Caruana:1995:JNI

[Car95] Louis Caruana. John von Neumann’s “impossibility proof” in a historical perspective. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 32(1):109–124, 1995. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Casti:2003:OTP

[Cas03] J. L. Casti. The one true platonic heaven: a scientific fiction on the limits of knowledge. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08547-0 (hardcover), 0-309-09510-7 (paperback). xviii + 160 pp. LCCN Q175 .C4339 2003. URL http://www.loc.

gov/catdir/toc/fy037/2003002279.html;   http://www.nap.
edu/catalog.php?record_id=10533.

Chambers:1984:DC

[CDJ84] Fred B. Chambers, David A. Duce, and Gillian P. Jones, editors. Distributed Computing, volume 20 of APIC studies in data processing. Academic Press, New York, NY, USA, 1984. ISBN 0-12-167350-2. xii + 327 pp. LCCN QA76.9.D5 D49 1984.

Calkin:1951:BDM

[CEE^{+}51] John Calkin, Cerda Evans, Foster Evans, John von Neumann, and Klari von Neumann. The burning of D–T mixtures in a spherical geometry. Classified report LA-1233, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April 23, 1951. 176 pp.

Ceruzzi:2000:NNS

[Cer00] Paul Ceruzzi. Nothing new since von Neumann: a historian looks at computer architecture, 1945–1995. In Rojas and Hashagen [RH00], pages 195–217. ISBN 0-262-18197-5 (hardcover), 0-585-35535-5 (electronic). LCCN QA76.17 .F57 2000.

Courant:1948:SFS

[CF48] Richard Courant and Kurt O. (Kurt Otto) Friedrichs, editors. Supersonic flow and shock waves, volume 1 of Pure and Applied Mathematics. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1948. xvi + 464 pp. LCCN ???? Originates from a report issued in 1944 under the auspices of the Office of Scientific Research and Development.

Ceruzzi:1991:RCK

[CFK^{+}91] Paul Ceruzzi, Kenneth Flamm, Peggy Aldrich Kidwell, Herbert R. J. Grosch, and John A. N. Lee. Reviews: Campbell-Kelly: ICL: A Business and Technical History; Aspray: Computing Before Computers; Watson and Petre: Father, Son & Co.; Asimov and Frenkel: Robots: Machines in Man’s Image; McNeil: An Encyclopedia of the History of Technology; Byte: Fifteenth Anniversary Summit; Deavours and Kruh: The Turing Bombe: Was it Enough?; Pearcey: A History of Australian Computing; Aspray: The Origins of John von Neumann’s Theory of Automata; Crossley and Henry: Thus Spake al-Khwarizmi: a Translation of the text of Cambridge University Library Ms. li.vi.5; Fauvel and Gerdes: African Slave and Calculating Prodigy: Bicentenary of the Death of Thomas Fuller; Marling: Maestro of Many Keyboards [brief biography of Donald Knuth]. Annals of the History of Computing, 13(1):111–117, January/March 1991. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1991/
pdf/a1111.pdf;   http://www.computer.org/annals/an1991/
a1111abs.htm.

Charney:1950:NIB

[CFvN50] J. G. Charney, R. Fjörtoft, and J. von Neumann. Numerical integration of the barotropic vorticity equation. Tellus, 2:237–254, 1950. ISSN 0040-2826. Reprinted in [Tau63c, Paper 30].

Chung:2006:RSS

[Chu14] Deborah D. L. Chung, editor. The Road to Scientific Success: Inspiring Life Stories of Prominent Researchers, volume 1. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006–2014. ISBN 981-256-600-7 (hardcover), 981-256-466-7 (paperback). xi + 230 pp. LCCN Q141 .R518 2006.

Chua:2021:DNE

[Chu21] Eugene Y. S. Chua. Does von Neumann entropy correspond to thermodynamic entropy? Philosophy of Science, 88(1):145–168, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cini:1985:CDC

[Cin85] Marcello Cini. The context of discovery and the context of validation: The proposals of von Neumann and Wiener in the development of 20th-Century physics. Rivista di Storia della Scienza, 2 (??):99–122, ???? 1985. CODEN ???? ISSN 1122-0228.

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437.

URL https://archive.siam.org/pdf/news/637.pdf.

Campbell-Kelly:2012:NCR

[CK12] Martin Campbell-Kelly. Von Neumann’s computer: [review of Turing’s Cathedral: the Origins of the Digital Universe, George Dyson, 2012 Pantheon Books £25.00 / $29.95 hardcover 423pp]. Physics World, 25(12):44–45, December 2012. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://iopscience.iop.org/pwa/full/pwa-pdf/
25/12/phwv25i12a42.pdf.

Campbell-Kelly:2013:RMM

[CK13] Martin Campbell-Kelly. Remembering Michael S. Mahoney. Perspectives on Science: Historical, Philosophical, Social, 21(3):379– 383, Fall 2013. CODEN PRSIEU. ISSN 1063-6145 (print),

1530-9274 (electronic). URL http://muse.jhu.edu/journals/
perspectives_on_science/v021/21.3.campbell-kelly.html;
 http://muse.jhu.edu/journals/perspectives_on_science/
v021/21.3.campbell-kelly.pdf.

Ceruzzi:1989:RNB

[CLBdB89] Paul Ceruzzi, John A. N. Lee, Susan A. Bradley, and Donald deB. Beaver. Reviews: a new beginning; Reid: Computers and Chips; Howell: John von Neumann, Inventors and Discoverers; Lundstrom: A Few Good Men from Univac; Perrolle: Computers and Social Change: Information, Property, and Power. Annals of the History of Computing, 11(4):329–338, October/ December 1989. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1989/pdf/a4329.pdf;
 http://www.computer.org/annals/an1989/a4329abs.htm.

Crank:1947:PMN

[CN47] John Crank and Phyllis Nicolson. A practical method for numerical evaluation of solutions of partial differential equations of the heatconduction type. Mathematical proceedings of the Cambridge Philosophical Society, 43(1):50–67, January 1947. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic). URL https://en.

wikipedia.org/wiki/Crank%E2%80%93Nicolson_method.

Coben:1972:SET

[Cob72] Stanley Coben. The scientific establishment and the transmission of quantum mechanics to the United States, 1919–32. The American Historical Review, 76(2):442–466, April 1972. ISSN 0002-8762 (print), 1937-5239 (electronic). URL http://www.jstor.

org/stable/1858707.

Cook:1957:RFP

[Coo57] J. M. Cook. Rational formulae for the production of a spherically symmetric probability distribution. Mathematical Tables and Other Aids to Computation, 11(58):81–82, April 1957. CO-DEN MTTCAS. ISSN 0891-6837. URL http://www.jstor.org/

stable/2002156.

Corry:2011:BRB

[Cor11] Leo Corry. Book review: The World as a Mathematical Game: John von Neumann and Twentieth Century Science. Isis, 102(1): 186–187, March 2011. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/10.
1086/660247.

Csaba:2002:TNM

[CR03] Ferenc Csaba and Miklós Rédei. “There is no more reason to reject intuitionism.” John von Neumann on Gödel’s incompleteness theorems and on the nature of mathematics. Mat. Lapok (N.S.), 11(2):16–25 (2006), 2002/03. ISSN 0025-519X.

Craig:1998:REP

[Cra98] Edward Craig, editor. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN

0-415-16916-X. LCCN B51. URL http://www.rep.routledge.
com/.

Csaszar:2003:SES

[Csá03] Császár Ákos. Száz éve született Neumann János. (Hungarian) [One hundred years ago came John von Neumann]. Középiskolai matematikai és fizikai lapok, 53(9):519–520, 2003. CODEN ???? ISSN ????

Csobadi:2004:SML

[Cso04] Susanne Csobádi. Spinozas mechanischer Lebensbegriff: Untersuchung e. Analogie d. Artificial Life Theory v. John von Neumann u. Spinozas Naturphilosophie. (German) [Spinoza’s mechanical view of life: Investigation of an analogy of artificial life theory by John von Neumann and Spinoza’s natural philosophy]. Magisterarbeit, Universität München, München, Germany, 2004. 100 pp.

Chandrasekhar:1942:SGF

[CvN42] Subramanyan Chandrasekhar and John von Neumann. The statistics of the gravitational field arising from a random distribution of stars. I. The speed of fluctuations. Astrophysical Journal, 95:489– 531, 1942. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). Reprinted in [Tau63c, Paper 12].

Chandrasekhar:1943:SGF

[CvN43] Subramanyan Chandrasekhar and John von Neumann. The statistics of the gravitational field arising from a random distribution of stars. II. The speed of fluctuations; dynamical friction; spatial correlations. Astrophysical Journal, 97:1–27, 1943. CODEN ASJOAB.

ISSN 0004-637X (print), 1538-4357 (electronic). Reprinted in [Tau63c, Paper 13].

Chandrasekhar:1944:SGFa

[CvN44a] Subramanyan Chandrasekhar and John von Neumann. The statistics of the gravitational field arising from a random distribution of stars. III. The correlations in the forces acting at two points separated by a finite distance. Astrophysical Journal, 99:25–46, 1944. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Chandrasekhar:1944:SGFb

[CvN44b] Subramanyan Chandrasekhar and John von Neumann. The statistics of the gravitational field arising from a random distribution of stars. IV. The stochastic variation of the force acting on a star. Astrophysical Journal, 97:47–53, 1944. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Ceruzzi:1991:RPI

[CW91] Paul Ceruzzi and Eric A. Weiss. Reviews: Penzias: Ideas and Information: Managing in a High-tech World; Rose: West of Eden: The End of Innocence at Apple Computer; Lindgren: Glory & Failure: the Difference Engines of Johann Müller, Charles Babbage, and Edvard Scheutz; Jennings: The Devouring Fungus: Tales of the Computer Age; Norberg: High Technology Calculation in the Early 20th Century: Punched Card Machinery in Business and Government; Hall and Barry: Sunburst: The Ascent of Sun Microsystems; Heppenheimer: How von Neumann Showed the Way; Aspray: Back to Basics: The Stored Program Concept; Rosen: The Origins of Modern Computing; Taylor: In Memoriam: J. C. R. Licklider, 1915–1990; Mounier-Kuhn: Genese de l’informatique en France (1945–1965): Diffusion de l’innovation et transfert de technologie; Eisler: My Life with the Printed Circuit. Annals of the History of Computing, 13(2):231–234, April/June 1991. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/books/
an1991/pdf/a2231.pdf;   http://www.computer.org/annals/
an1991/a2231abs.htm.

Czeizel:2008:NJM

[Cze08] Czeizel Endre. Neumann János (1903–1957), mint az egyik tudósgéniusz-modell megtestesítője. (Hungarian) [John von Neumann (1903–1957) embodies a model of a scientific genius]. Psychiatria Hungarica, 23(2):72–84, 2008. CODEN ???? ISSN ????

Dalmedico:2001:HEM

[Dal01] Amy Dahan Dalmédico. History and epistemology of models: Meteorology (1946–1963) as a case study. Archive for History of Exact Sciences, 55(5):395–422, April 2001. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=5&spage=395.

Dantzig:1985:ILP

[Dan85] George B. Dantzig. Impact of linear programming on computer development. Technical Report SOL 85-7, Systems Optimization Laboratory, Stanford University, Stanford, CA, USA, June

1985. 20 pp. URL https://apps.dtic.mil/dtic/tr/fulltext/
u2/a157659.pdf.

Danilov:1990:DFN

[Dan90] Yu. A. Danilov. Dzhon fon Neiman, volume 90 of Novoe v Zhizni, Nauke, Tekhnike: Seriya “Matematika, Kibernetika” [Current Life, Science and Technology: Series “Mathematics and Cybernetics”]. “Znanie”, Moscow, 1990. ISBN 5-07-001585-0. 47 pp. With a preface by I. M. Yaglom.

Dore:1989:JNM

[DCG89] Mohammed H. I. Dore, Sukhamoy Chakravarty, and Richard M. (Richard Murphey) Goodwin, editors. John von Neumann and modern economics. Clarendon Press, Oxford, UK, 1989. ISBN 0-19-828554-X. xvi + 241 pp. LCCN HB102.V65 J64 1989. US$27.50.

DeWitt:2005:SRI

[DCH05] David DeWitt, Michael Carey, and Joseph M. Hellerstein. Stonebraker receives IEEE John von Neumann Medal. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):13– 15, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deville:2013:CLB

[DD13] Alain Deville and Yannick Deville. Clarifying the link between von Neumann and thermodynamic entropies. European Physical Journal H, 38(1):57–81, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://link.

springer.com/article/10.1140/epjh/e2012-30032-0.

DellAglio:1995:DHM

[Del95] Luca Dell’Aglio. Divergences in the history of mathematics: Borel, von Neumann and the genesis of game theory. Rivista di Storia della Scienza. Serie 2. Università degli Studi di Roma “La Sapienza”, 3(2):1–45 (1997), 1995. ISSN 1122-0228.

Devroye:2017:EBC

[DG17] Luc Devroye and Claude Gravel. The expected bit complexity of the von Neumann rejection algorithm. Statistics and Computing, 27(3):699–710, May 2017. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11222-016-9648-z.

Dulmage:1955:TFK

[DH55] L. Dulmage and I. Halperin. On a theorem of Frobenius–König and J. von Neumann’s game of hide and seek. Transactions of the Royal Society of Canada, Section 3, Third Series, 49:23–25, 1955.

Diaz:2005:JNP

[Día05] J. I. Díaz. John von Neumann: from pure mathematics to applied mathematics. Boletín de la Sociedad Española de Matemática Aplicada. SēMA, 32(??):149–169, 2005. ISSN 1575-9822.

Dickson:2003:BRM

[Dic03] Michael Dickson. Book review: Miklós Rédei and Michael Stölzner, (eds.), John von Neumann and the Foundations of Physics. Philosophy of Science, 70(4):855–859, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/378876.

Dieks:2017:NIP

[Die17] Dennis Dieks. Von Neumann’s impossibility proof: Mathematics in the service of rhetorics. Studies in History and Philosophy of Modern Physics, 60(??):136–148, November 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301307.

Duncan:2013:NMY

[DJ13] A. Duncan and M. Janssen. (Never) mind your p’s and q’s: von Neumann versus Jordan on the foundations of quantum theory. European Physical Journal H, 38(2):175–259, March 2013.

CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2012-30024-5.

Dederick:1941:OSB

[DK41] L. S. Dederick and R. H. Kent. Optimum spacing of bombs of shots in the presence of systematic errors. Report 241, Ballistic Research Laboratory, Aberdeen, MD, USA, July 3, 1941.

Doran:2004:OAQ

[DK04] Robert S. Doran and Richard V. Kadison, editors. Operator algebras, quantization, and non-commutative geometry: a centennial celebration honoring John von Neumann and Marshall H. Stone: AMS special session on operator algebras, quantization, and noncommutative geometry, a centennial celebration honoring John von Neumann and Marshall H. Stone, January 15–16, 2003, Baltimore, Maryland, volume 365 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-3402-9 (print), 0-8218-7955-3 (electronic). ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA612.33 .O64 2004. URL http:/

/www.ams.org/books/conm/365/.

Devinatz:1955:PSA

[DNvN55] A. Devinatz, A. E. Nussbaum, and J. von Neumann. On the permutability of self-adjoint operators. Annals of Mathematics (2), 62:199–203, 1955. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 28].

Domolki:2003:OAN

[Döm03] Dömölki Bálint. Összefoglaló áttekintés Neumann János életútjáról. (Hungarian) [Executive summary: John von Neumann’s path of life]. Fizikai Szemle (Budapest), 53(12):416–418, 2003. CODEN FISZA6. ISSN ????

DeBenedictis:2016:HWM

[DW16] Erik P. DeBenedictis and R. Stanley Williams. Help wanted: A modern-day Turing. Computer, 49(10):76–79, October 2016. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://www.computer.org/csdl/mags/co/2016/
10/mco2016100076-abs.html.

Dyadkin:1971:ANA

[Dya71] I. G. Dyadkin. An analogue of von Neumann’s algorithm for simulating a normal distribution. In Monte Carlo methods and applications (Metody Monte-Karlo i ikh primeneniya USSR Academy of Sciences, Siberian Branch. Computing Center (VTs SO AN SSSR). Novosibirsk, 30 August–3 September 1971 (Vychislitelnyi Tsentr, Sibirskoye Otdelenie. Akademiia Nauk. SSSR)), pages 53– 54. Akademiia Nauk SSSR, Novosibirsk, USSR, 1971.

Dyson:1981:BRBa

[Dys81] Freeman J. Dyson. Book review: John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death, by Steve J. Heims, MIT Press. Technology Review (M.I.T.), 83(4): 16–??, February/March 1981. CODEN TEREAU. ISSN 0040-1692.

Dyson:2012:MC

[Dys12a] George Dyson. Monte Carlo. In Turing’s cathedral: the origins of the digital universe [Dys12c], chapter 10, pages 175–199. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012. Pages 191– 192 describe the origin of the Monte Carlo method.

Dyson:2012:NJ

[Dys12b] George Dyson. Neumann János. In Turing’s cathedral: the origins of the digital universe [Dys12c], chapter 4, pages 40–63. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012.

Dyson:2012:TCO

[Dys12c] George Dyson. Turing’s cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Dyson:2012:UD

[Dys12d] George Dyson. Ulam’s demons. In Turing’s cathedral: the origins of the digital universe [Dys12c], chapter 11, pages 200–224. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012. Pages 191– 192 describe the origin of the Monte Carlo method.

Dyson:2013:WTJ

[Dys13] Freeman Dyson. A walk through Johnny von Neumann’s garden. Notices of the American Mathematical Society, 60(2):154–161, February 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/201302/

rnoti-p154.pdf. Talk given at Brown University, Providence, Rhode Island, May 4, 2010.

Esser:2000:WMD

[E^{+}00] Rüdiger Esser et al., editors. Workshop on Molecular Dynamics on Parallel Computers (1999: John von Neumann Institute for Computing) Workshop on Molecular Dynamics on Parallel Computers: John von Neumann Institute for Computing (NIC) Research Centre, Jülich, Germany, 8–10 February 1999. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-4232-8. LCCN QP517.M65 W67 1999.

Eberl:1967:BRJ

[Ebe67] W. Eberl. Book report: J. von Neumann und O. Morgenstern, Theory of Games and Economic Behavior. Computing, 2(2):175, 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckhardt:1987:SUJ

[Eck87] Roger Eckhardt. Stan Ulam, John von Neumann, and the Monte Carlo method. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):131–137, 1987. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
00326867.pdf;   http://library.lanl.gov/cgi-bin/getfile?

15-12.pdf. With contributions by Tony Warnock, Gary D. Doolen and John Hendricks, Stanisław Ulam 1909–1984.

Emch:1997:FQM

[Emc97] Gérard G. Emch. Foundations of quantum mechanics: Building on von Neumann’s heritage. International Journal of Quantum Chemistry, 65(5):379–387, December 5, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http:/

/www3.interscience.wiley.com/cgi-bin/abstract?ID=42784;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?ID=

42784&PLACEBO=IE.pdf. Special Issue: Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods. Issue Edited by Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, Michael C. Zerner.

Emch:2002:BRB

[Emc02] Gérard G. Emch. Book review: John von Neumann and the Foundations of Quantum Physics. Edited by Miklós Rédei and Michael Stöltzner. Kluwer Academic Publishers, Dordrecht, Boston, London, 2001, ix + 371 pages, ISBN 0-7923-6812-6, 134.00 EUR / 119.00 USD /89.00 GBP (hardcover). Foundations of Physics, 32(6):981–985, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1016067531153.

Eatwell:1989:GT

[EMN89] John Eatwell, Murray Milgate, and Peter Newman, editors. Game Theory. The new Palgrave. Macmillan Reference Books, London, UK, 1989. ISBN 0-333-49536-5, 0-333-49537-3 (paperback). xii + 264 pp. LCCN HB144 .N48 1989.

Erdi:1985:ANN

[Érd85] Érdi Péter. Egy analógia nyomában: Neumann János a számológépr l és az agyról. (Hungarian) [A footsteps analogy of John von Neumann, the calculator and the computer brain]. Világosság, 26(2):81–85, ???? 1985. CODEN ???? ISSN ????

Erickson:2015:WGT

[Eri15] Paul Erickson. The World the Game Theorists Made. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-09703-X (hardcover), 0-226-09717-X (paperback), 0-226-09720-X (e-book). 390 pp. LCCN HB144 .E75 2015.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4. ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006934045-b.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006934045-d.html;   http://www.loc.gov/catdir/toc/fy0708/
2006934045.html.

Fleming:1969:IME

[FB69] Donald Fleming and Bernard Bailyn, editors. The intellectual migration: Europe and America, 1930–1960. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 1969. ISBN 0-674-33411-6, 0-674-33412-4. 748 pp. LCCN E169.1 .F6.

Freire:2022:OHH

[FBD^{+}22] Olival Freire, Guido Bacciagaluppi, Olivier Darrigol, Thiago Hartz, Christian Joas, Alexei Kojevnikov, and Osvaldo Pessoa, editors. The Oxford Handbook of the History of Quantum Interpretations. Oxford handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2022. ISBN 0-19-884449-2 (hardcover). xiv + 1296 pp. LCCN QC173.98 .O94 2022. URL https://global.oup.com/

academic/product/the-oxford-handbook-of-the-history-of-
quantum-interpretations-9780198844495.

Fernandez:1999:AER

[Fer99] Julio F. Fernández. Algorithm for exponential random numbers. Computer Physics Communications, 121–122:78–82, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002854.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Fryer:1955:CTJ

[FH55] K. D. Fryer and Israel Halperin. On the coordinatization theorem of J. von Neumann. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):432–444, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Field:2014:SNE

[Fie14] Alexander Field. Schelling, von Neumann, and the event that didn’t occur. GAMES, 5(1):53–89, February 25, 2014. ISSN 2073-

4336. URL https://www.mdpi.com/2073-4336/5/1/53.

Fields:2015:CAP

[Fie15] Chris Fields. Co-authorship proximity of A. M. Turing Award and John von Neumann Medal winners to the disciplinary boundaries of computer science. Scientometrics, 104(3):809–825, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1575-9.

Filep:2003:NJR

[Fil03] Filep László. Neumann János és a Riesz testvérek. (Hungarian) [John von Neumann and the Riesz brothers]. Természet világa, 134(3):80–??, 2003. CODEN ???? ISSN ????

Fitzpatrick:1999:ILE

[Fit99] Anne Fitzpatrick. Igniting the light elements: The Los Alamos Thermonuclear Weapon Project, 1942–1952. Thesis LA-13577-T, Los Alamos National Laboratory, Los Alamos, NM, USA, July

1999. xi + 45 pp. URL https://fas.org/sgp/othergov/doe/
lanl/docs1/00460048.pdf.

Fitzpatrick:2013:ILE

[Fit13] Anne C. Fitzpatrick. Igniting the light elements: the Los Alamos Thermonuclear Weapon Project, 1942-1952. Biblioscholar, ????, 2013. ISBN 1-288-82498-X. x + 325 (est.) pp. LCCN ????

Feuchtwang:1986:GGI

[FKC86] T. E. Feuchtwang, E. Kazes, and P. H. Cutler. Generalized gauge independence and the physical limitations on the von Neumann measurement postulate. Foundations of Physics, 16(12):1263–1284, December 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00732120.

Forsythe:1951:SJN

[For51] George E. Forsythe. Summary of John von Neumann’s lecture, various techniques used in connection with random digits. Applied mathematics series / National Bureau of Standards, 12(??):36–38, ???? 1951. ISSN 1049-4685.

Forsythe:1972:NCMa

[For72a] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Technical Report CS-TR-72-254, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1972. 21 pp.

Forsythe:1972:NCMb

[For72b] George E. Forsythe. Von Neumann’s comparison method for random sampling from the normal and other distributions. Mathematics of Computation, 26(120):817–826, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Formica:2010:NMS

[For10] Giambattista Formica. Von Neumann’s methodology of science: From incompleteness theorems to later foundational reflections. Perspectives on Science: Historical, Philosophical, Social, 18(??): 480–499, ???? 2010. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Fernandez:1996:FAR

[FR96] Julio F. Fernández and Juan Rivero. Fast algorithms for random numbers with exponential and normal distributions. Computers in Physics, 10(1):83–88, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168562. See [Fer99] for a proof

of the algorithm proposed here for exponential random numbers.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://adsabs.harvard.edu/abs/2005PhP.....7..204F;
 http://www.springerlink.com/content/1422-6960/.

Frank:2007:SCH

[Fra07] Tibor Frank. The Social Construction of Hungarian Genius 1867– 1930. Eötvös Loránd University, Budapest, 2007. 108 pp. LCCN DB988 .T526 2007.

Freeman:2013:JNC

[Fre13] W. J. Freeman. John von Neumann: The computer and the brain. In Palm and Aertsen [PA13], pages 239–240. ISBN 3-642-70913-

3. LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-3-642-70911-1_18.

Friedrichs:1980:NHS

[Fri80] K. O. Friedrichs. Von Neumann’s Hilbert space theory and partial differential equations. SIAM Review, 22(4):486–493, October 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frohlich:1993:HHS

[Frö93] Fröhlich Pál. Három hazai szakvélemény a fiatal wignerről és neumannról. (Hungarian) [Three young domestic opinions (??) and Wigner and von Neumann]. Fizikai Szemle (Budapest), 43(5):197– 201, 1993. CODEN FISZA6. ISSN ????

Frommer:2000:NCL

[Fro00] Andreas Frommer, editor. Numerical challenges in lattice quantum chromodynamics: joint interdisciplinary workshop of John von Neumann Institute for Computing, Jülich, and Institute of Applied Computer Science, Wuppertal University, August 1999, volume 15 of Lecture notes in computational science and engineering, 1439-7358. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67732-1. LCCN QC793.3.G38 N86 2000.

Fagin:2011:IPE

[FS11] Barry S. Fagin and Dale J. Skrien. IASSim: a programmable emulator for the Princeton IAS/von Neumann Machine. In Thomas J. Cortina, editor, Proceedings of the 42nd ACM Technical Symposium on Computer Science Education (SIGCSE 11), pages 359– 364. ACM Press, New York, NY, USA, 2011. ISBN 1-4503-0500-

8 (print). URL http://dl.acm.org/citation.cfm?id=1953163;
 http://www.cs.colby.edu/djskrien/IASSim/. See [FS12] for

an analysis and debugging of von Neumann’s computer programs.

Fagin:2012:DSG

[FS12] Barry Fagin and Dale Skrien. Debugging on the shoulders of giants: von Neumann’s programs 65 years later. Computer, 45 (11):59–68, November 2012. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.cs.colby.edu/ djskrien/IASSim/. See [FS11] for a description of the emulator on which von Neumann’s programs were run and debugged.

Fekete:1922:LNG

[FvN22] M. Fekete and John von Neumann. Über die Lage der Nullstellen gewisser Minimumpolynome. (German) [On the condition of the zeros of certain minimum polynomials]. Jahresbericht der Deutschen Mathematiker-Vereinigung (Stuttgart), 31(??):125–138, ???? 1922. Reprinted in [Tau61a, Paper 2].

Fermi:1951:TIIb

[FvN51] Enrico Fermi and John von Neumann. Taylor instability at the boundary of two incompressible liquids. Document AECU-2979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1951. Part 2.

Fermi:1955:TII

[FvN55] Enrico Fermi and John von Neumann. Taylor instability of incompressible liquids. U.S. Government Document AECU-2979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, November

19, 1955. 17 pp. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0042.pdf. Reprinted in [Tau63c, Paper

31].

Galway:1993:ESCb

[Gal93] Lionel Galway. On the edge: Statistics and computing John von Neumann and the origin of pseudo-random number generators. Chance, 6(3):57–59, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ginev:1997:IIP

[GC97] Dimitri Ginev and R. S. (Robert Sonné) Cohen, editors. Issues and images in the philosophy of science: scientific and philosophical essays in honour of Azarya Polikarov, volume 192 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4444-8 (hardcover), 94-011-5788-X. ISSN 0068-0346. xvii + 403 pp. LCCN Q174 .B67 vol. 192. URL https://link.springer.

com/book/10.1007/978-94-011-5788-9.

Grattan-Guinness:2005:LWW

[GGC^{+}05] I. Grattan-Guinness, Roger Cooke, et al., editors. Landmark writings in Western mathematics 1640–1940. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-444-50871-6 (cloth), 0-08-045744-4 (e-book). xvii + 1022 pp. LCCN QA21.

Godfrey:1993:CNP

[GH93] M. D. Godfrey and D. F. Hendry. The computer as von Neumann planned it. IEEE Annals of the History of Computing, 15(1):11–21, January–March 1993. CODEN IAHCEX. ISSN 1058-6180 (print),

1934-1547 (electronic). URL http://dlib.computer.org/an/
books/an1993/pdf/a1011.pdf;   http://qss.stanford.edu/
~godfrey/vonNeumann/edv-an.pdf;   http://www.computer.
org/annals/an1993/a1011abs.htm.

Glimm:1990:LJN

[GIS90] James Glimm, John Impagliazzo, and I. M. (Isadore Manuel) Singer, editors. The Legacy of John von Neumann: Proceedings of the Summer Research Institute on the Legacy of John von Neumann held at Hofstra University, Hempstead, New York, May 29–June 4, 1988, volume 50 of Proceedings of Symposia in Pure Mathematics. American Mathematical Society, Providence, RI, USA, 1990. ISBN 0-8218-1487-7. vii + 334 pp. LCCN QA29 .V66 L44 1990.

Giuntini:2001:ICN

[GL01] Roberto Giuntini and Federico Laudisa. The impossible causality: the no hidden variables theorem of John von Neumann. In John von Neumann and the foundations of quantum physics (Budapest, 1999), volume 8 of Vienna Circ. Inst. Yearb., pages 173– 188. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Guerraggio:2004:OQC

[GM04] Angelo Guerraggio and Elena Molho. The origins of quasiconcavity: a development between mathematics and economics. Historia Mathematica, 31(1):62–75, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086003000570.

Gillies:1953:TVP

[GMvN53] D. B. Gillies, J. P. Mayberry, and J. von Neumann. Two variants of poker. In Contributions to the theory of games, vol. 2, Annals of Mathematics Studies, no. 28, pages 13–50. Princeton University Press, Princeton, NJ, 1953. Reprinted in [Tau63c, Paper 6].

Goldstine:1959:JMR

[GMvN59] H. H. Goldstine, F. J. Murray, and J. von Neumann. The Jacobi method for real symmetric matrices. Journal of the ACM, 6(1): 59–96, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Reprinted in [Tau63b, Paper 16].

Godfrey:1974:BRB

[God74] Michael D. Godfrey. Book review: The Computer from Pascal to von Neumann, by H. H. Goldstine. Journal of the Royal Statistical Society. Series A (General), 137(4):638–639, ???? 1974. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344738.

Goldstine:1960:FRP

[Gol60] H. H. Goldstine. Footnote to a recent paper. Journal of the ACM, 7(1):78–79, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstine:1972:CPN

[Gol72] Herman Heine Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, 1972. ISBN 0-691-02367-0 (paperback), 0-691-08104-2. xii + 378 pp. LCCN

TK7885.A5 G64 1993. URL http://www.jstor.org/stable/j.

ctt7rvp0. Second printing, 1973. Paperback edition 1980. Fifth printing, 1993 with new preface. Reprint 2000 by Books on Demand, Ann Arbor, MI, USA.

Goldstine:1983:FTN

[Gol83a] Herman Goldstine. Faster than von Neumann? Two-Year College Mathematics Journal, 14(5):389, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00494925.1983.11972725.

Goldstine:1983:RJN

[Gol83b] Herman H. Goldstine. The role of John von Neumann in the computer field. In Aris et al. [ADS83], pages 308–327. ISBN 0-8166-1087-8. LCCN QC7 .S77 1983.

Goldman:1984:CTS

[Gol84] Louis Goldman. Commentary: Is there a Soviet bomb in Wichita? Bulletin of the Atomic Scientists, 40(1):54–56, January 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstine:1987:RED

[Gol87a] H. H. Goldstine. Reflections on the early days of the department. IBM Journal of Research and Development, 31(2):154–157, March 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Goldstine:1987:SPN

[Gol87b] Hermann H. Goldstine. A számítógép Pascaltól Neumannig. (Hungarian) [The computer from Pascal to von Neumann]. Muszaki Konyvkiado, Budapest, Hungary, second edition, 1987. ISBN 963-16-2771-3. 368 pp. LCCN ????

Golostine:1987:NJS

[Gol87c] Golostine Herman. Neumann János és a számítógép. (Riporter: Kenéz János). (Hungarian) [John von Neumann and the computer. (reporter: John Kenéz)]. Új impulzus, 3(9):17–??, ???? 1987. CO-DEN ???? ISSN ????

Goldstine:1993:CPN

[Gol93] Herman H. (Herman Heine) Goldstine. The computer from Pascal to von Neumann. Princeton University Press, Princeton, NJ, 1993. ISBN 0-691-02367-0 (paperback), 0-691-08104-2 (hardback). xii + 378 pp. LCCN TK7885.A5 G64 1993. URL http://www.jstor.

org/stable/j.ctt7rvp0.

Goldstine:2003:SPN

[Gol03] Hermann H. Goldstine. A számítógép Pascaltól Neumannig. (Hungarian) [The computer from Pascal to von Neumann]. Muszaki Konyvkiado, Budapest, Hungary, 2003. ISBN 963-10-7277-0. 375 pp. LCCN ????

Gotz:2003:NJK

[Göt03] Götz Gusztáv. Neumann János közreműködése a meteorológiában. (Hungarian) [John von Neumann’s co-operation in meteorology]. Természet világa, 134(3):65–69, 2003. CODEN ???? ISSN ????

Grafton:1956:MMW

[Gra56] Samuel Grafton. Married to a man who believes the mind can move the world. Good Housekeeping, ??(??):??, September 1956. Interview with Klari von Neumann.

Gray:1963:DCE

[Gra63] Harry J. (Harry Joshua) Gray. Digital Computer Engineering. Prentice-Hall international series in electrical engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963. 381 pp. LCCN TK7888.3 .G7.

Grcar:2005:JNB

[Grc05] J. F. Grcar. John von Neumann birthday centennial. SIAM News, 38(2):??, March 2005. ISSN 0036-1437. URL http://www.siam.

org/news/news.php?id=36.

Grcar:2011:JNA

[Grc11] Joseph F. Grcar. John von Neumann’s analysis of Gaussian elimination and the origins of modern numerical analysis. SIAM Review, 53(4):607–682, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://epubs.siam.org/

sirev/resource/1/siread/v53/i4/p607_s1.

Grcar:2012:JNA

[Grc12] Joseph F. Grcar. John von Neumann’s analysis of Gaussian elimination and the origins of modern numerical analysis. SIAM Review, 53(4):607–682, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://epubs.siam.org/

sirev/resource/1/siread/v53/i4/p607_s1.

Green:2010:SUS

[Gre10] David Green. The Sydney University SILLIAC. Web site, Au-

gust 14, 2010. URL http://members.iinet.net.au/~dgreen/

silliac.html. The SILLIAC was the first computer installed at Sydney University, and was operational from 1956 to 1968. The Web site links to the SILLIAC Emulator, a C program for Microsoft Windows.

Grier:2003:CJN

[Gri03] David Alan Grier. The computer: From John von Neumann to Irma S. Rombauer. IEEE Annals of the History of Computing, 25(4):104, 103, October/December 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

csdl.computer.org/dl/mags/an/2003/04/a4104.htm;   http:
//csdl.computer.org/dl/mags/an/2003/04/a4104.pdf.

Grzesik:1989:NRT

[Grz89] J. A. Grzesik. Von Neumann’s rejection technique reexamined. SIAM Review, 31(3):486–489, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenberger:1999:EEP

[GRZ99] Daniel M. Greenberger, Wolfgang L. Reiter, and Anton Zeilinger, editors. Epistemological and experimental perspectives on quantum physics, volume 7 of Vienna Circle Institute yearbook. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6063-X. ISSN 0929-6328. x + 377 pp. LCCN QC174.13 .E65 1999.

Goldstine:1945:PLS

[GvN45] Herman H. Goldstine and John von Neumann. On the principles of large scale computing machines. Unpublished. Reprinted in [GIS90, pp. 179–183]., 1945.

Goldstine:1947:MLA

[GvN47a] Herman H. Goldstine and John von Neumann. The mathematical and logical aspects of an electronic computing instrument. Technical report, Institute for Advanced Study, Princeton, NJ, USA, 1947.

Goldstine:1947:PCPa

[GvN47b] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 1, vol.

1. Technical report, Institute for Advanced Study, Princeton, NJ, USA, ???? 1947. ?? pp.

Goldstine:1947:PCPc

[GvN47c] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 1, vol.

2. Technical report, Institute for Advanced Study, Princeton, NJ, USA, ???? 1947. ?? pp.

Goldstine:1947:PCPe

[GvN47d] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 1, vol.

3. Technical report, Institute for Advanced Study, Princeton, NJ, USA, ???? 1947. ?? pp.

Goldstine:1947:PCPb

[GvN47e] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 2, vol. 1. Technical report 1, Institute for Advanced Study, Princeton, NJ, USA, April 1, 1947. 69 pp. Report prepared for U.S. Army Ordnance Department under contract W-36-034-OKD-7481. Reprinted in [Tau63b, 80–151]. Knuth [?, p. 278] cites pp. 142–151 of this report as the first published treatment of double-precision arithmetic on digital computers.

Goldstine:1947:PCPd

[GvN47f] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 2, vol.

2. Technical report, Institute for Advanced Study, Princeton, NJ, USA, ???? 1947. ?? pp.

Goldstine:1947:PCPf

[GvN47g] Herman H. Goldstine and John von Neumann. Planning and coding of problems for an electronic computing instrument. Part 2, vol.

3. Technical report, Institute for Advanced Study, Princeton, NJ, USA, ???? 1947. ?? pp.

Goldstine:1947:SEN

[GvN47h] Herman H. Goldstine and John von Neumann. Some estimates on the numerical stability of the elimination method for inverting matrices of high order. Bulletin of the American Mathematical Society, 53(??):1123–??, ???? 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Goldstine:1951:NIM

[GvN51] Herman H. Goldstine and John von Neumann. Numerical inverting of matrices of high order. II. Proceedings of the American Mathematical Society, 2:188–202, 1951. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://www.

jstor.org/view/00029939/di970628/97p0185x/0. See [vNG47]

for Part I. Reprinted in [Tau63b, Paper 15, pp. 558–572].

Goldstine:1955:BWC

[GvN55] Herman H. Goldstine and John von Neumann. Blast wave calculation. Communications on Pure and Applied Mathematics (New York), 8(2):327–353, May 1955. CODEN CPA-MAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/

cpa.3160080207/abstract. Reprinted in [Tau63c, Paper 29, pages 386–412].

Goldstine:1963:PLS

[GvN63] H. Goldstine and J. von Neumann. On the principles of large-scale computing machines. In Taub [Tau63b], pages ix + 784. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Goldstine:1957:SWJ

[GW57] Herman H. Goldstine and Eugene P. Wigner. The scientific work of John von Neumann. Science, 125(3250):683–684, April 12, 1957. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/125/3250/
683.full.pdf.

Goldstine:1996:SWJ

[GW96] H. H. Goldstine and E. P. Wigner. The scientific work of John von Neumann. In Wigner [Wig96c], pages 123–126. ISBN 3-540-56972-3. Reprinted from [GW57].

Goldstine:2003:NJT

[GW03a] Herman H. Goldstine and Wigner Jenő. Neumann János tudományos munkássága. (Hungarian) [Scientific work of John von Neumann]. Természet Világa [Natural World], 134(3):17–19, ???? 2003. CODEN ???? ISSN ????

Goldstine:2003:NJM

[GW03b] Goldstine Herman and Wigner Jenő. Neumann János munkássága. (Hungarian) [John von Neumann’s work]. Természet világa, 134(3): 17–19, 2003. CODEN ???? ISSN ????

Gyorgyi:1975:NJL

[Gyö75] Györgyi Géza. Neumann János levelei Ortvay Rudolfhoz. (Hungarian) [John von Neumann leaves rudolt ortvay]. Fizikai Szemle (Budapest), 25(5):166–179, ???? 1975. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Hajnal:1959:WJN

[Haj59] András Hajnal. The work of J. von Neumann in the axiomatic set-theory. Mat. Lapok, 10:5–11, 1959. ISSN 0025-519X.

Halmos:1958:NME

[Hal58] Paul R. Halmos. Von Neumann on measure and ergodic theory. Bulletin of the American Mathematical Society, 64(3 (part 2)):86– 94, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.
bams/1183522373.

Halmos:1973:LJN

[Hal73] Paul R. Halmos. The legend of John von Neumann. American Mathematical Monthly, 80(4):382–394, April 1973. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Halmos:1977:NLH

[Hal77] Halmos Pál. A Neumann-legenda. (Hungarian) [The Neumann legend]. Természet világa, 108(1):14–17, ???? 1977. CODEN ???? ISSN ????

Halperin:1985:BRS

[Hal85] Israel Halperin. Books in review: a survey of John von Neumann’s books on continuous geometry. Order (Dordrecht), 1(3):301–305, 1985. CODEN ORDRE5. ISSN 0167-8094 (print), 1572-9273 (electronic).

Halperin:1987:JNM

[Hal87] Israel Halperin. John von Neumann memorabilia. Historia Mathematica, 14(3):287, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900504.

Halperin:1990:EIJ

[Hal90] Israel Halperin. The extraordinary inspiration of John von Neumann. In Glimm et al. [GIS90], pages 15–17. ISBN 0-8218-1487-7. LCCN QA29 .V66 L44 1990.

Halmos:1991:LJN

[Hal91] P. R. Halmos. Legend of John von Neumann. In Ferris and Fadiman [FF91], pages 614–628. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Halmos:2003:NLH

[Hal03] Halmos Pál. A Neumann-legenda. (Hungarian) [The von Neumann legend]. Természet világa, 134(3):8–13, 2003. CODEN ???? ISSN ????

Halmos:2004:LJN

[Hal04] Paul R. Halmos. The legend of John von Neumann. In Doran and Kadison [DK04], pages 1–13. ISBN 0-8218-3402-9 (print), 0-8218-7955-3 (electronic). ISSN 0271-4132 (print), 1098-3627 (electronic).

LCCN QA612.33 .O64 2004. URL http://www.ams.org/books/

conm/365/. Reprint of [Hal73].

Hamming:1992:RJN

[Ham92] R. W. Hamming. Review of John von Neumann and the Origins of Modern Computing (Aspray, W., 1990). Computer, 25(10):126– 127, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hart:1942:TPR

[Har42] B. I. Hart. Tabulation of the probabilities for the ratio of the mean square successive difference to the variance. Note by John von Neumann. Annals of mathematical statistics, 13(2):207–214, 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/

1177731606. Reprinted in [Tau62, Paper 36].

Harrison:1967:BRB

[Har67] Michael A. Harrison. Book review: 4/67-1R Theory of Self-Reproducing Automata. 1966. John von Neumann. Arthur W. Burks, Editor. University of Illinois Press. American Documentation, 18(4):254, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Hartmanis:1989:GNPa

[Har89a] Juris Hartmanis. Gödel, von Neumann and the P =? NP problem. Technical Report TR89-994, Cornell University, Computer Science Department, April 1989. 8 pp.

Hartmanis:1989:GNPb

[Har89b] Juris Hartmanis. Gödel, von Neumann and the P =? NP problem. Bulletin of the European Association for Theoretical Computer Science, 38:101–107, June 1989. CODEN BEASDU. ISSN

0252-9742. URL http://people.cs.uchicago.edu/~fortnow/

beatcs/column38.ps. Columns: Structural Complexity.

Hargittai:2006:OVM

[Har06a] István Hargittai. Az öt világformáló marslakó. Vince, Budapest, Hungary, 2006. ISBN 963-9552-77-1. 397 + 32 pp. LCCN QC15 .H27155 2006.

Hargittai:2006:MSF

[Har06b] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp. LCCN QC15 .H27 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005029427-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005029427-b.html;   http://www.loc.gov/catdir/toc/ecip061/
2005029427.html.

Hashagen:2010:HJN

[Has10a] Ulf Hashagen. Die Habilitation von John von Neumann an der Friedrich-Wilhelms-Universität in Berlin: Urteile über einen ungarisch-jüdischen Mathematiker in Deutschland im Jahr 1927. (German) [The Habilitation of John von Neumann at the Friedrich Wilhelm University in Berlin: Judgments about a Hungarian Jewish mathematician in Germany in 1927]. Historia Mathematica, 37(2):242–280, May 2010. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0315086009000317. See erratum [Has10b].

Hashagen:2010:EHJ

[Has10b] Ulf Hashagen. Erratum to “Die Habilitation von John von Neumann an der Friedrich-Wilhelms-Universität in Berlin: Urteile über einen ungarisch-jüdischen Mathematiker in Deutschland im Jahr 1927” [Historia Mathematica 37 (2) (2010) 242–280] [MR2609069]. Historia Mathematica, 37(4):723, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086010000352. See [Has10a].

Hawkins:1945:BRB

[Haw45] David Hawkins. Book review: Theory of Games and Economic Behavior by John von Neumann; Oskar Morgenstern. Philosophy of Science, 12(3):221–227, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/184828.

Heims:1982:BRJ

[HB82] Steve J. Heims and Duane W. Bailey. Book review: John von Neumann and Norbert Wiener, From Mathematics to the Technologies of Life and Death. American Journal of Physics, 50(4):383, April 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/50/4/10.1119/1.12849.

Harman:2013:OSR

[HD13] Oren Solomon Harman and Michael R. Dietrich, editors. Outsider scientists: routes to innovation in biology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07837-X (hardcover), 0-226-07840-X (paperback), 0-226-07854-X (e-book). 374 pp. LCCN QH26 .O98 2013.

Heims:1977:GBM

[Hei77] Steve J. Heims. Gregory Bateson and the mathematicians: From interdisciplinary interaction to societal functions. Journal of the History of the Behavioral Sciences, 13(??):141–159, ???? 1977. CO-DEN JHBSA5. ISSN 0022-5061 (print), 1520-6696 (electronic).

Heims:1980:JNN

[Hei80] Steve J. Heims. John von Neumann and Norbert Wiener: from mathematics to the technologies of life and death. MIT Press, Cambridge, MA, USA, 1980. ISBN 0-262-08105-9. xviii + 547 pp. LCCN QA28 .H44. From mathematics to the technologies of life and death.

Henderson:2003:NER

[Hen03] Leah Henderson. The von Neumann entropy: a reply to Shenker. British Journal for the Philosophy of Science, 54(2):291–296, June 2003. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
54/2/291.full.pdf+html;   http://www.jstor.org/stable/
3541968.

Heppenheimer:1990:HNS

[Hep90] T. A. Heppenheimer. How von Neumann showed the way. American Heritage of Invention and Technology, 6(2):8–16, Fall 1990. CODEN AHETEU. ISSN 8756-7296.

Herken:1988:UTM

[Her88] Rolf Herken, editor. The Universal Turing machine: a half-century survey. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. ISBN 0-19-853741-7. xiv + 661 pp. LCCN QA267 .U55 1988.

Herken:1995:UTM

[Her95] Rolf Herken, editor. The Universal Turing machine: a half-century survey, volume 2 of Computerkultur, 0946-9613. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 3-211-82637-8 (paperback), 3-211-82628-9. xvi + 611 pp. LCCN QA267 .U55 1995.

Heywood:1947:WM

[Hey47] Robert B. Heywood, editor. The works of the mind. University of Chicago Press, Chicago, IL, USA, 1947. xi + 245 pp. LCCN BF408 .C5. Mortimer J. Adler and others, Ed. for the Committee on Social Thought. With a preface by John U. Nef.

Householder:1951:MCM

[HFG51] Alston S. Householder, George E. Forsythe, and Hallett-Hunt Germond, editors. Monte Carlo method. Proceedings of a Symposium Held June 29, 30 and July 1, 1949 in Los Angeles, California, volume 12 of Applied Mathematics Series / National Bureau of Standards. United States Government Printing Office, Washington, DC, USA, 1951.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hicks:1993:ANC

[HKLS93] D. L. Hicks, K. L. Kuttler, L. M. Liebrock, and C. E. Southwell. Answering von Neumann’s conjecture on convergence of averages and analyzing the von Neumann–Richtmyer scheme via a material averaging method. Applied Mathematics and Computation, 55(2–3):265–295, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/009630039390025A.

Hodes:1981:BRB

[Hod81] Elizabeth Hodes. Book review: John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death by Steve J. Heims. Isis, 72(3):500–501, September 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230276.

Holland:1962:OLT

[Hol62] John H. Holland. Outline for a logical theory of adaptive systems. Journal of the ACM, 9(3):297–314, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Early work on genetic algorithms. Reprinted in [Bur70].

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (e-book). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hon:2009:LEF

[Hon09] Giora Hon. Living extremely flat: The life of an automaton; John von Neumann’s conception of error of (in)Animate systems. In Going Amiss In Experimental Research [HSS09], chapter 5, pages 55–71. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346. LCCN Q172.5.E77 G647 2009. URL https://

link.springer.com/chapter/10.1007/978-1-4020-8893-3_5.

Horvath:1997:TKE

[Hor97] T. Horvath. Theodore Kármán, [Eugene] Paul Wigner, John [von] Neumann, Leo Szilard, Edward Teller and their ideas of ultimate reality and meaning. Ultimate Reality and Meaning, 20(2–3):123– 146, June/September 1997. ISSN 0709-549X.

Howell:1988:JN

[How88] Catherine Herbert Howell. John von Neumann. In Inventors and discoverers: changing our world [Nat88], page ?? ISBN 0-87044-751-3, 0-87044-752-1, 0-87044-753-X. LCCN T18 .I57 1988.

Haigh:2020:HRN

[HP20] Thomas Haigh and Mark Priestley. Historical reflections: von Neumann thought Turing’s universal machine was ‘simple and neat.’: but that didn’t tell him how to design a computer. Communications of the ACM, 63(1):26–32, January 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3372920.

Haigh:2014:ABE

[HPR14a] T. Haigh, M. Priestley, and C. Rope. Los Alamos bets on ENIAC: Nuclear Monte Carlo simulations, 1947–1948. IEEE Annals of the History of Computing, 36(3):42–63, July 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2014:EXM

[HPR14b] Thomas Haigh, Mark Priestley, and Crispin Rope. Engineering “The miracle of the ENIAC”: Implementing the modern code paradigm. IEEE Annals of the History of Computing, 36(2):41– 59, 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Haigh:2014:RSP

[HPR14c] Thomas Haigh, Mark Priestley, and Crispin Rope. Reconsidering the stored-program concept. IEEE Annals of the History of Computing, 36(1):4–17, 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hon:2009:GAE

[HSS09] Giora Hon, Jutta Schickore, and Friedrich Steinle. Going Amiss In Experimental Research, volume 267 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-8892-2 (hardcover), 1-4020-8893-0 (e-book). ISSN 0068-0346. x + 277 pp. LCCN Q172.5.E77 G647 2009. URL https://link.springer.

com/book/10.1007/978-1-4020-8893-3.

Hurwicz:1948:RJN

[Hur48] Leonid Hurwicz. Review: John von Neumann, Oskar Morgenstern, The Theory of Games and Economic Behavior. Annals of mathematical statistics, 19(3):436–437, September 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177730216.

Hurd:1985:NEM

[Hur85] Cuthbert C. Hurd. A note on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):141– 155, April/June 1985. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1985/pdf/a2141.pdf;
 http://www.computer.org/annals/an1985/a2141abs.htm. In-

cludes typeset reprint of [RUvN47].

Halmos:1942:OMC

[HvN42] Paul R. Halmos and John von Neumann. Operator methods in classical mechanics. II. Annals of Mathematics, 43(??):332–350, ???? 1942. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 23].

Hilbert:1927:GQG

[HvNN27] David Hilbert, John von Neumann, and L. Nordheim. Über die Grundlagen der Quantenmechanik. (German) [On the foundations of quantum mechanics]. Mathematische Annalen, 98(??):1–30, ???? 1927. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61a, Paper 7].

Israel:2002:NMP

[IG02] Giorgio Israel and Ana Millán Gasca. Von Neumann: La matematica per il dominio della realtà. (Italian) [Von Neumann: Mathematics for the dominion of reality]. Grandi della Scienza, 5(??): 4–95, ???? 2002.

Israel:1995:MCG

[IM95] Giorgio Israel and Ana Millán Gasca. Il mondo come gioco matematico: John von Neumann, scienziato del Novecento. (Italian) [The world as mathematical game: John von Neumann, scientist of the 1900s], volume 249 of Studi superiori NIS. La Nuova Italia Scientifica, Roma, Italy, 1995. ISBN 88-430-0319-4. 158 pp. LCCN QA29.V66 I87 1995.

Israel:2001:MCJ

[IM01] Giorgio Israel and Ana Millán Gasca. El mundo como un juego matemático: John von Neumann, un científico del siglo XX. (Spanish) [The world as a mathematical game: John von Neumann, a scientist of the Twentieth Century], volume 2 of Ciencia abierta. Nivola, Madrid, Spain, 2001. ISBN 84-95599-11-2. 182 pp. LCCN QA8.4 .I77 2001.

Israel:2008:MCG

[IM08] Giorgio Israel and Ana Millán Gasca. Il mondo come gioco matematico: la vita e le idee di John von Neumann. (Italian) [The world as a mathematical game: the life and ideas of John von Neumann]. Saggi. Scienze. Bollati Boringhieri, Torino, Italy, 2008. ISBN 88-339-1871-8. 273 pp. LCCN QA29.V66 I87 2008. 22.00 EUR. URL

http://www.loc.gov/catdir/toc/casalini07/08667829.pdf.

Israel:2009:BMN

[IM09a] Giorgio Israel and Ana Millán Gasca. Beyond mathematics: von Neumann’s scientific activity in the 1940s and 1950s. In The World as a Mathematical Game: John von Neumann and Twentieth Century Science [IM09b], pages 121–175. ISBN 3-7643-9895-7 (hardcover). LCCN QA29.V66 I8713 2009.

Israel:2009:WMG

[IM09b] Giorgio Israel and Ana Millán Gasca. The World as a Mathematical Game: John von Neumann and Twentieth Century Science, volume 38 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2009. ISBN 3-7643-9895-7 (hardcover). xii + 207 pp. LCCN QA29.V66 I8713 2009.

Istrail:2013:ATJ

[IM13] Sorin Istrail and Solomon Marcus. Alan Turing and John von Neumann — their brains and their computers. Lecture Notes in Computer Science, 7762:26–35, 2013. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-36751-9_2.

IAS:2012:CSI

[Ins12] Institute for Advanced Study, editor. A community of scholars: impressions of the Institute for Advanced Study. Princeton University Press, Princeton, NJ, 2012. ISBN 0-691-15136-9 (hardcover), 1-4008-3979-3 (e-book). xvi + 110 pp. LCCN LD2523.I33 C66 2012. Photographs by Serge J.-F. Levy.

Pae:2013:EOR

[iP13] Sung il Pae. Exact output rate of Peres’s algorithm for random number generation. Information Processing Letters, 113(5–6):160– 164, March 15, 2013. CODEN IFPLAT. ISSN 0020-0190 (print),

1872-6119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0020019013000033.

Isaacson:2014:IHG

[Isa14] Walter Isaacson. The Innovators: How a Group of Hackers, Geniuses and Geeks Created the Digital Revolution. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4711-3879-8 (hardcover), 1-4767-0869-X (cloth), 1-4711-3897-6 (paperback), 1-4104-7497-6 (cloth), 1-4767-0870-3. viii + 542 pp. LCCN QA76.2.A2.

Jammer:1974:HVT

[Jam74a] Max Jammer. Hidden-variable theories. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74c], chapter 7, pages 252–339. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:MCQ

[Jam74b] Max Jammer. Measurement in classical and in quantum physics. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74c], chapter 7, pages 471–473. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74c] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

James:2002:RME

[Jam02] I. M. (Ioan Mackenzie) James. Remarkable mathematicians: from Euler to von Neumann. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2002. ISBN 0-521-81777-3 (hardcover), 0-521-52094-0 (paperback). xiv + 433 pp. LCCN QA28 .J36 2002. URL http://www.loc.

gov/catdir/description/cam022/2002022266.html;   http:
//www.loc.gov/catdir/enhancements/fy0731/2002022266-b.
html;   http://www.loc.gov/catdir/samples/cam033/2002022266.
html;   http://www.loc.gov/catdir/toc/cam026/2002022266.
html.

Center:1987:SJN

[JC87] Science at the John von Neumann National Supercomputer Center: annual research report, 1987. ISSN 1046-0632. Consortium for Scientific Computing, Princeton, NJ, USA.

Jeffress:1951:CMB

[Jef51] Lloyd A. Jeffress, editor. Cerebral mechanisms in behavior: the Hixon symposium, September 20, 1948, Pasadena. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1951. LCCN BF181 .J4; QP351. Reprinted 1967.

Jeki:2003:NJNa

[Jék03a] Jéki László. Neumann János és a nukleáris fegyverek. (Hungarian) [John von Neumann and nuclear weapons]. Fizikai Szemle (Budapest), 53(2):423–425, 2003. CODEN FISZA6. ISSN ????

Jeki:2003:NJNb

[Jék03b] Jéki László. Neumann János és a nukleáris fegyverek. (Hungarian) [John von Neumann and nuclear weapons]. Természet világa, 134 (3):62–64, 2003. CODEN ???? ISSN ????

Jun:1999:IRN

[JK99] Benjamin Jun and Paul Kocher. The Intel random number generator. White paper prepared for Intel Corporation, Cryptography Research, Inc., Menlo Park, CA, USA, April 22, 1999. URL http:

//www.cryptography.com/intelRNG.pdf.

Ji:2013:CAM

[JML13] Hao Ji, Michael Mascagni, and Yaohang Li. Convergence analysis of Markov Chain Monte Carlo linear solvers using Ulam–von Neumann algorithm. SIAM Journal on Numerical Analysis, 51(4): 2107–2122, ???? 2013. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

XXX:1986:JNN

[Joh86] The John von Neumann National Supercomputer Center, Princeton, NJ, USA. The John von Neumann National Supercomputer Center, 1986. 4 pp.

Jones:2004:BRI

[Jon04] Claire Jones. Book review: Ioan James, Remarkable Mathematicians: From Euler to von Neumann. Spectrum Series of the Mathematical Association of America. Cambridge: Cambridge University Press, 2002. Pp. xiv + 433. ISBN 0-521-52094-0. £19.95, $25.00 (paperback). British Journal for the History of Science, 37(4): 487–488, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4028663.

Jordan:1935:IPL

[JvN35] P. Jordan and J. von Neumann. On inner products in linear, metric spaces. Annals of Mathematics (2), 36(3):719–723, ???? 1935. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 4].

Jordan:1934:AGQ

[JvNW34] P. Jordan, J. von Neumann, and E. Wigner. On an algebraic generalization of the quantum mechanical formalism. Annals of Mathematics (2), 35(1):29–64, ???? 1934. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 21].

K:1944:BRBa

[K.44] M. G. K. Book review: Theory of Games and Economic Behaviour, by John von Neumann; Oskar Morgenstern. Journal of the Royal Statistical Society, 107(3/4):293, ???? 1944. ISSN 0952-8385. URL

http://www.jstor.org/stable/2981222.

Kac:1982:RJN

[Kac82] Mark Kac. Reviews: John von Neumann and Norbert Wiener. From mathematics to the technologies of life and death. American Mathematical Monthly, 89(9):713–714, 1982. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kadison:1958:TOP

[Kad58] Richard V. Kadison. Theory of operators. Part II. Operator algebras. Bulletin of the American Mathematical Society, 64(3 (part 2)):61–85, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183522372.

Kahan:1997:JNL

[Kah97] W. Kahan. The John von Neumann lecture on the baleful effect of computer languages and benchmarks upon applied mathematics, physics and chemistry. World-Wide Web document., July 15, 1997.

URL http://www.cs.berkeley.edu/~wkahan/SIAMjvnl.ps. Pre-

sented at the SIAM 45th annual meeting, Stanford University.

Kalai:2013:FHP

[Kal13] Ehud Kalai. Foreword: The high priest of game theory. American Mathematical Monthly, 120(5):384–385, May 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.4169/amer.
math.monthly.120.05.384.pdf.

Kaplansky:1995:CNT

[Kap95] Irving Kaplansky. A contribution to von Neumann’s theory of games. II. Linear Algebra and its Applications, 226/228(1–3): 371–373, September/October 1995. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/tree/store/laa/cas_sub/browse/
browse.cgi?year=1995&volume=226-228&issue=1-3&aid=9500167.

Kardos:2004:UZN

[Kar04] Kardos István. Egy univerzális zseni : Neumann János (1903– 1957). (Hungarian) [A universal genius John von Neumann (1903– 1957)]. Ezredvég, 2(??):94–96, 2004. CODEN ???? ISSN ????

Kato:1960:EIM

[Kat60] Tosio Kato. Estimation of iterated matrices, with application to the von Neumann condition. Numerische Mathematik, 2:22–29, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kay:1997:CIL

[Kay97] Lily E. Kay. Cybernetics, information, life: The emergence of scriptural representations of heredity. Configurations: a Journal of Literature, Science, and Technology, 5(??):23–91, ???? 1997. CODEN ???? ISSN 1063-1801 (print), 1080-6520 (electronic).

Kemeny:1955:MVM

[Kem55] John G. Kemeny. Man viewed as a machine. Scientific American, 192(4):58–67, April 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v192/n4/pdf/
scientificamerican0455-58.pdf.

Kevles:1981:JNN

[Kev81] Daniel J. Kevles. Book review: John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death. By Steve J. Heims. Cambridge: MIT Press, 1980. xix + 547 pp. Charts, tables, illustrations, notes, and index. $19.95. The Journal of American History, 68(2):432–433, 1981. ISSN 0021-8723 (print), 1945-2314 (electronic). URL http:/

/jah.oxfordjournals.org/content/68/2/432.full.pdf+html;
 http://jah.oxfordjournals.org/content/68/2/432.short.

Khrennikov:2008:RNL

[Khr08] Andrei Khrennikov. The role of von Neumann and Lüders postulates in the Einstein, Podolsky, and Rosen considerations: Comparing measurements with degenerate and nondegenerate spectra. Journal of Mathematical Physics, 49(5):052102, May 2008. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/
i5/p052102_s1.

Kidwell:1992:JNO

[Kid92] Peggy Aldrich Kidwell. John von Neumann and the Origins of Modern Computing by William Aspray (review). Technology and Culture, 33(1):174–175, January 1992. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/888361/pdf.

Kjeldsen:2001:JNC

[Kje01] Tinne Hoff Kjeldsen. John von Neumann’s conception of the Minimax Theorem: a journey through different mathematical contexts. Archive for History of Exact Sciences, 56(1): 39–68, November 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=56&
issue=1&spage=39.

Krips:1995:SRR

[KMM95] Henry Krips, J. E. McGuire, and Trevor Melia, editors. Science, Reason, and Rhetoric. Pittsburgh-Konstanz series in the philosophy and history of science. University of Pittsburgh Press, Pittsburgh, PA, USA, 1995. ISBN 0-8229-3912-6. xix + 322 pp. LCCN Q175 .S4225 1995. URL http://www.gbv.de/

dms/bowker/toc/9780822939122.pdf;   http://www.loc.gov/
catdir/enhancements/fy0709/95033117-d.html.

Knuth:1970:VNF

[Knu70] Donald E. Knuth. Von Neumann’s first computer program. ACM Computing Surveys, 2(4):247–260, December 1970. CODEN CMSVAN. ISSN 0010-4892. Reprinted in [AB87].

Knuth:1998:SA

[Knu98] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Korek:1986:NJP

[Kor86] Korek Valéria. Neumann János papucsban. (közr. és bev. Vezér Erzsébet). (Hungarian) [John von Neumann’s slippers. (?? and ?? by Elizabeth Vezer)]. Élet és Irodalom [Life and Literature], 30(10): 4–??, ???? 1986. CODEN ???? ISSN 0424-8848 (print), 1588-0362 (electronic).

Korosne:1999:ANH

[Kőr99] Kőrösné Mikis Márta. Akik nyomot hagytak a 20. századon : Neumann Jánostól az internetig. (Hungarian) [Those who left their mark on the 20th Century: John von Neumann and the Internet]. Új pedagógiai szemle, 49(7-8):229–231, 1999. CODEN ???? ISSN ????

Kovacs:1984:HLN

[Kov84] Kovács Gyözö. Hogyan lett Neumann Jánosból a ’számítógépek’ atyja?. (Hungarian) [How did John von Neumann become the father of computers?]. Élet és Tudomány [Life and Science], 39(42): 1315–1317, October 19, 1984.

Gyozo:1987:ENJ

[Kov87a] Kovács Gyözö. Emlekkepek Neumann Jánosrol. (Hungarian) [Memories of John von Neumann]. Mikroszámítógép magazin: a Neumann János Számítógéptudományi Társaság kiadványa, ?? (??):??, April 1987. CODEN ???? ISSN 0236-6088.

Kovacs:1987:NJH

[Kov87b] Kovács Gyözö. Neumann János. (Hungarian) [John von Neumann]. Magyar feltalálók, találmányok. Müszaki Kvk., Budapest, Hungary, 1987. ISBN 963-16-1272-4. ISSN 1416-7840. 43 pp. LCCN ????

Kovacs:1989:MSH

[Kov89] Kovács Gyözö. Magyarok a szamitastechnikaban. (Hungarian) [Hungarian Pioneers in Computer Technology]. Technikatorteneti Szemle XVII 1988–1989. Müszaki Kvk., Budapest, Hungary, 1989. ISBN ???? ???? pp. LCCN ????

Kovacs:1994:JNH

[Kov94] Gyözö Kovács. John von Neumann, A Hungarian born computer pioneer. In Brunnstein and Raubold [BR94], pages 285–292. ISBN 0-444-81990-8 (set), 0-444-81987-8 (vol. 2), 0-444-81988-6 (vol. 3). ISSN 0926-5473. LCCN QA75.5 .I38 1994.

Kovacs:1999:KVN

[Kov99] Kovács Győző. Ki volt Neumann János?. (Hungarian) [Who was John von Neumann?]. Élet és Tudomány [Life and Science], 37 (??):1158–1161, 1999. CODEN ???? ISSN ????

Kovacs:2003:LRJ

[Kov03a] László Kovács. László Rátz and John von Neumann: a gifted teacher and his brilliant pupil. Faculty of Education, University of Manitoba, Winnipeg, MB, Canada, 2003. ISBN 0-9695481-5-X (paperback). 107 pp. LCCN QA29 V66K68 2003.

Kovacs:2003:MEE

[Kov03b] Kovács Győző. Egy marslakó emlékezete. 100 éve született Neumann János a modern számítógép több fontos alapelvének megalkotója. (Hungarian) [A Martian memory. born 100 years ago, John von Neumann, the important principle architect of the modern computer]. Természet világa, 134(3):31–35, 2003. CODEN ???? ISSN ????

Kovacs:2003:NJE

[Kov03c] Kovács Győző. Neumann János élete és munkássága. (Hungarian) [John von Neumann: Life and works]. IME, 2(8):50, 2003. CODEN ???? ISSN ????

Kovacs:2003:ZPH

[Kov03d] Kovács Győző. Zseni és polihisztor. (Hungarian) [genius and polymath]. Élet és Tudomány [Life and Science], 51–52(??):1642–1646, 2003. CODEN ???? ISSN ????

Kovacs:2003:NJM

[Kov03e] Kovács László. Neumann János magyar tanárai. (Hungarian) [John von Neumann’s Hungarian teachers]. Természet világa, 134(3):36– 42, 2003. CODEN ???? ISSN ????

Kowald:2015:WTN

[Kow15] Axel Kowald. Why is there no von Neumann probe on Ceres? Error catastrophe can explain the Fermi–Hart paradox. Journal of the British Interplanetary Society, 68(??):383–338, ???? 2015. CODEN JBISAW. ISSN 0007-084X. URL https://arxiv.org/

abs/1605.02169.

Krips:1995:RID

[Kri95] Henry Krips. Rhetoric, ideology, and desire in von Neumann’s ‘Gründlagen’. In Krips et al. [KMM95], pages 279–293. ISBN 0-8229-3912-6. LCCN Q175 .S4225 1995. URL http://www.gbv.de/

dms/bowker/toc/9780822939122.pdf;   http://www.loc.gov/
catdir/enhancements/fy0709/95033117-d.html.

Kennaway:1984:TSN

[KS84] J. Kennaway and M. Sleep. Towards a successor to von Neumann. In Chambers et al. [CDJ84], pages 111–124. ISBN 0-12-167350-2. LCCN QA76.9.D5 D49 1984.

Kuhn:1958:JNW

[KT58] H. W. Kuhn and A. W. Tucker. John von Neumann’s work in the theory of games and mathematical economics. Bulletin of the American Mathematical Society, 64(3 (part 2)):100–122, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.
bams/1183522375.

Kumar:2021:MFA

[Kum21] Manjit Kumar. The Man from the Future by Ananyo Bhattacharya review — the life of John von Neumann. The Guardian, ??(??):??, October 13, 2021. ISSN 0261-3077 (print), 1756-3224 (electronic).

URL https://www.theguardian.com/books/2021/oct/13/the-
man-from-the-future-by-ananyo-bhattacharya-review-the-
life-of-john-von-neumann.

Koopman:1932:DSC

[KvN32] B. O. Koopman and J. von Neumann. Dynamical systems of continuous spectra. Proceedings of the National Academy of Sciences of the United States of America, 18(??):255–263, ???? 1932. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau61b, Paper 14].

Kuratowski:1937:SAS

[KvN37] C. Kuratowski and J. von Neumann. On some analytic sets defined by transfinite induction. Annals of Mathematics (2), 38(2):521– 525, ???? 1937. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 19].

Kunze:1997:LCP

[KvN97] M. Kunze and J. von Neumann. Linear complementary problems and the simulation of the motion of rigid body systems subject to Coulomb friction. Zeitschrift für Angewandte Mathematik und Mechanik, 77(11):833–838, 1997. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Kovetz:1963:EES

[KvNS63] A. Kovetz, John von Neumann, and R. Suzi. Extension of Ekman’s spiral to conditions of near-neutral static stability and a constant horizontal temperature gradient. Beitr. Phys. Atmos., 36:201–212, 1963.

Kraft:1964:MCM

[KW64] Rose Kraft and Carl J. Wensrich. Monte Carlo methods: a bibliography covering the period 1949 to 1963. Report UCRL-7823 and UC-32 and TID-4500 (32nd edition), University of California Radiation Laboratory, Berkeley, CA, USA, April 1, 1964.

iii + 110 pp. URL https://catalog.hathitrust.org/Record/

007292355. Supersedes report UCRL-658.

Labatut:2023:M

[Lab23] Benjamín Labatut. The Maniac. Penguin, London, UK and New York, NY, USA, 2023. ISBN 0-593-65447-1 (hardcover), 0-593-65448-X (e-book). LCCN PR9309.9.L33.

Lacki:2000:EAQ

[Lac00] Jan Lacki. The early axiomatizations of quantum mechanics: Jordan, von Neumann and the continuation of Hilbert’s program. Archive for History of Exact Sciences, 54(4):279– 318, January 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=54&
issue=4&spage=279.

Ladanyi:1993:HHS

[Lad93] Ladányi Andor. Három hazai szakvélemény a fiatal Wignerről és Neumannról. (Hungarian) [Three young domestic opinions (??) and Wigner and von Neumann]. Hiány, 4(6-7):56–60, 1993. CO-DEN ???? ISSN ????

Laffont:1989:EUI

[Laf89] Jean-Jacques Laffont. The Economics of Uncertainty and Information. The MIT Press, Cambridge, Massachusetts, London, England, 1989. 2nd printing, 1990, originally in French, 1986, Economica Paris.

Lanouette:1989:VCL

[Lan89] William Lanouette. The varied career of Leo Szilard. Rockefeller Archive Center Newsletter, ??(??):7–??, Summer 1989. An article about Szilard‘s many affiliations with the Rockefeller Foundation and University, as revealed in the Rockefeller Archive Center holdings. Items included Szilard’s physics research in New York in the 1930s, settling refugee scholars, friendships with Einstein and von Neumann, information theory, microbiology, nuclear power, and arms-control initiatives.

Lanouette:1993:HGT

[Lan93a] William Lanouette. Hungarian geniuses and their bomb. Seminar on Science and Social Responsibility, about Szilard, Teller, von Neumann, and Wigner. Georgetown University, Washington, DC, USA., May 10, 1993.

Lanouette:1993:WFM

[Lan93b] William Lanouette. Where the four Martians of the Manhattan Project landed: The national laboratory connections of Teller, Wigner, von Neumann, and Szilard. Talk at the American Physical Society, Washington, DC, USA., April 13, 1993.

Lanouette:1999:WFM

[Lan99] William Lanouette. Where the four Martians of the Manhattan Project landed: The National Laboratory connections of Teller, Wigner, von Neumann, and Szilard. A Talk at a Conference on Laboratory History and Sociology, State University of New York at Stony Brook & Brookhaven National Laboratory, June 10, 1999.

Lance:2008:JNS

[Lan08] Christopher Lance. John von Neumann: selected letters. Bulletin of the London Mathematical Society, 40(6):1096–1098, 2008. CO-DEN LMSBBT. ISSN 0024-6093 (print), 1469-2120 (electronic). Book review of [vN05].

Lanouette:2011:FHM

[Lan11] William Lanouette. The Four Hungarians of the Manhattan Project. Washington Magyar Club. Described the roles of Leo Szilard, Eugene Wigner, John von Neumann, and Edward Teller in the Manhattan Project., January 15, 2011.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN QA76 .G686 1996.

US$23.95. URL http://bit.csc.lsu.edu/~chen/GreatPapers.
html.

Lax:1990:RJN

[Lax90] Peter D. Lax. Remembering John von Neumann. In The legacy of John von Neumann (Hempstead, NY, 1988), volume 50 of Proc. Sympos. Pure Math., pages 5–7. American Mathematical Society, Providence, RI, USA, 1990.

Lax:2003:NJK

[Lax03] Lax Péter. Neumann János korai évei, a los alamos-i évek és a számítástechnikához vezető út. (Hungarian) [John von Neumann’s early years, the Los Alamos years, and leading the computing road]. Magyar Tudomány [Hungarian Science], 48(12):1508–1512, 2003. CODEN ???? ISSN ????

Lax:2005:JNE

[Lax05] Peter Lax. John von Neumann: The early years, the years at Los Alamos, and the road to computing. SIAM News, 38(2):??, March

2005. ISSN 0036-1437. URL http://www.siam.org/news/news.
php?id=39.

Leary:1955:UK

[Lea55] Lewis Leary, editor. The Unity of Knowledge. Doubleday and Company, New York, NY, USA, 1955. xiv + 306 pp.

Leonard:2006:CCP

[Leo06] Robert Leonard. From chess to catastrophe: Psychology, politics and the genesis of von Neumann’s game theory. Note de recherche 2006-04, Centre interuniversitaire de recherche sur la science et la technologie, Université de Québec à Montréal, Montréal, Québec, Canada, April 2006. ISBN 2-923333-21-7. iv + 82 pp. URL https:

//depot.erudit.org/bitstream/001204dd/1/2006_04.pdf.

Leonard:2010:NMC

[Leo10] Robert Leonard. Von Neumann, Morgenstern, and the creation of game theory: from chess to social science, 1900–1960. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-56266-X (hardback). x + 390 pp. LCCN HB144 .L46 2010.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Liebig:2011:KZE

[Lie11] Hans Liebig. Konrad Zuse, Erfinder des Computers — im Vergleich mit Alan Turing und John v. Neumann. (German) [Konrad Zuse, inventor of the computer — in comparison with Alan Turing and John von Neumann]. Informatik Spektrum, 34(6): 553–564, December 2011. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL http://www.springerlink.

com/content/h1j0r5m668715865/. Special issue: Konrad Zuse.

Lighthill:1976:FHJ

[Lig76] James Lighthill. Flagellar hydrodynamics — John von Neumann Lecture, 1975. SIAM Review, 18(2):161–230, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lim:2022:AAP

[LLY22] Zhao Ging Lim, Chen-Tuo Liao, and Yi-Ching Yao. Asymptotic analysis of Peres’ algorithm for random number generation. Probability in the Engineering and Informational Sciences, 36(2):341– 356, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951

(electronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/asymptotic-analysis-of-peres-algorithm-for-random-
number-generation/3B02FFEA433E0788CFD05B30974E83D3. See

Yuval Peres, Iterating von Neumann’s Procedure for Extracting Random Bits, Annals of Statistics 20(1) 590–597, March 1992, doi:10.1214/aos/1176348543.

Lorenzo:2019:AMH

[Lor19] Mark Jones Lorenzo. Abstracting Away the Machine: The History of the Fortran Programming Language (FORmula TRANslation). SE Books, 5307 West Tyson Street, Philadelphia, PA 19107, USA, 2019. ISBN 1-0823-9594-3. 325 pp. LCCN QA76.73.F25 BT22 2019.

Lieberman:1967:EDC

[LR67] Henry D. Lieberman and Louis Robinson. The electronic digital computer: How it started, how it works and what it does. New York Times, ??(??):136–137, January 9, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/118029058?accountid=14677.

Legendi:1983:LWJ

[LS83] Tamás Legendi and Tibor Szentivanyi, editors. Leben und Werk von John von Neumann: ein zusammenfassender überblick. (German) [The Life and Work of John von Neumann: a summary overview]. Bibliographisches Institut, Mannheim, Germany; Wien, Austria; Zürich, Switzerland, 1983. ISBN 3-411-01639-6. 151 pp. LCCN QA76.2.V6 N414 1983g. German translation by Rozsa Nienhaus of Hungarian original, Neumann János élete és munkássága.

Lucchetti:2011:TIP

[Luc11] Roberto Lucchetti. The theoretical intelligence and the practical vision of John von Neumann. In Mathematical lives, pages 63–68. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011.

Lupher:2003:BRB

[Lup03] Tracy Lupher. Book review: John von Neumann and the foundations of quantum physics: (Vienna Circle Institute yearbook (2000), 8) Miklos Redei and Michael Stoltzner (Eds.); Kluwer Academic Publishers, Dordrecht, 2001, pp., US $125, ISBN 0-7923-6812-6. Studies in History and Philosophy of Modern Physics, 34(4):684–687, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219803000704.

See [RS01a].

Lupacchini:2018:JNM

[Lup18] Rossella Lupacchini. John von Neumann: Metamatematica hilbertiana e fondamenti della meccanica quantistica. (Italian) [John von Neumann: Hilbertian metamathematics and foundations of quantum mechanics], volume 11 of Saggi di scienze e filosofia naturale. Bibliopolis, Napoli, Italia, 2018. ISBN 88-7088-661-1. 178 pp. LCCN ????

Lusito:2019:BRJ

[Lus19] Fabio Lusito. Book review: John von Neumann, Matematica hilbertiana e fondamenti della meccanica quantistica, by Rossella Lupacchini (ed.). Nuncius, 34(2):475–477, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/2/article-p475_19.xml.

Lorch:1947:ECP

[LvN47] E. R. Lorch and John von Neumann. On the Euclidean character of the perpendicularity relation. Bulletin of the American Mathematical Society, 53(??):489–??, ???? 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Li:2013:NAM

[LY13] Jiequan Li and Zhicheng Yang. The von Neumann analysis and modified equation approach for finite difference schemes. Applied Mathematics and Computation, 225(??):610–621, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313010254.

MacCagg:1972:JNG

[Mac72] William O. MacCagg, Jr. Jewish nobles and geniuses in modern Hungary, volume 3 of East European monographs. East European Monographs, Boulder, CO, USA, 1972. 254 pp. LCCN DS135.H9 M38 1972.

MacCagg:1986:JNG

[Mac86] William O. MacCagg, Jr. Jewish nobles and geniuses in modern Hungary, volume 3 of East European monographs. East European Monographs, Boulder, CO, USA, reprint edition, 1986. ISBN 0-88033-092-9. 254 pp. LCCN DS135.H9 M38 1972.

Macrae:1992:JNSa

[Mac92a] Norman Macrae. John von Neumann: The Scientific Genius Who Pioneered the Modern Computer, Game Theory, Nuclear Deterrence, and Much More. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-41308-1. x + 405 pp. LCCN QA29.V66 M33 1992. US$25.00 (US$31.50 Can.).

Macrae:1992:JNSb

[Mac92b] Norman Macrae. John von Neumann: The Scientific Genius Who Pioneered the Modern Computer, Game Theory, Nuclear Deterrence, and Much More. American Mathematical Society, Providence, RI, USA, 1992. ISBN 0-8218-2064-8, 0-8218-2676-X. x + 405 pp. LCCN 01.50.

Macrae:1994:JNM

[Mac94]  Norman Macrae.  John von Neumann: Mathematik und

Computerforschung-Facetten eines Genies. (German) [John von Neumann: Mathematics and Computer-Research Facets of a Genius]. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 3-7643-2974-2. 349 pp. LCCN ???? Translated by Monika Niehaus-Osterloh.

Macrae:1999:JN

[Mac99a] Norman Macrae. John von Neumann. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-2064-8. x + 406 pp. The scientific genius who pioneered the modern computer, game theory, nuclear deterrence, and much more, Reprint of the 1992 original.

Macrae:1999:JNS

[Mac99b] Norman Macrae. John von Neumann: the scientific genius who pioneered the modern computer, game theory, nuclear deterrence, and much more. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-2064-8. x + 405 pp. LCCN QA29.V66 M334 1999. Originally published in [Mac92b].

Madden:2012:JTD

[Mad12] Niall Madden. John Todd and the development of modern numerical analysis. Irish Mathematical Society Bulletin, 69:11–24, Summer 2012. ISSN 0791-5578. URL http://www.maths.tcd.

ie/pub/ims/bull69/Madden.pdf.

Mahoney:1993:BRBb

[Mah93] Michael S. Mahoney. Book review: John von Neumann and the Origins of Modern Computing by William Aspray. Isis, 84(2):408– 409, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236298.

Mahoney:2011:CMP

[Mah11] Michael S. (Michael Sean) Mahoney. Computing and mathematics at Princeton in the 1950s. In Mahoney and Haigh [MH11], page ?? ISBN 0-674-05568-3. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [CK13] for comments on the author and this book.

Mandelbrot:2017:TGN

[Man17] Benoît Mandelbrot. A touching gesture by von Neumann. Youtube video (2m44s), August 21, 2017.

Marschak:1946:NMN

[Mar46] Jacob Marschak. Neumann’s and Morgenstern’s new approach to static economics. Journal of Political Economy, 54(2):97–115, April 1946. CODEN ???? ISSN 0022-3808 (print), 1537-534X

(electronic). URL https://www.journals.uchicago.edu/doi/
10.1086/256327;   https://www.jstor.org/stable/1825790.

Marlow:1965:UDN

[Mar65] A. R. Marlow. Unified Dirac–von Neumann formulation of quantum mechanics. I. Mathematical theory. Journal of Mathematical Physics, 6(6):919–927, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v6/i6/p919_s1.

Marx:1996:MMG

[Mar96] George Marx. The myth of the Martians and the golden age of Hungarian science. Science & Education (Springer), 5(3):225–234, July 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1996Sc%26Ed..
.5..225M.

Marchal:1998:JNF

[Mar98] Pierre Marchal. John von Neumann: the founding father of artificial life. Artificial Life, 4(3):229–235, Summer 1998. CODEN ARLIEY. ISSN 1064-5462, 1530-9185.

Marton:2006:GEN

[Mar06] Kati Marton. The great escape: nine Jews who fled Hitler and changed the world. Simon and Schuster, New York, NY, USA, 2006. ISBN 0-7432-6115-1 (hardcover), 0-7432-6116-X (paperback). 271 + 16 pp. LCCN DS135.H93 A153 2006. URL http://catdir.

loc.gov/catdir/enhancements/fy0665/2006049162-d.html;
http://catdir.loc.gov/catdir/enhancements/fy0668/2006049162-
s.html;   http://catdir.loc.gov/catdir/enhancements/fy0668/
2006049162-t.html;   http://catdir.loc.gov/catdir/enhancements/
fy0912/2006049162-b.html.

Matolcsi:2002:JNF

[Mat03] Máté Matolcsi. John von Neumann and the foundation of the abstract theory of Hilbert spaces. Mat. Lapok (N.S.), 11(2):26–35 (2006), 2002/03. ISSN 0025-519X.

Macdonald:2000:BRE

[MC00] J. Ross Macdonald and Harvey G. Cragon. Book review: ENIAC: The Triumphs and Tragedies of the World’s First Computer, by Scott McCartney, Walker and Co., New York, 1999. 262 pp. $23.00 hardcover. ISBN 0-8027-1348-3. Physics Today, 53(7):58–59, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

McClelland:1983:RCC

[McC83] William F. McClelland. Report of 10/18/51 conference on checking with von Neumann. Annals of the History of Computing, 5(2):153– 154, April/June 1983. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1983/pdf/a2153.pdf;
 http://www.computer.org/annals/an1983/a2153abs.htm.

McCartney:2001:ETT

[McC01] Scott McCartney. ENIAC: the Triumphs and Tragedies of the World’s First Computer. Berkley Books, New York, NY, USA, 2001. ISBN 0-425-17644-4. vii + 262 pp. LCCN QA76.5.

Montuschi:2023:MCE

[MCP23] Paolo Montuschi, Yuan-Hao Chang, and Vincenzo Piuri. Inmemory computing: The emerging computing topic in the postvon Neumann era. Computer, 56(10):4–6, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mahoney:2011:HC

[MH11] Michael S. (Michael Sean) Mahoney and Thomas Haigh, editors. Histories of computing. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05568-3. 250 pp. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [CK13] for comments on the author and this book.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Mirsky:1975:TIJ

[Mir75] L. Mirsky. A trace inequality of John von Neumann. Monatshefte für Mathematik, 79(4):303–306, 1975. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Mirowski:1992:WWN

[Mir92] Philip Mirowski. What were von Neumann and Morgenstern trying to accomplish? In Weintraub [Wei92], pages 113–147. ISBN 0-8223-1253-0, 99-2403131-8. ISSN 0018-2702. LCCN HB1 .H55 suppl. vol. 24 1992; HB144 .T68 1992.

Mitchell:1980:CSI

[Mit80] Janet A. Mitchell, editor. A community of scholars: the Institute for Advanced Study, faculty and members 1930–1980. The Institute for Advanced Study, Princeton, NJ, USA, 1980. xxiii + 549 pp. LCCN LD2535.I31 I67 1980. With a foreword by Harry Woolf. Biographies and bibliographies of The Institute for Advanced Study faculty and members, in celebration of the 50th anniversary of the Institute.

Mitsch:2022:HSA

[Mit22] Chris Mitsch. Hilbert-style axiomatic completion: On von Neumann and hidden variables in quantum mechanics. Studies in History and Philosophy of Science Part A, 95(??):84–95, October 2022. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368122000978.

Mumford:2009:CIC

[MJ09] Christine L. Mumford and L. C. Jain, editors. Computational Intelligence: Collaboration, Fusion and Emergence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-01798-3 (print), 3-642-01799-1 (e-book). LCCN

Q342 .C66 2009. URL http://link.springer.com/openurl?
genre=book\%26isbn=978-3-642-01798-8.

Moldauer:1972:RNT

[Mol72] P. A. Moldauer. A reinterpretation of von Neumann’s theory of measurement. Foundations of Physics, 2(1):41–47, January 1972. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00708618.

Monahan:1979:ENM

[Mon79] John F. Monahan. Extensions of von Neumann’s method for generating random variables. Mathematics of Computation, 33(147): 1065–1069, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Monastyrskii:2006:JNM

[Mon06] M. I. Monastyrskiĭ. John von Neumann — the mathematician and the man. Istor.-Mat. Issled. (2), 11(46):240–266, 359, 2006. ISBN 5-8037-0349-4.

Morita:2002:CEN

[Mor02] Morita Tsuneo. ’Egy csodálatos elme’ — és Neumann János. (Hungarian) [“A beautiful mind’ — and John von Neumann]. Fizikai Szemle (Budapest), 52(6):177–180, 2002. CODEN FISZA6. ISSN ????

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Metropolis:1950:STV

[MRvN50] N. C. Metropolis, G. Reitwiesner, and J. von Neumann. Statistical treatment of values of first 2,000 decimal digits of e and of π calculated on the ENIAC. Mathematical Tables and Other Aids to Computation, 4(30):109–111, 1950. CODEN MTTCAS. ISSN 0891-6837. Reprinted in [Tau63b, Paper 22].

Miklos:1973:SNJ

[MS73] Miklós Philip and Szentiványi Tibor, editors. Szemelvények Neumann János életéből. (Hungarian) [Selected Passages from the Life of John von Neumann]. Neumann János Számítógéptudományi Társaság, Budapest, Hungary, 1973. 35 pp. LCCN ????

Mermin:2018:HNN

[MS18] N. David Mermin and Rüdiger Schack. Homer nodded: Von Neumann’s surprising oversight. Foundations of Physics, 48(9): 1007–1020, September 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10701-018-0197-5.pdf.

Muhlenbein:2009:CIL

[Müh09] Heinz Mühlenbein. Computational intelligence: The legacy of Alan Turing and John von Neumann. In Mumford and Jain [MJ09], pages 23–43. ISBN 3-642-01798-3 (print), 3-642-01799-1 (e-book). LCCN Q342 .C66 2009. URL http://link.springer.

com/chapter/10.1007/978-3-642-01799-5_2.

Murray:1958:TOS

[Mur58] F. J. Murray. Theory of operators. Part I. Single operators. Bulletin of the American Mathematical Society, 64(3 (part 2)):57–60, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.
bams/1183522371.

Murawski:2004:JNH

[Mur04] Roman Murawski. John von Neumann and Hilbert’s school of foundations of mathematics. Report, Adam Mickiewicz University, Faculty of Mathematics and Computer Science, ul. Umultowska 87, 61614 Poznań, Poland, September 23, 2004. 18 pp. URL http://

www.calculemus.org/studies04/murawski.pdf.

Murray:1936:RO

[MvN36] F. J. Murray and J. von Neumann. On rings of operators. Annals of Mathematics (2), 37(1):116–229, ???? 1936. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 2].

Murray:1937:ROI

[MvN37] F. J. Murray and J. von Neumann. On rings of operators. II. Transactions of the American Mathematical Society, 41(2):208–248, ???? 1937. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). Reprinted in [Tau63a, Paper 3].

Murray:1943:ROI

[MvN43] F. J. Murray and J. von Neumann. On rings of operators. IV. Annals of Mathematics (2), 44:716–808, 1943. CODEN ANMAAH.

ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 5].

Mendelsohn:1977:SPS

[MWW77] Everett Mendelsohn, Peter Weingart, and Richard Whitley, editors. The Social Production of Scientific Knowledge, volume 1 of Sociology of the sciences. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0775-8, 90-277-0776-6 (paperback). vi + 294 pp. LCCN Q175.5 .S6.

Nagy:1987:JNH

[Nag87a] Ferenc Nagy. Neumann János és a “Magyar Titok’, A Dokumentumok Tükrében. (Hungarian) [John von Neumann and the Hungarian Secret: documents in the light]. Országos Müszaki Információs Központ és Könyvtár, Budapest, Hungary, 1987. ISBN 963-592-595-6. 247 pp. LCCN QA29.V66 A4 1987.

Nagy:1987:JAN

[Nag87b] Nagy Ferenc. Új adatok Neumannról (riporter: Kis János). (Hungarian) [New data on von Neumann (reporter: John kis)]. Új impulzus, 3(13):13–14, ???? 1987. CODEN ???? ISSN ????

Nagy:2003:NJK

[Nag03] Nagy Tibor. Neumann János és a kvantumelmélet. (Hungarian) [John von Neumann and quantum theory]. Fizikai Szemle (Budapest), 53(12):419–420, 2003. CODEN FISZA6. ISSN ????

Nagy:1989:JNC

[NAH^{+}89] D. Nagy, W. Aspray, P. Horvath, E. Teller, N. Vonneuman, and E. P. Wigner. John von Neumann — a case-study of scientific creativity — discussion. Annals of the History of Computing, 11 (3):165–169, ???? 1989. CODEN AHCOE5. ISSN 0164-1239.

Nasar:1998:BM

[Nas98a] Sylvia Nasar. A Beautiful Mind. Simon and Schuster, New York, NY, USA, 1998. ISBN 0-684-81906-6 (paperback), 0-684-85370-1 (paperback), 0-7432-2457-4 (paperback). 459 + 8 pp. LCCN QA29.N25 N37 1998.

Nasar:1998:JNP

[Nas98b] Sylvia Nasar. John von Neumann (Princeton, 1948–49). In A Beautiful Mind [Nas98a], pages 79–82. ISBN 0-684-81906-6 (paperback), 0-684-85370-1 (paperback), 0-7432-2457-4 (paperback). LCCN QA29.N25 N37 1998.

NGBS:1988:IDC

[Nat88] National Geographic Book Service. Inventors and discoverers: changing our world. National Geographic Society, Washington, DC, USA, 1988. ISBN 0-87044-751-3, 0-87044-752-1, 0-87044-753-X. 320 pp. LCCN T18 .I57 1988.

Nemeth:1984:DEN

[Ném84] Németh Sándor. Diákkori emlékek Neumann Jánosról. (Hungarian) [Student memories of John von Neumann]. Új tükör, 21(7):21–??, ???? 1984. CODEN ???? ISSN ????

Nemeth:1993:FGV

[Ném93] Németh József. A Fasori Gimnáziumtól a világhírig. (Hungarian) [The grammar school avenue became world famous]. Diakonia, 15 (2):47–52, 1993. CODEN ???? ISSN ????

Neumann:1933:KPH

[Neu33] Neumann János. A koordináta-mérés pontosságának határai az elektron dirac-félé elméletében. (Hungarian) [Über die Grenzen der Koordinatenmessungs-Genauigkeit in der Diracschen Theorie des Elektrons]. (German) [On the limits of coordinate measurement accuracy in the Dirac theory of the electron]. Mat. es Termeszettut … Ertesito, 50(??):366–385, ???? 1933. Reprinted in [Tau61b, Paper 20].

Neumann:1964:SAH

[Neu64] Neumann János. A szamologep es az agy. (Hungarian) [The Computer and the Brain]. Gondolat, Budapest, Hungary, 1964. 130 pp. LCCN ????

Neumann:1972:SAH

[Neu72] Neumann János. A szamologep es az agy. (Hungarian) [The Computer and the Brain]. Gondolat, Budapest, Hungary, second edition, 1972. 133 pp. LCCN ????

Neumann:1989:MMS

[Neu89a] Neumann Miklós. Még monográfia sincs róla. a testvéröccs Neumann Jánosról. (Hungarian) [I have no monograph. The ???? of John von Neumann]. Népszava, 117(203):9, 1989. CODEN ???? ISSN ????

Neumann:1989:TNE

[Neu89b] Neumann Miklós. A tudós nem élhet elefántcsonttoronyban. (Hungarian) [The scientist cannot live in an ivory tower]. Magyar hírlap, 22(199):8, 1989. CODEN ???? ISSN ????

Neumann:1990:TTF

[Neu90] Neumann Miklós. A tudósok társadalmi felel ssége. Neumann János üzenete. (Hungarian) [The social responsibility of scientists. John von Neumann’s message]. Magyar hírlap, 23(17):3–??, 1990. CO-DEN ???? ISSN ????

Neumann:2003:ETH

[Neu03] Neumann Nicolas A. Emlékeim testvéremről. (Hungarian) [Memories of my brother]. Természet világa, 134(3):22–30, 2003. CODEN ???? ISSN ????

Neumann:2006:SAH

[Neu06] Neumann János. A szamologep es az agy. (Hungarian) [The Computer and the Brain]. NetAcademia Oktatóközp., Budapest, Hungary, second edition, 2006. ISBN 963-06-1523-1. iii + 90 pp. LCCN ????

Nagy:1989:NOC

[NHN89] Dénes Nagy, Péter Horvath, and Ferenc Nagy. The von Neumann– Ortvay connection. Annals of the History of Computing, 11 (3):183–188, July/September 1989. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1989/
pdf/a3183.pdf;   http://www.computer.org/annals/an1989/
a3183abs.htm.

Northrup:1998:ACP

[Nor98] Mary Northrup. American computer pioneers. Collective biographies. Enslow Publishers, Springfield, NJ, USA, 1998. ISBN 0-7660-1053-8. 112 pp. LCCN QA76.17 .N68 1998.

Nowak:1966:JND

[NPP66] Amrak Nowak, Patricia Powell, and Jerry Pantzer. John Von Neumann, a documentary. Video cassette (60 minutes), 1966. Uses rare photographs and films to present a biography of the mathematician John Von Neumann. Includes personal recollections of friends and colleagues as well as detailed descriptions and discussions of his contributions in various fields of science.

Oxtoby:1958:JN

[OPP58] J. C. Oxtoby, B. J. Pettis, and G. B. Price. John von Neumann 1903–1957. Bulletin of the American Mathematical Society, 64(3 (part 2)):1–129, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). With a portrait and a biography.

Oxtoby:1988:JN

[OPP88] J. C. Oxtoby, B. J. Pettis, and G. B. Price, editors. John von Neumann, 1903–1957. American Mathematical Society, Providence, RI, USA, 1988. ISBN 0-8218-0021-3. vi + 129 pp. Reprint of Bull. Amer. Math. Soc. 64 (1958), no. 3, Part 2.

Palm:2013:BTP

[PA13] Günther Palm and Ad Aertsen, editors. Brain theory: proceedings of the first Trieste meeting on brain theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-70913-3. LCCN ????

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/toc/fy02/99046603.html.

Palugyai:1987:JVT

[Pal87] Palugyai István. Jelenség volt a tudományban. Neumann János emlékezete. (Hungarian) [It was a phenomenon in science: John von Neumann’s memory]. Magyar hírlap, 20(33):4–??, ???? 1987. CODEN ???? ISSN ????

Pallo:2003:NJH

[Pal03a] Palló Gábor. Neumann János (1903–1957). (Hungarian) [John von Neumann (1903–1957)]. Fizikai Szemle (Budapest), 53(12):413– 414, 2003. CODEN FISZA6. ISSN ????

Palugyai:2003:NJO

[Pal03b] Palugyai István. Neumann János öröksége. (Hungarian) [John von Neumann’s legacy]. Népszabadság, 12(3):8–??, 2003. CODEN ???? ISSN ????

Panday:1992:BRJ

[Pan92] Rajeev Panday. Book review: John von Neumann and the Origins of Modern Computing by William Aspray (MIT Press 1990, 376 pages, 40 illustrations, $35.00). ACM SIGCAS Computers and Society, 22(1–4):29, October 1992. CODEN CMSCD3. ISSN 0095-2737 (print), 2167-3055 (electronic).

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005002173-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/
2005002173-b.html;   http://www.loc.gov/catdir/toc/ecip056/
2005002173.html.

Pelaez:1999:SPC

[Pel99] Eloína Peláez. The stored-program computer: Two conceptions. Social Studies of Science, 29(3):359–389, June 1, 1999. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/030631299029003002.

Pellicer:2003:CJN

[Pel03] Manuel López Pellicer, editor. Centenario de John von Neumann: (1903–1957). (Spanish) [The Centenary of John von Neumann: (1903–1957)], volume 692 of Revista Arbor: Ciencia pensamiento y cultura. CSIC, Madrid, Spain, August 2003. ISBN ???? ISSN 0518-2654. xxiii + 1369–1496 pp. LCCN ???? URL http://www.

csic.es/arbor/sumarios_2003.html.

Peres:1992:INP

[Per92] Yuval Peres. Iterating von Neumann’s procedure for extracting random bits. Annals of Statistics, 20(1):590–197, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://www.jstor.org/stable/2242181.

Pesavento:1995:INS

[Pes95] Umberto Pesavento. An implementation of von Neumann’s self-reproducing machine. Artificial Life, 2(4):337–354, July 1995. CODEN ARLIEY. ISSN 1064-5462 (print), 1530-9185

(electronic). URL https://direct.mit.edu/artl/article/2/4/
337/2260/An-Implementation-of-von-Neumann-s-Self.

Petz:2001:ENN

[Pet01] Dénes Petz. Entropy, von Neumann and the von Neumann entropy. In Rédei and Stöltzner [RS01b], pages 83–96. ISBN 0-7923-6812-6. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN ????

Petz:2005:NJT

[Pet05] Petz Dénes. Neumann János tudományos öröksége. (Hungarian) [John von Neumann’s scientific heritage]. Debreceni szemle, 13(1): 15, 2005. CODEN ???? ISSN ????

Phillips:1988:SHM

[Phi88] Esther R. Phillips, editor. Studies in the history of mathematics, volume 26 of Studies in Mathematics. Mathematical Association of America, Washington, DC, USA, 1988. ISBN 0-88385-128-8. 308 pp. LCCN ????

Pinch:1977:WDP

[Pin77] Trevor J. Pinch. What does a proof do if it does not prove? A study of the social conditions and metaphysical divisions leading to David Bohm and John von Neumann failing to communicate in quantum physics. In Mendelsohn et al. [MWW77], pages 171–216. ISBN 90-277-0775-8, 90-277-0776-6 (paperback). LCCN Q175.5 .S6.

Pronold:2022:RBT

[PJW^{+}22] J. Pronold, J. Jordan, B. J. N. Wylie, I. Kitayama, M. Diesmann, and S. Kunkel. Routing brain traffic through the von Neumann bottleneck: Efficient cache usage in spiking neural network simulation code on general purpose computers. Parallel Computing, 113 (??):??, October 2022. CODEN PACOEJ. ISSN 0167-8191 (print),

1872-7336 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167819122000461.

Post:1995:QRE

[Pos95a] Evert Jan Post. Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics, volume 181 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346. xiv + 320 pp. LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL

http://books.google.com/books?id=lK3aAAAAMAAJ;   http:/
/catalog.hathitrust.org/api/volumes/oclc/32469387.html;
 http://catdir.loc.gov/catdir/enhancements/fy0823/95017463d.
html;   http://catdir.loc.gov/catdir/enhancements/fy0823/
95017463t.html;   https://link.springer.com/book/10.1007/
978-94-015-8410-4.

Post:1995:NPE

[Pos95b] Evert Jan Post. Von Neumann, Popper–EPR, Bohm, Bell, Aspect. In Quantum Reprogramming: Ensembles and Single Systems: a Two-Tier Approach to Quantum Mechanics [Pos95a], chapter 5, pages 74–83. ISBN 0-7923-3565-1 (hardcover), 94-015-8410-9. ISSN 0068-0346. LCCN Q174 .B67 vol. 181; Q174; Q174.12; QC174.12 P675 1995. URL https://link.springer.

com/chapter/10.1007/978-94-015-8410-4_5.

Poundstone:1993:PDJ

[Pou93] William Poundstone. Prisoner’s dilemma: John von Neumann, game theory, and the puzzle of the bomb. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-286162-X (paperback), 0-385-41580-X. xi + 294 + 8 pp. LCCN QA29.V66 P68 1993. US$12.95 (US$15.95 Can.).

Poundstone:2003:DDP

[Pou03] William Poundstone. Le dilemme du prisonnier: Von Neumann, la théorie des jeux et la bombe. (French) [Prisoner’s dilemma: John von Neumann, game theory, and the puzzle of the bomb]. Cassini Éditeurs, 75001 Paris, France, 2003. ISBN 2-84225-143-1 (paperback), 2-84225-046-X. viii + 348 pp. LCCN QA29 .V66 P6814

2003. 15 EUR. URL http://www.cassini.fr/#dilemme_du_

prisonnier. Translated by Oristelle Bonis.

Penney:1944:SHB

[PPvNF44] W. G. Penney, Rudolf E. Peierls, John von Neumann, and Klaus Fuchs. Shock hydrodynamics and blast waves. Report AECD 2860, U.S. Atomic Energy Commission, Technical Information Division, Oak Ridge, TN, USA, October 28, 1944. 60 pp. URL http://hdl.

handle.net/2027/mdp.39015086444083. Notes written by J. O.

Hirschfelder. Declassified 22 April 1949. This report is based on LADC-549.

Priestley:2010:SOM

[Pri10] Mark Priestley. A Science of Operations: Machines, Logic and the Invention of Programming. History of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-554-4 (hardcover), 1-84882-555-2 (e-book). ix + 341 pp. LCCN QA76.6 .P737 2010.

Randell:1972:ATO

[Ran72] Brian Randell. On Alan Turing and the origins of digital computers. Technical report 33, University of Newcastle upon Tyne, Computing Laboratory, Newcastle upon Tyne, UK, 1972. 36 pp.

Randell:1974:BRT

[Ran74] Brian Randell. Book review: Technology The Computer from Pascal to von Neumann. By Herman H. Goldstine. Princeton: Princeton University Press, and London: Oxford University Press, 1973. Pp. xii + 378. £6.25. British Journal for the History of Science, 7(3):289–290, November 1974. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4025539.

Randell:1982:ODC

[Ran82] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1982. ISBN 0-387-11319-3, 3-540-11319-3. xvi + 580 pp. LCCN TK7885.A5 O741 1982.

Rashid:2007:JNS

[Ras07] Salim Rashid. John von Neumann and scientific method. Journal of the History of Ideas, 68(3):501–527, July 2007. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL http:/

/www.jstor.org/stable/30140935;   https://muse.jhu.edu/
pub/56/article/219822.

Rechenberg:2007:BRB

[Rec07] Helmut Rechenberg. Book review: John von Neumann: Selected Letters, vol. 27, Miklós Rédei (Ed.), in: History of Mathematics. American Mathematical Society, London Mathematical Society, Providence, RI (2005), ISBN 0-8218-3776-1. Historia Mathematica, 34(3):354–356, August 2007. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S031508600600084X.

Redei:1959:JNW

[Réd59] László Rédei. John von Neumann’s work in algebra and number theory. Mat. Lapok, 10:226–230, 1959. ISSN 0025-519X.

Redei:1996:WJN

[Réd96] Miklós Rédei. Why John von Neumann did not like the Hilbert Space formalism of quantum mechanics (and what he liked instead). Studies in History and Philosophy of Modern Physics, 27(4):493–510 (1997), December 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000172.

Redei:1999:UPM

[Réd99] Miklós Rédei. “Unsolved problems in mathematics”: J. von Neumann’s address to the International Congress of Mathematicians, Amsterdam, September 2–9, 1954. The Mathematical Intelligencer, 21(4):7–12, 1999. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Redei:2001:NCQ

[Réd01] Miklós Rédei. Von Neumann’s concept of quantum logic and quantum probability. In Rédei and Stöltzner [RS01b], pages 153–172. ISBN 0-7923-6812-6. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN ????

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Regis:1987:WGE

[Reg87] Edward Regis. Who got Einstein’s office?: eccentricity and genius at the Institute for Advanced Study. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-12065-8. xv + 316 pp. LCCN Q180.U5 R35 1987. US$17.95.

Reitwiesner:1950:EDM

[Rei50] George W. Reitwiesner. An ENIAC determination of π and e to more than 2000 decimal places. Mathematical Tables and Other Aids to Computation, 4(29):11–15, January 1950. CODEN MTTCAS. ISSN 0891-6837.

Rellstab:1992:NIC

[Rel92] Urs Rellstab. New insights into the collaboration between John von Neumann and Oskar Morgenstern on the Theory of games and economic behavior. In Toward a history of game theory, pages 77– 93. Duke University Press, Durham, NC, USA, 1992.

Revesz:1983:JNR

[Rév83] Győrgy Révész. John von Neumann und der Rechner. (German) [John von Neumann and the computer]. In Life and work of John von Neumann, pages 99–113. Bibliographisches Institut, Mannheim, Germany; Wien, Austria; Zürich, Switzerland, 1983.

Rojas:2000:FCH

[RH00] Raúl Rojas and Ulf Hashagen, editors. The first computers: history and architectures. History of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-18197-5 (hardcover), 0-585-35535-5 (electronic). xii + 457 pp. LCCN QA76.17 .F57 2000.

RileyKnoedler:2021:UMJ

[RKH^{+}21] Molly Riley Knoedler, Julianna C. Kostas, Caroline Mary Hogan, Harper Kerkhoff, and Chad M. Topaz. An unpublished manuscript of John von Neumann on shock waves in boostered detonations: historical context and mathematical analysis. Archive for History of Exact Sciences, 75(1):83–108, January 2021. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-020-00258-9.pdf.

Rowe:1989:HMM

[RMK94] David E. Rowe, John McCleary, and Eberhard Knobloch, editors. The history of modern mathematics: proceedings of the Symposium on the History of Modern Mathematics, Vassar College, Poughkeepsie, New York, June 20–24, 1989. Academic Press, New York, NY, USA, 1989–1994. ISBN 0-12-599661-6 (vol. 1), 0-12-599662-4 (vol. 2), 0-12-599663-2 (vol. 3). LCCN QA21 .S98 1989. URL http:

//www.loc.gov/catdir/description/els032/89017766.html;
 http://www.loc.gov/catdir/toc/els031/89017766.html.

Robinson:2012:SED

[Rob12] Andrew Robinson, editor. The scientists: an epic of discovery. Thames and Hudson, New York, NY, USA, 2012. ISBN 0-500-25191-6. 304 pp. LCCN Q141 .S3712 2012.

Rojas:1994:WIC

[Roj94] Raúl Rojas. Who invented the computer? The debate from the viewpoint of computer architecture. Proceedings of symposia in applied mathematics, 48:361–365, 1994. CODEN PSYMA7. ISSN

0160-7634. URL http://page.mi.fu-berlin.de/rojas/1993/
Who_invented_the_computer.pdf. Proceedings of Mathematics

of Computation 1943–1993: a half-century of computational mathematics (Vancouver, BC, 1993).

Redei:2001:JNF

[RS01a] Miklós Rédei and Michael Stöltzner, editors. John von Neumann and the foundations of quantum physics, volume 8 of Vienna Circle Institute Yearbook. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6812-6. ix + 371 pp. LCCN QC173.98 .J65 2001. Including papers from the workshop held at Loránd Eötvös University, Budapest, February 1999.

Redei:2001:AGR

[RS01b] Miklós Rédei and Michael Stöltzner, editors. John von Neumann and the foundations of quantum physics (Budapest, 1999), volume 8 of Vienna Circle Institute yearbook. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6812-6. ISSN 0929-6328 (print), 2215-1818 (electronic). ix + 371 pp. LCCN ????

Rutkovszky:2003:NJI

[Rut03] Rutkovszky Edéné. Neumann János az interneten. (Hungarian) [John von Neumann on the Internet]. Természet világa, 134(3): 70–??, 2003. CODEN ???? ISSN ????

Richtmyer:1947:SMN

[RUvN47] Robert D. Richtmyer, Stanisław Ulam, and John von Neumann. Statistical methods in neutron diffusion. Technical Report LAMS-551, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April 9, 1947. 22 pp. Republished in typeset form in [Hur85].

Richtmeyer:1985:SMN

[RvN85] R. D. Richtmyer and J. von Neumann. Statistical methods in neutron diffusion. Annals of the History of Computing, 7(2):148–155, April/June 1985. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1985/
pdf/a2141.pdf;   http://www.computer.org/annals/an1985/

a2141abs.htm. Reprint of report with letters of March 11, 1947 and April 2, 1947.

Redei:2012:HIP

[RW12] Miklós Rédei and Charlotte Werndl. On the history of the isomorphism problem of dynamical systems with special regard to von Neumann’s contribution. Archive for History of Exact Sciences, 66(1):71–93, January 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=66&
issue=1&spage=71.

Samuelson:1992:ETN

[Sam92] Paul A. Samuelson. Economics and thermodynamics: von Neumann’s problematic conjecture. In Rational interaction: essays in honor of John C. Harsanyi [Sel92], pages 377–389. ISBN 0-387-55067-4 (New York), 3-540-55067-4 (Berlin), 3-662-09664-1. LCCN

QA269 .R375 1992. URL http://link.springer.com/content/
pdf/10.1007/978-3-662-09664-2_22.pdf.

Schnelle:1988:TNN

[Sch88] Helmut Schnelle. Turing naturalized: von Neumann’s unfinished project. In Herken [Her88], pages 539–559. ISBN 0-19-853741-7. LCCN QA267 .U55 1988.

Scheibe:1991:JVN

[Sch91] Erhard Scheibe. J. v. Neumanns und J. S. Bells Theorem. Ein Vergleich. (German) [J. v. Neumann’s and J. S. Bell’s Theorem. A comparison]. Philosophia Naturalis, 28(1):35–53, 1991. CODEN ???? ISSN 0031-8027. 2nd FEST Philosophy and Physics Workshop (Heidelberg, 1989).

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv– xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Selten:1992:RIE

[Sel92] Reinhard Selten. Rational interaction: essays in honor of John C. Harsanyi. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55067-4 (New York), 3-540-55067-4 (Berlin), 3-662-09664-1. ix + 438 pp. LCCN QA269 .R375 1992. URL http://catdir.loc.

gov/catdir/enhancements/fy0815/91045698-d.html;   http:
//www.gbv.de/dms/bowker/toc/9780387550671.pdf;   http:/
/www.zentralblatt-math.org/zmath/en/search/?an=0829.00011.

Shannon:1958:NCA

[Sha58] Claude E. Shannon. Von Neumann’s contributions to automata theory. Bulletin of the American Mathematical Society, 64(3 (part 2)):123–129, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183522376.

Shelburne:2012:ED

[She12] Brian Shelburne. The ENIAC’s 1949 determination of π. IEEE Annals of the History of Computing, 99(PrePrints), 2012. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shurkin:1984:EMH

[Shu84a] Joel N. Shurkin. Engines of the mind: a history of the computer. W. W. Norton & Co., New York, NY, USA, 1984. ISBN 0-393-01804-0. 352 pp. LCCN QA76.17 .S49 1984.

Shurkin:1984:N

[Shu84b] Joel N. Shurkin. Von Neumann. In Engines of the mind: a history of the computer [Shu84a], chapter 7, pages 173–208. ISBN 0-393-01804-0. LCCN QA76.17 .S49 1984.

Shurkin:1985:EMH

[Shu85a] Joel N. Shurkin. Engines of the mind: a history of the computer. Washington Square Press, New York, NY, USA, revised edition, 1985. ISBN 0-671-60036-2. xiii + 335 pp. LCCN QA76.17 .S49 1985.

Shurkin:1985:N

[Shu85b] Joel N. Shurkin. Von Neumann. In Engines of the mind: a history of the computer [Shu85a], chapter 7, page ?? ISBN 0-671-60036-2. LCCN QA76.17 .S49 1985.

Shurkin:1996:EME

[Shu96a] Joel N. Shurkin. Engines of the mind: the evolution of the computer from mainframes to microprocessors. W. W. Norton & Co., New York, NY, USA, updated edition, 1996. ISBN 0-393-31471-5. 363 pp. LCCN QA76.17 .S49 1996.

Shurkin:1996:N

[Shu96b] Joel N. Shurkin. Von Neumann. In Engines of the mind: the evolution of the computer from mainframes to microprocessors [Shu96a], chapter 7, pages 173–208. ISBN 0-393-31471-5. LCCN QA76.17 .S49 1996.

Szentivanyi:1979:NJE

[SI^{+}79] Szentiványi Tibor, Ádám András István, et al., editors. Neumann János élete és munkássága: a különböz tudományterületeken elért eredményeinek összefoglaló áttekintése. (Hungarian) [The life and works of John von Neumann]. Számok Soksz., Budapest, Hungary, 1979. ISBN 963-8431-19-9. 175 pp. LCCN QA7 .N48.

Silagadze:2023:EQM

[Sil23] Zurab K. Silagadze. Evading quantum mechanics à la Sudarshan: Quantum-mechanics-free subsystem as a realization of Koopman– von Neumann mechanics. Foundations of Physics, 53(6):??, December 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-023-00734-6.

Simonits:2003:NJJ

[Sim03] Simonits András. Neumann János és a játékelmélet. (Hungarian) [John von Neumann and game theory]. Természet világa, 134(3): 56–60, 2003. CODEN ???? ISSN ????

Sipos:2005:HEN

[Sip05] Sipos László. Hatvan éves Neumann János korszakalkotó programja. (Hungarian) [John von Neumann’s sixty-year-old groundbreaking program]. Mérnök újság, ??(??):??, December 2005. CO-DEN ???? ISSN ????

Slater:1987:WIC

[Sla87] Robert Slater. Who invented the computer? Computers in Physics, 1(1):44–49, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903432.

Shannon:1956:AS

[SM56] C. E. Shannon and J. McCarthy. Automata Studies, volume 34 of Annals of Mathematics Studies. Princeton University Press, Princeton, NJ, 1956. CODEN ANMAAH. ISSN 0066-2313.

Smagorinsky:1983:BNW

[Sma83] Joseph Smagorinsky. The beginnings of numerical weather prediction and general circulation modeling: Early recollections. Advances in Geophysics, 25:3–37, 1983. CODEN AD-

GOAR. ISSN 0065-2687. URL http://www.sciencedirect.com/
science/article/pii/S0065268708601703. Theory of Climate:

Proceedings of a Symposium Commemorating the Two-Hundredth Anniversary of the Academy of Sciences of Lisbon.

Smithies:1959:JN

[Smi59] F. Smithies. John von Neumann. Journal of the London Mathematical Society, 34:373–384, 1959. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Smiley:2010:MWI

[Smi10] Jane Smiley. The man who invented the computer: the biography of John Atanasoff, digital pioneer. Doubleday, New York, NY, USA, 2010. ISBN 0-385-52713-6, 0-385-53372-1 (e-book), 1-299-11995-6 (e-book). 246 + 8 pp. LCCN QA76.2.A75 S64 2010. US$25.95.

Szokefalvi-Nagy:1957:JNW

[SN57] Béla Szökefalvi-Nagy. J. von Neumann’s work in the theory of operators. Mat. Lapok, 8:185–210, 1957. ISSN 0025-519X.

Sterbenz:1974:FPC

[Ste74] Pat H. Sterbenz. Floating Point Computation. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-322495-3. xiv + 316 pp. LCCN QA76.8.I12 S771 1974.

Stern:1980:JNI

[Ste80] Nancy Stern. John von Neumann’s influence on electronic digital computing, 1944–1946. Annals of the History of Computing, 2 (4):349–362, October/December 1980. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1980/
pdf/a4349.pdf;   http://www.computer.org/annals/an1980/
a4349abs.htm.

Stern:1981:EUA

[Ste81] Nancy B. Stern. From ENIAC to UNIVAC: an appraisal of the Eckert–Mauchly computers. Digital Press history of computing series. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1981. ISBN 0-932376-14-2. ix + 286 pp. LCCN QA76.5 .S775 1981.

Stern:1982:JNN

[Ste82] Nancy Stern. John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death by Steve J. Heims (review). Technology and Culture, 23(1):108–110, January 1982. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/891121/
pdf.

Stone:1983:BRJ

[Sto83] Marshall H. Stone. Book review: John von Neumann and Norbert Wiener, from mathematics to the technologies of life and death. Bulletin of the American Mathematical Society (new series), 8(2): 395–399, 1983. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Stoltzner:1999:WJN

[Stö99] Michael Stöltzner. What John von Neumann thought of the Bohm interpretation. In Greenberger et al. [GRZ99], pages 257–262. ISBN 0-7923-6063-X. ISSN 0929-6328. LCCN QC174.13 .E65 1999.

Stoltzner:2004:OOA

[Stö04] Michael Stöltzner. On optimism and opportunism in applied mathematics: Mark Wilson meets John von Neumann on mathematical ontology. Erkenntnis, 60(1):121–143, 2004. CODEN ERKEDQ. ISSN 0165-0106 (print), 1572-8420 (electronic).

Stoltzner:2007:NGJ

[Stö07] Michael Stöltzner. A new glimpse of John von Neumann’s thought laboratory. Studies in History and Philosophy of Modern Physics, 38(4):938–947, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219807000196.

Strauss:1963:MDa

[Str63a] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Strauss:1963:N

[Str63b] Lewis L. Strauss. Von Neumann. In Men and decisions [Str63a], pages 231–236. LCCN E741 .S78.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Schatten:1946:CSLa

[SvN46a] Robert Schatten and John von Neumann. The cross-space of linear transformations, II. Bulletin of the American Mathematical Society, 52(??):67–??, ???? 1946. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Schatten:1946:CSLb

[SvN46b] Robert Schatten and John von Neumann. The cross-space of linear transformations. II. Annals of Mathematics, 47:608–630, 1946. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 29].

Schatten:1948:CSLa

[SvN48a] Robert Schatten and John von Neumann. The cross-space of linear transformations, III. Bulletin of the American Mathematical Society, 54(??):637–??, ???? 1948. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Schatten:1948:CSLb

[SvN48b] Robert Schatten and John von Neumann. The cross-space of linear transformations. III. Annals of Mathematics, 49:557–582, 1948. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 30].

Segal:1950:TUR

[SvN50] I. E. Segal and John von Neumann. A theorem on unitary representations of semisimple Lie groups. Annals of Mathematics, 52: 509–517, 1950. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 25].

Salzman:1989:GSL

[SvNG89] David Salzman, John von Neumann, and Jack Grimes. Graphics superworkstations and the last hurrah. IEEE Computer Graphics and Applications, 9(4):27–29, July 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stern:1980:CQD

[SW80] Nancy Stern and Maurice V. Wilkes. Comments, queries, and debate: Who “Invented” the first electronic digital computer?; Mauchly’s position on von Neumann’s role in drafting the EDVAC Report. Annals of the History of Computing, 2(4):375–377, October/December 1980. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1980/pdf/a4375.pdf;
 http://www.computer.org/annals/an1980/a4375abs.htm.

Swartzlander:1976:CDD

[Swa76] Earl E. Swartzlander, Jr. Computer design development: principal papers. Hayden Book Co., Rochelle Park, NJ, USA, 1976. ISBN 0-8104-5988-4. 310 pp. LCCN QA76.5 .C612565.

Szabo:2003:NJS

[Sza03a] Szabó Péter Gábor. Neumann János és Szeged. (Hungarian) [John von Neumann and Szeged]. Természet világa, 134(3):48–51, 2003. CODEN ???? ISSN ????

Szabo:2003:NJH

[Sza03b] Szabó Tímea. Neumann János 1903–1957. (Hungarian) [John von Neumann 1903–1957]. A fizika tanítása, 11(4):25–28, 2003. CO-DEN ???? ISSN ????

Szasz:2003:MHM

[Szá03c] Szász Domokos. A matematikus. (Hungarian) [the mathematician]. Természet világa, 134(3):3–7, 2003. CODEN ???? ISSN ????

Szasz:2011:JNM

[Szá11] Domokos Szász. John von Neumann, the mathematician. The Mathematical Intelligencer, 33(2):42–51, 2011. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Szentivanyi:1976:SAE

[Sze76] Szentiványi Tibor. A számítógép-agyú ember. Neumann János élete és úttör munkássága. (Hungarian) [The computer-minded people. John von Neumann and his ground-breaking work]. Népszava, 106(303):10–??, ???? 1976. CODEN ???? ISSN ????

Szekely:1978:ESN

[Szé78] Székely J. Gábor. 75 éve született Neumann János. (Hungarian) [John von Neumann was born 75 years ago]. Magyar nemzet, 34 (299):8–??, ???? 1978. CODEN ???? ISSN ????

Szelezsan:1983:JNR

[Sze83a] János Szelezsán. John von Neumann, ,,der Rechentechniker“(sein Werk auf dem Gebiet der numerischen Methoden). (German) [John von Neumann, “the computer technician” (his work in the field of numerical methods)]. In Life and work of John von Neumann, pages 115–151. Bibliographisches Institut, Mannheim, Germany; Wien, Austria; Zürich, Switzerland, 1983.

Szentgyorgyi:1983:JST

[Sze83b] Szentgyörgyi Zsuzsa. A játékelmélett l a számítógép-tudományig. Neumann Jánosról, születésének 80. évfordulóján. (Hungarian) [Game theory in computer science. John von Neumann, the 80th anniversary of his birth]. Népszabadság, 41(304):6–??, ???? 1983. CODEN ???? ISSN ????

Szentgyorgyi:1983:SAN

[Sze83c] Szentgyörgyi Zsuzsa. A szellem aranycsinálója — Neumann János (1903–1957). (Hungarian) [The ???? spirit: John von Neumann (1903–1957)]. Nyelvünk és kultúránk, 53(??):72–78, ???? 1983. CODEN ???? ISSN ????

Szelezsan:1987:NJM

[Sze87] Szelezsán János. Neumann János munkássága a numerikus módszerek területén. (Hungarian) [John von Neumann’s work in the field of numerical methods]. Információ, elektronika, 22(3): 143–147, ???? 1987. CODEN ???? ISSN ????

Szenassy:1988:NJE

[Szé88] Szénássy Barna. Neumann János életének “els félideje”. levelek fejér lipót hagyatékából. (Hungarian) [John von Neumann “the lifetime of the first half.” the legacy of lipot fejér (??)]. Természet világa, 119(8):352–356, 1988. CODEN ???? ISSN ????

Szentgyorgyi:1997:NEH

[Sze97] Szentgyörgyi Zsuzsa. Negyven éve halt meg Neumann János. (Hungarian) [he died forty years ago: John von Neumann]. Magyar hírlap, 2(06):11–??, 1997. CODEN ???? ISSN ????

Szelezsan:2003:NJE

[Sze03a] Szelezsán János. Neumann János az első, számítógépet alkalmazó fizikus. (Hungarian) [John von Neumann, the first physicist to use a computer]. Fizikai Szemle (Budapest), 53(12):425–429, 2003. CO-DEN FISZA6. ISSN ????

Szenassy:2003:NJE

[Szé03b] Szénássy Barna. Neumann János életének “első félideje”: Levelek fejér lipót hagyatékából. (Hungarian) [John von Neumann’s life “in the first half time”: Letters from the legacy of Leopold Fejér]. Természet világa, 134(3):43–47, 2003. CODEN ???? ISSN ????

Tarjan:1958:WJN

[Tar58] Rezső Tarján. The work of J. von Neumann concerning electronic computing machines. Mat. Lapok, 9:6–18, 1958. ISSN 0025-519X.

Tarjan:1972:JNT

[Tar72] Rezső Tarján. John von Neumann and the theory of automata. Informác.-Elektron., 7:83–88, 1972. ISSN 0019-9753.

Tarwater:1977:BTA

[Tar77] Dalton Tarwater, editor. The bicentennial tribute to American mathematics 1776–1976. Papers presented at the fifty-nineth annual meeting of the Mathematical Association of America commemorating the nation’s bicentennial. Mathematical Association of America, Washington, DC, USA, 1977. URL https://maa.org/

sites/default/files/pdf/ABOUTMAA/History/Bicentennial_
Tribute.pdf.

Taub:1960:BRB

[Tau60] A. H. Taub. Book review: The Computer and the Brain by John von Neumann. Isis, 51(1):94–96, March 1960. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227613.

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon Press, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon Press, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon Press, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon Press, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon Press, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Teller:1957:JND

[Tel57] Edward Teller. John von Neumann [28 December 1903–8 February 1957]. Bulletin of the Atomic Scientists, 13(4):150–151, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teleman:1970:RRJ

[Tel70] Silviu Teleman. On the regular rings of John von Neumann. Rev. Roumaine Math. Pures Appl., 15:735–742, 1970. ISSN 0035-3965.

Thompson:1989:JN

[Tho89] Gerald L. Thompson. John von Neumann. In Eatwell et al. [EMN89], pages 242–252. ISBN 0-333-49536-5, 0-333-49537-3 (paperback). LCCN HB144 .N48 1989.

Tucker:1985:IHG

[TN85] Albert Tucker and Frederik Nebeker. Interview with Herman Goldstine. The Princeton Mathematics Community in the 1930s Transcript Number 15 (PMC15), March 22, 1985. URL https://web.

math.princeton.edu/oral-history/c14.pdf.

Todd:1974:JNN

[Tod74] John Todd. John von Neumann and the National Accounting Machine. SIAM Review, 16(4):526–530, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://

www.jstor.org/stable/2028693.

Tompkins:1957:JN

[Tom57] C. B. Tompkins. John von Neumann, 1903–1957. Mathematical Tables and Other Aids to Computation, 11:127–128, 1957. CODEN MTTCAS. ISSN 0891-6837.

Toszegi:2004:JND

[Tós04] Zsuzsanna Tószegi. The John von Neumann Digital Library (www.neumann-haz.hu). International Journal on Digital Libraries, 4(1):13–18, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL https://link.springer.

com/article/10.1007/s00799-003-0077-1.

Toth:1993:ITH

[Tót93] Maria Tóth, editor. Information technology in Hungary: marking the 25th anniversary of the John von Neumann Society for Computing Sciences. NJSZT, Budapest, 1993. ISBN 963-8431-79-2. 63 pp. LCCN ????

Tropp:1975:CPN

[Tro75] Henry S. Tropp. The Computer from Pascal to von Neumann by Herman H. Goldstine (review). Technology and Culture, 16 (4):660–662, October 1975. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/891621/pdf.

Tropp:1976:BRB

[Tro76] Henry S. Tropp. Book review: The Computer from Pascal to von Neumann by Herman H. Goldstine. Isis, 67(2):295–297, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230938.

Turing:1960:KMD

[Tur60a] A. M. Turing. Kann eine Maschine denken? [Mozet li masina myslit’?] Mit einem Anhang von J. von Neumann: Allgemeine und logische Theorie der Automaten. [Obscaja i logiceskaja teorija avtomatov.]. (German) [Can a machine think? With an appendix by J. von Neumann: General and Logical Theory of Automata]. Staatsverlag für physikalisch-mathematische Literatur, Moscow, Russia, 1960. 112 pp.

Turing:1960:MLM

[Tur60b] Alan M. Turing. Mozhet li mashina myslit? With a supplement by J. von Neumann. Translated from the English by Ju. A. Danilov; edited and with a foreword by S. A. Janovskaja. Gosudarstv. Izdat. Fiz.-Mat. Lit., Moscow, 1960. 112 pp.

Turing:1999:MLM

[TvN99] A. Turing and J. von Neumann. Mozhet li mashina myslit’?. (Russian) [Can the machine think?]. Nauchno-Izdatel’skij Tsentr “Regulyarnaya i Khaoticheskaya Dinamika”, Izhevsk, Russia, 1999. 104 pp.

Taub:1934:DEP

[TVvN34] A. H. Taub, O. Veblen, and J. von Neumann. The Dirac equation in projective relativity. Proceedings of the National Academy of Sciences of the United States of America, 20(??):383–388, ???? 1934. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau61b, Paper 24].

Tropp:1987:AIM

[TWA^{+}87] Henry S. Tropp, Michael R. Williams, Isaac Auerbach, Irving S. Reed, John von Neumann, and Bobbi Mapstone. Anecdotes: Introduction; the MADDIDA; 1982 Claude E. Shannon Lecture: Application transforms to coding and related topics; the von Neumann letter; interview with Irving S. Reed. Annals of the History of Computing, 9(3/4):357–369, July/September 1987. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/books/
an1987/pdf/a3357.pdf;   http://www.computer.org/annals/
an1987/a3357abs.htm.

Tweedale:1993:BRW

[Twe93] Geoffrey Tweedale. Book review: William Poundstone, Prisoner’s Dilemma: John von Neumann, Game Theory, and the Puzzle of the Bomb. Oxford: Oxford University Press, 1992. Pp. xi + 290. ISBN 0-19-286162-X. £7.99 (paperback edition). British Journal for the History of Science, 26(3):375–376, September 1993. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4027424.

Ulam:1968:JN

[UKTS68] S. Ulam, H. W. Kuhn, A. W. Tucker, and Claude E. Shannon. John von Neumann, 1903–1957. Perspectives in American History, 2(??):235–269, ???? 1968. CODEN ???? ISSN 0079-0990.

Ulam:1969:JN

[UKTS69] Stanisław M. Ulam, H. W. Kuhn, A. W. Tucker, and Claude E. Shannon. John von Neumann, 1903–1957. In Fleming and Bailyn [FB69], pages 235–268. ISBN 0-674-33411-6, 0-674-33412-4. LCCN E169.1 .F6.

Ulam:1958:JN

[Ula58a] S. Ulam. John von Neumann, 1903–1957. Bulletin of the American Mathematical Society, 64(3 (part 2)):1–49 + 1, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL

http://projecteuclid.org/euclid.bams/1183522369.

Ulam:1958:LJN

[Ula58b] Stanisław M. Ulam. The late John von Neumann on computers and the brain. Scientific American, ??(??):127–??, June 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ulam:1980:NIM

[Ula80] Stanisław Marcin Ulam. Von Neumann: The interaction of mathematics and computing. In Metropolis et al. [MHR80], pages 93–99. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Ulam:1982:JN

[Ula82a] S. Ulam. John von Neumann, 1903–1957. Annals of the History of Computing, 4(2):157–181, 1982. CODEN AHCOE5. ISSN 0164-1239. Reprint.

Ulam:1982:BJN

[Ula82b] Stanisław Ulam. Biographies: John von Neumann, 1903– 1957. Annals of the History of Computing, 4(2):157–181, April/ June 1982. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1982/pdf/a2157.pdf;
 http://www.computer.org/annals/an1982/a2157abs.htm.

Ulam:1986:JN

[URR86a] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. John von Neumann, 1903–1957. In Science, computers, and people: from the tree of mathematics [URR86c], chapter 16, pages 169–214. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:JNC

[URR86b] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. John von Neumann on computers and the brain. In Science, computers, and people: from the tree of mathematics [URR86c], chapter 18, pages 223–230. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:SCP

[URR86c] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Science, computers, and people: from the tree of mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1986. ISBN 0-8176-3276-X. xxii + 264 pp. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:NIM

[URR86d] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Von Neumann: The interaction of mathematics and computing. In Science, computers, and people: from the tree of mathematics [URR86c], chapter 17, pages 215–222. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1947:GHS

[UvN47a] S. Ulam and John von Neumann. On the group of homeomorphisms of the surface of the sphere. Bulletin of the American Mathematical Society, 53(??):506–??, ???? 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1947:CSD

[UvN47b] S. M. Ulam and John von Neumann. On combinations of stochastic and deterministic processes. Bulletin of the American Mathematical Society, 53(11):1120, November 1947. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL http://www.ams.org/journals/bull/1947-53-11/S0002-
9904-1947-08918-7/S0002-9904-1947-08918-7.pdf. The ab-

stract notes “… starting with almost every x_{1} (in the sense of Lebesgue measure) and iterating the function f(x) = 4 · (1 − x) one obtains a sequence of numbers on (0, 1) with a computable algebraic distribution. By playing suitable games with numbers ‘drawn’ in this fashion, one can obtain various other distributions, either given explicitly or satisfying given differential or integral equations.”.

Valente:2008:JNM

[Val08] Giovanni Valente. John von Neumann’s mathematical “Utopia” in quantum theory. Studies in History and Philosophy of Modern Physics, 39(4):860–871, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219808000403.

Vargha:1988:CBP

[Var88] Dénes Vargha, editor. COLING Budapest: proceedings of the 12th International Conference on Computational Linguistics, 22–27 August 1988. John von Neumann Society for Computing Sciences, Budapest, Hungary, 1988. ISBN 963-8431-56-3. LCCN P98 .I55 1988. Two volumes.

Varga:1994:NJH

[Var94] Varga Antal. Neumann János. (Hungarian) [John von Neumann]. Polygon, 4(1):1–18, 1994. CODEN ???? ISSN ????

Veblen:1937:GCD

[Veb37] Oswald Veblen. Geometry of complex domains: a seminar conducted by Professors Oswald Veblen and John von Neumann, 1935– 36. The Institute for Advanced Study, Princeton, NJ, USA, 1937. 231 pp. LCCN QA608 .V4. First term lectures by Professor Veblen, second term lectures by Mr. J. W. Givens, notes by Dr. A. H. Taub and Mr. J. W. Givens.

Veisdal:2019:UGJ

[Vei19] Jørgen Veisdal. The unparalleled genius of John von Neumann. Web site., November 11, 2019.

Venezia:2002:CSI

[Ven02] Antonio Venezia. Un caso storico irrisolto: Le critiche di Popper alla logica quantistica di Birkhoff e von Neumann. (Italian) [a case of unsolved history: The criticism of Popper of the quantum logic of Birkhoff and von Neumann]. Historia Mathematica, 29(??):138– 156, ???? 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Venetianer:2003:NJK

[Ven03] Venetianer Pál. Neumann János és korunk biológiája. (Hungarian) [John von Neumann and biology of our time]. Magyar Tudomány [Hungarian Science], 48(12):1529–1532, 2003. CODEN ???? ISSN ????

Veblen:1955:GCD

[VG55] Oswald Veblen and J. W. Givens. Geometry of complex domains, a seminar conducted by Professors Oswald Veblen and John von Neumann, 1935–36. The Institute for Advanced Study, Princeton, NJ, USA, 1955. iii + 259 pp. Lectures by O. Veblen and J. W. Givens. Notes by A. H. Taub and J. W. Givens. Revised edition.

vanHove:1958:NCQ

[vH58] Léon van Hove. Von Neumann’s contributions to quantum theory. Bulletin of the American Mathematical Society, 64(3 (part 2)):95– 99, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.
bams/1183522374.

vanHeijenoort:1967:FGS

[vH67] Jean van Heijenoort. From Frege to Gödel: a source book in mathematical logic, 1879–1931. Harvard University Press, Cambridge, MA, USA, 1967. ISBN 0-674-32450-1 (hardcover), 0-674-32449-8 (paperback). viii + 664 pp. LCCN QA9 .V35 1967. Reprinted 2002.

vonNeumann:1923:ETZ

[vN23] John von Neumann. Zur Einführung der transfiniten Zahlen. (German) [On the introduction of transinite numbers]. Acta litterarum ac scientiarum Regiae Universitatis Hungaricae Franscisco-Josephinae, Section scientiarum mathematicarum, 1(??):199–208, ???? 1923. Reprinted in [Tau61a, Paper 3]. English translation in [vH67, 346–354].

vonNeumann:1925:ESS

[vN25a] John von Neumann. Egyenletesen sürü számsorozatok. (Hungarian) [Gleichmässig dichte Zahlenfolgen. (German) [Uniformly dense number sequences]]. Mat. Lapok, 32(??):32–40, ???? 1925. ISSN 0025-519X. Reprinted in [Tau61a, Paper 5].

vonNeumann:1925:AM

[vN25b] John von Neumann. Eine Axiomatisierung der Mengenlehre. (German) [An axiomatization of set theory]. Journal für die reine und angewandte Mathematik, 154(??):219–240, ???? 1925. CODEN JR-MAA8. ISSN 0075-4102 (print), 1435-5345 (electronic). Reprinted in [Tau61a, Paper 4].

vonNeumann:1926:XXX

[vN26a] John von Neumann. Az általános halmazelmélet axiomatikus felépítése. Doctoral dissertation, University of Budapest, Budapest, Hungary, 1926.

vonNeumann:1926:PTI

[vN26b] John von Neumann. Zur Prüferschen Theorie der idealen Zahlen. (German) [On the Pruefer theory of ideal numbers]. Acta litterarum ac scientiarum Regiae Universitatis Hungaricae Franscisco-Josephinae, Section scientiarum mathematicarum, 2(??):193–227, ???? 1926. Reprinted in [Tau61a, Paper 6].

vonNeumann:1927:MBQ

[vN27a] John von Neumann. Mathematische Begründung der Quantenmechanik. (German) [The mathematical basis of quantum mechanics]. Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen, 1(??):1–57, ???? 1927. Reprinted in [Tau61a, Paper 9].

vonNeumann:1927:TQG

[vN27b] John von Neumann. Thermodynamik quantenmechanischer Gesamtheiten. (German) [Thermodynamics of quantum-mechan-ical populations]. Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen, 1(??):273–291, ???? 1927. Reprinted in [Tau61a, Paper 11].

vonNeumann:1927:WAQ

[vN27c] John von Neumann. Wahrscheinlichkeitstheoretischer Aufbau der Quantenmechanik. (German) [Probabilistic structure of quantum mechanics]. Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen, 1(??):245–272, ???? 1927. Reprinted in [Tau61a, Paper 10].

vonNeumann:1927:HB

[vN27d] John von Neumann. Zur Hilbertschen Beweistheorie. (German) [On Hilbert proof theory]. Mathematische Zeitschrift, 26(??):1– 46, ???? 1927. CODEN MAZEAX. ISSN 0025-5874, 1432-1823. Reprinted in [Tau61a, Paper 12].

vonNeumann:1927:TDK

[vN27e] John von Neumann. Zur Theorie der Darstellungen kontinuierlicher Gruppen. (German) [On the theory of representations of continuous groups]. Sitzungsberichte der Preussische Akademie, ??(??): 76–90, 1927. Reprinted in [Tau61a, Paper 8].

vonNeumann:1928:AES

[vN28a] John von Neumann. Allgemeine Eigenwerttheorie symmetrischer Funktionaloperatoren. (German) [General Eigenvalue Theory of Symmetric Functional Operators]. Habilitationsschrift [postdoctoral thesis required to quality as a university lecturer], University of Berlin, Berlin, Germany, ???? 1928.

vonNeumann:1928:AMG

[vN28b] John von Neumann. Die Axiomatisierung der Mengenlehre. (German) [The axiomatization of set theory]. Mathematische Zeitschrift, 27(??):669–752, ???? 1928. CODEN MAZEAX. ISSN 0025-5874, 1432-1823. Reprinted in [Tau61a, Paper 16].

vonNeumann:1928:ZIA

[vN28c] John von Neumann. Die Zerlegung eines Intervalles in abzählbar viele kongruente Teilmengen. (German) [Splitting the interval into countable many congruent subsets]. Fundamenta mathematicae, 11(??):230–238, ???? 1928. ISSN 0016-2736 (print), 1730-6329 (electronic). Reprinted in [Tau61a, Paper 13].

vonNeumann:1928:ESF

[vN28d] John von Neumann. Eigenwerteproblem symmetrischer Functionaloperatoren. (German) [Eigenvalue problems for symmetric operators]. Jahresbericht der Deutschen Mathematiker-Vereinigung (Stuttgart), 37(??):11–14, ???? 1928.

vonNeumann:1928:SAU

[vN28e] John von Neumann. Ein System algebraisch unabhängiger Zahlen. (German) [A system of algebraically independent numbers]. Mathematische Annalen, 99(??):134–141, ???? 1928. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61a, Paper 14].

vonNeumann:1928:EBD

[vN28f] John von Neumann. Einige Bemerkungen zur Diracschen Theorie des Drehelektrons. (German) [Some remarks on the Dirac theory of electron spin]. Zeitschrift für Physik, 48(??):868–881, ???? 1928. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 17].

vonNeumann:1928:TJF

[vN28g] John von Neumann. Sur la théorie des jeux. (French) [On the theory of games]. Comptes rendus de l’Académie des sciences, Paris, 186(??):1689–1691, June 18, 1928. URL http://gallica.bnf.fr/

ark:/12148/bpt6k3139f/f1691.image.pagination.

vonNeumann:1928:DDT

[vN28h] John von Neumann. Über die Definition durch transfinite Induktion und verwandte Fragen der allgemeinen Mengenlehre. (German) [On the definition by transfinite induction and related questions of general set theory]. Mathematische Annalen, 99(??):373– 391, ???? 1928. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61a, Paper 15].

vonNeumann:1928:TGG

[vN28i] John von Neumann. Zur Theorie der Gesellschaftsspiele. (German) [On the theory of games of strategy]. Mathematische Annalen, 100(??):295–320, ???? 1928. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). English translation in [Bar59]. Reprinted in [Tau63c, Paper 1].

vonNeumann:1929:AEH

[vN29a] John von Neumann. Allgemeine Eigenwerttheorie Hermitescher Funktionaloperatoren. (German) [General eigenvalue theory of Hermitian functional operators]. Mathematische Annalen, 102(??): 49–131, ???? 1929. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61b, Paper 1].

vonNeumann:1929:BET

[vN29b] John von Neumann. Beweis des Ergodensatzes und des H-Theorems in der neuen Mechanik. (German) [Proof of ergodic theories and H theorems in the new mechanics]. Zeitschrift für Physik, 57(??):30–70, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 25].

vonNeumann:1929:AEG

[vN29c] John von Neumann. Über die analytischen Eigenschaften von Gruppen linearer Transformationen und ihrer Darstellungen. (German) [On the analytic characteristics of groups of linear transformations and their representations]. Mathematische Zeitschrift, 30 (??):3–42, ???? 1929. CODEN MAZEAX. ISSN 0025-5874, 1432-1823. Reprinted in [Tau61a, Paper 22].

vonNeumann:1929:WAM

[vN29d] John von Neumann. Über eine Widerspruchfreiheitsfrage in der axiomatischen Mengenlehre. (German) [On a contradiction of the consistency question in axiomatic set theory]. Journal für die reine und angewandte Mathematik, 160(??):227–241, ???? 1929. CO-DEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic). Reprinted in [Tau61a, Paper 21].

vonNeumann:1929:BMS

[vN29e] John von Neumann. Zum Beweise des Minkowskischen Sätzes über Linearformen. (German) [On the proof of the Minkowski theorem over linear forms]. Mathematische Zeitschrift, 30(??):1–2, ????

1929. CODEN MAZEAX. ISSN 0025-5874, 1432-1823. Reprinted in [Tau61b, Paper 10].

vonNeumann:1929:AFT

[vN29f] John von Neumann. Zur Algebra der Funktionaloperatoren und Theorie der normalen Operatoren. (German) [On the algebra for functional operators and the theory of normal operators]. Mathematische Annalen, 102(??):370–427, ???? 1929. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61b, Paper 2].

vonNeumann:1929:ATM

[vN29g] John von Neumann. Zur allgemeinen Theorie des Maßes. (German) [On the general theory of measure]. Fundamenta mathematicae, 13 (??):73–116, ???? 1929. ISSN 0016-2736 (print), 1730-6329 (electronic). Reprinted in [Tau61a, Paper 26].

vonNeumann:1929:TUM

[vN29h] John von Neumann. Zur Theorie der unbeschränkten Matrizen. (German) [On the theory of unrestricted stencils]. Journal für die reine und angewandte Mathematik, 161(??):208–236, ???? 1929. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic). Reprinted in [Tau61b, Paper 3].

vonNeumann:1929:ZAA

[vN29i] John von Neumann. Zusatz zur Arbeit ,,Zur allgemeinen Theorie des Maßes“. (German) [Addition to the work “On the general theory of measure”]. Fundamenta mathematicae, 13(??):333, ???? 1929. ISSN 0016-2736 (print), 1730-6329 (electronic). Reprinted in [Tau61a, Paper 27].

vonNeumann:1930:HVG

[vN30] John von Neumann. Über einen Hilfssatz der Variationsrechnung. (German) [On a lemma of the calculus of variations]. Abhandlungen Hamburg, 8(??):28–31, ???? 1930. Reprinted in [Tau61b, Paper 4].

vonNeumann:1931:ARS

[vN31a] John von Neumann. Algebraische Repräsentanten der Funktionen ,,bis auf eine Menge vom Maße Null“. (German) [Algebraic representations of functions “Up to a quantity of zero mass”]. Journal für die reine und angewandte Mathematik, 165(??):109–115, ???? 1931. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic). Reprinted in [Tau61b, Paper 6].

vonNeumann:1931:BAH

[vN31b] John von Neumann. Bemerkungen zu den Ausführungen von Herrn St. Lesniewski über meine Arbeit ,,Zur Hilbertschen Beweistheorie“. (German) [Remarks on the comments of Mr. St. Lesniewski on my work “On Hilbert proof theory”]. Fundamenta mathematicae, 17(??):331–334, ???? 1931. ISSN 0016-2736 (print), 1730-6329 (electronic). Reprinted in [Tau61b, Paper 8].

vonNeumann:1931:ESO

[vN31c] John von Neumann. Die Eindeutigkeit der Schrödingerschen Operatoren. (German) [The uniqueness of Schrödinger operators]. Mathematische Annalen, 104(??):570–578, ???? 1931. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic). Reprinted in [Tau61b, Paper 7].

vonNeumann:1931:FGM

[vN31d] John von Neumann. Die formalistische Grundlegung der Mathematik. (German) [The formal basis of mathematics]. Erkennttliss, zugleich Annalen der Philosophie (Leipzig), 2(??):116–121, 1931. Reprinted in [Tau61b, Paper 9].

vonNeumann:1931:FFG

[vN31e] John von Neumann. Über Funktionen von Funktionaloperatoren. (German) [On functions of functional operators]. Annals of Mathematics (2), 32(2):191–226, ???? 1931. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 5].

vonNeumann:1932:ESM

[vN32a] John von Neumann. Einige Sätze über messbare Abbildungen. (German) [Some theorems of measurable diagrams]. Annals of Mathematics (2), 33(3):574–586, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 16].

vonNeumann:1932:MGQ

[vN32b] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical Foundations of Quantum Mechanics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1932. ??? + ???? pp. LCCN ???? Reprinted: Dover Publications (1943); Presses Universitaires de France (1947); Madrid, Institute de Matematicas “Jorge Juan”

(1949). Translated from German by Robert T. Beyer, Princeton University Press (1955).

vonNeumann:1932:PAE

[vN32c] John von Neumann. Physical applications of the ergodic hypothesis. Proceedings of the National Academy of Sciences of the United States of America, 18(??):263–266, ???? 1932. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau61b, Paper 13].

vonNeumann:1932:PQE

[vN32d] John von Neumann. Proof of the quasi-ergodic hypothesis. Proceedings of the National Academy of Sciences of the United States of America, 18(??):70–82, ???? 1932. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau61b, Paper 12].

vonNeumann:1932:AF

[vN32e] John von Neumann. Über adjungierte Funktionaloperatoren. (German) [On adjoint functional operators]. Annals of Mathematics (2), 33(2):294–310, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 11].

vonNeumann:1932:AFG

[vN32f] John von Neumann. Über adjungierte Funktionaloperatoren. (German) [On adjoint functional operators]. Annals of Mathematics, 33(??):294–310, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

vonNeumann:1932:SHM

[vN32g] John von Neumann. Über einen Satz von Herrn M. H. Stone. (German) [On a theorem of Mr. M. H. Stone]. Annals of Mathematics (2), 33(3):567–573, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 15].

vonNeumann:1932:OKM

[vN32h] John von Neumann. Zur Operatorenmethode in der klassischen Mechanik. (German) [On the operator methods of classical mechanics]. Annals of Mathematics (2), 33(3):587–642, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). See additions [vN32i]. Reprinted in [Tau61b, Paper 17].

vonNeumann:1932:ZAG

[vN32i] John von Neumann. Zusätze zur Arbeit ,,Zur Operatorenmethode in der klassischen Mechanik“. (German) [Additions to the work “On operator methods in classical mechanics”]. Annals of Mathematics (2), 33(4):789–791, ???? 1932. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). See [vN32h]. Reprinted in [Tau61b, Paper 18].

vonNeumann:1933:EAP

[vN33] John von Neumann. Die Einführung analytischer Parameter in topologischen Gruppen. (German) [The introduction of analytic parameters in topological groups]. Annals of Mathematics (2), 34 (1):170–190, ???? 1933. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau61b, Paper 19].

vonNeumann:1934:APFa

[vN34a] John von Neumann. Almost periodic functions in a group, I. Bulletin of the American Mathematical Society, 40(??):224–??, ???? 1934. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1934:APFb

[vN34b] John von Neumann. Almost periodic functions in a group. I. Transactions of the American Mathematical Society, 36(3):445–492, ???? 1934. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850

(electronic). URL http://www.ams.org/profession/prizes-
awards/ams-prizes/bocher-prize. Reprinted in [Tau61b, Paper

23]. This two-part paper won the American Mathematical Society Bôcher Memorial Prize in 1938.

vonNeumann:1934:HMT

[vN34c] John von Neumann. Zum Haarschen Mass in topologischen Gruppen. (German) [On the Haar measure in topological groups]. Compositio Mathematica, 1(??):106–114, ???? 1934. CODEN CMP-MAF. ISSN 0010-437X (print), 1570-5846 (electronic). A Russian translation appears in Uspekhi Mat. Nauk, 2(1936):168–176. Reprinted in [Tau61b, Paper 22].

vonNeumann:1934:MFS

[vN35a] John von Neumann. Measure in functional spaces. Unpublished manuscript. Reviewed by I. Halperin. Reprinted in [Tau62, Paper 31, pp. 435–438]., 1934–1935.

vonNeumann:1935:CSI

[vN35b] John von Neumann. Charakterisierung des Spektrums eines Integraloperators. (German) [Characterization of the spectra of integral operators]. Actualités Scientifiques et Industriels. Exposés Math., publiés à la mémoire de J. Herbrand, 229(13):??, 1935. Reprinted in [Tau62, Paper 2].

vonNeumann:1935:CTSa

[vN35c] John von Neumann. On complete topological spaces. Bulletin of the American Mathematical Society, 41(??):35–??, ???? 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1935:CTSb

[vN35d] John von Neumann. On complete topological spaces. Transactions of the American Mathematical Society, 37(1):1–20, ???? 1935. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). Reprinted in [Tau61b, Paper 25].

vonNeumann:1935:NO

[vN35e] John von Neumann. On normal operators. Proceedings of the National Academy of Sciences of the United States of America, 21(??):366–369, ???? 1935. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 3].

vonNeumann:1935:RGD

[vN35f] John von Neumann. On relativistic gas-degeneracy and the collapsed configurations of stars. Unpublished notes. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 15, pp. 173–174]., 1935.

vonNeumann:1935:RRR

[vN35g] John von Neumann. Representations and ray-representations in quantum mechanics. Bulletin of the American Mathematical Society, 41(??):305–??, ???? 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1935:SSE

[vN35h] John von Neumann. Static solutions of Einstein field equations for perfect fluid with T_{ρ}^{ρ} = 0. Unpublished manuscript. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 14, p. 172]., 1935.

vonNeumann:1936:CG

[vN36a] John von Neumann. Continuous geometry. Proceedings of the National Academy of Sciences of the United States of America, 22(??): 92–100, ???? 1936. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 8].

vonNeumann:1936:ECG

[vN36b] John von Neumann. Examples of continuous geometries. Proceedings of the National Academy of Sciences of the United States of America, 22(??):101–108, ???? 1936. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 9].

vonNeumann:1936:CTR

[vN36c] John von Neumann. On a certain topology for rings of operators. Annals of Mathematics (2), 37(1):111–115, ???? 1936. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 1].

vonNeumann:1936:AGQ

[vN36d] John von Neumann. On an algebraic generalization of the quantum mechanical formalism. I. Rec. Math. Moscou, n. Ser., 1(??):415– 482, ???? 1936. Reprinted in [Tau63a, Paper 9].

vonNeumann:1936:RR

[vN36e] John von Neumann. On regular rings. Proceedings of the National Academy of Sciences of the United States of America, 22(??):707– 713, ???? 1936. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 10].

vonNeumann:1936:UIL

[vN36f] John von Neumann. On the uniqueness of invariant Lebesgue measures. Bulletin of the American Mathematical Society, 42(??):343– ??, ???? 1936. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1936:UHM

[vN36g] John von Neumann. The uniqueness of Haar’s measure. Rec. Math. Moscou, n. Ser., 1(??):721–734, ???? 1936. Reprinted in [Tau62, Paper 6].

vonNeumann:1937:ATC

[vN37a] John von Neumann. Algebraic theory of continuous geometries. Proceedings of the National Academy of Sciences of the United States of America, 23(??):16–22, ???? 1937. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 11].

vonNeumann:1937:CGT

[vN37b] John von Neumann. Continuous geometries with a transition probability. Unpublished manuscript. Reviewed by I. Halperin. Reprinted in [Tau62, Paper 16, pp. 191–194] and [vNH81]., 1937.

vonNeumann:1937:CRT

[vN37c] John von Neumann. Continuous rings and their arithmetics. Proceedings of the National Academy of Sciences of the United States of America, 23(??):341–349, ???? 1937. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau62, Paper 12].

vonNeumann:1937:QLS

[vN37d] John von Neumann. Quantum logistics (strict- and probabilitylogics). Unfinished manuscript. Reviewed by I. Halperin. Reprinted in [Tau62, Paper 17, pp. 195–197]., 1937.

vonNeumann:1937:SMI

[vN37e] John von Neumann. Some matrix-inequalities and metrization of matrix-space. Tomskii Univ. Rev., 1(??):286–300, ???? 1937. Reprinted in [Tau62, Paper 20, pp. 205–219].

vonNeumann:1937:OGV

[vN37f] John von Neumann. Über ein ökonomisches Gleichungssystem und eine Verallgemeinerung des Brouwerschen Fixpunktsatzes. (German) [On an economics equation system and a generalization of the Brouwer Fixed-Point Theorem]. Ergebn. eines Math. Kolloqu., 8(??):73–83, ???? 1937. Reprinted in [Tau63c, Paper 3]. English translation in [vN45b].

vonNeumann:1938:IDP

[vN38] John von Neumann. On infinite direct products. Compositio mathematica, Groningen, 6(??):1–77, ???? 1938. URL http://www.

numdam.org/item?id=CM_1939__6__1_0. Reprinted in [Tau63a,

Paper 6].

vonNeumann:1939:RCL

[vN39] John von Neumann. Representation of certain linear groups by unitary operators in Hilbert space. Unpublished manuscript. Reviewed by G. W. Mackey. Reprinted in [Tau62, Paper 32, pp. 439–441]., 1939.

vonNeumann:1940:DSS

[vN40a] John von Neumann. Discussion of De Sitter’s space and of Dirac’s equation in it. Unpublished manuscript. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 18, p. 177]., 1940.

vonNeumann:1940:LAG

[vN40b] John von Neumann. Lattice Abelian groups. Unpublished manuscript. Reviewed by G. Birkhoff. Reprinted in [Tau62, Paper 18, pp. 198–199]., 1940.

vonNeumann:1940:ROI

[vN40c] John von Neumann. On rings of operators. III. Annals of Mathematics, 41(??):94–161, ???? 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 4].

vonNeumann:1941:AOA

[vN41a] John von Neumann. Appendix: Optimum aiming at an imperfectly located target. Report 241, Ballistic Research Laboratory, Aberdeen, MD, USA, July 3, 1941. Reprinted in [Tau62, Paper 37].

vonNeumann:1941:DRM

[vN41b] John von Neumann. Distribution of the ratio of the mean square successive difference to the variance. Annals of mathematical statistics, 12(4):367–395, ???? 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177731677. Reprinted in

[Tau62, Paper 34].

vonNeumann:1941:GLT

[vN41c] John von Neumann. The general and logical theory of automata. In Cerebral Mechanisms in Behavior, pages 1–41. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1941.

vonNeumann:1941:PSS

[vN41d] John von Neumann. The point source solution. U.S. Government Document AM-9, National Defense Research Council, Division B, Washington, DC, USA, June 30, 1941.

vonNeumann:1942:APM

[vN42a] John von Neumann. Approximative properties of matrices of high finite order. Portugaliae Mathematica, 3(??):1–62, ???? 1942. CO-DEN POMAAJ. ISSN 0032-5155. Reprinted in [Tau62, Paper 24].

vonNeumann:1942:FRC

[vN42b] John von Neumann. A further remark concerning the distribution of the ratio of the mean square successive difference to the variance. Annals of mathematical statistics, 13(1):86–88, 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://projecteuclid.org/euclid.aoms/

1177731645. Reprinted in [Tau62, Paper 35].

vonNeumann:1942:TDW

[vN42c] John von Neumann. Theory of detonation waves. U.S. Government Document PB 31090, ????, ????, May 4, 1942. 34 pp. A progress report to the National Defense Research Committee Div. B, OSRD-549, April 1, 1942. Reprinted in [Tau63c, Paper 20].

vonNeumann:1943:ORS

[vN43a] John von Neumann. Oblique reflection of shocks. U.S. Government Document PB 37079, ????, ????, October 12, 1943. Reprinted in [Tau63c, Paper 22].

vonNeumann:1943:SAP

[vN43b] John von Neumann. On some algebraical properties of operator rings. Annals of Mathematics, 44:709–715, 1943. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 8].

vonNeumann:1943:SWB

[vN43c] John von Neumann. [Shock waves in boostered detonations]. The provenance of this undated, and previously unknown, handwritten work is documented in [RKH^{+}21], along with a transcription of the manuscript. There is an accompanying supplement written by Raymond J. Seeger, but his work is not included in the 2021 transcription. The manuscript is believed to be from 1942 or 1943., 1943.

vonNeumann:1943:TSW

[vN43d] John von Neumann. Theory of shock waves. U.S. Government Document PB 32719, ????, ????, January 29, 1943. Progress report to August 31, 1942. Reprinted in [Tau63c, Paper 19].

vonNeumann:1944:IRS

[vN44a] John von Neumann. Introductory remarks (Sec. I), theory of the spinning detonation (Sec. XII), theory of the intermediate product (Sec. XIII). U.S. Government Document AM-570, ????, ????, April 10, 1944. 19 pp. Report of Informal Technical Conference on the Mechanism of Detonation.

vonNeumann:1944:PAN

[vN44b] John von Neumann. Proposal and analysis of a numerical method for the treatment of hydrodynamical shock problems. U.S. Government Document OSRD-3617, ????, ????, March 20, 1944. 31 pp. Reprinted in [Tau63c, Paper 27].

vonNeumann:1944:RMS

[vN44c] John von Neumann. Riemann method: Shock waves and discontinuities (one-dimensional), two-dimensional hydrodynamics. In Lectures in Shock Hydrodynamics and Blast Waves [BFvN^{+}44], page 106. Notes by J. O. Hirschfelder.

vonNeumann:1945:FDR

[vN45a] John von Neumann. First draft of a report on the EDVAC. Technical report, University of Pennsylvania, June 30, 1945. Report prepared for U.S. Army Ordinance Department under Contract W-670-ORD-4926. Reprinted in [Ste81, pp. 177–246], [Ran82, pp. 399–413], [vN93b], and [Lap96].

vonNeumann:1945:MGE

[vN45b] John von Neumann. A model of general economic equilibrium. Review of Economic Studies, 13(1):1–9, ???? 1945. CODEN ???? ISSN 0034-6527 (print), 1467-937X (electronic). URL https://

www.jstor.org/stable/2296111. English translation by G. Mor-

genstern of German original [vN37f].

vonNeumann:1945:RIR

[vN45c] John von Neumann. Refraction, intersection and reflection of shock waves. In Conference on Supersonic Flow and Shock Waves (March 15, Navy Bureau of Ordnance), pages 4–12. ????, ????, July 16, 1945. U.S. Government Document AM-1663. NAVORD Report 203, July 16. Reprinted in [Tau63c, Paper 23].

vonNeumann:1945:UVM

[vN45d] John von Neumann. Use of variational methods in hydrodynamics. Memorandum to O. Veblen. Reprinted in [Tau63c, Paper 26, pp. 357–360]., March 26, 1945.

vonNeumann:1946:SBS

[vN46a] John von Neumann. Statement before the Special Senate Committee on Atomic Energy. Unpublished manuscript, prepared prior to the January 31, 1946, hearing. For minutes of the actual testimony, see Atomic Energy Act of 1946, U.S. Printing Office. Reprinted in [Tau63c, Paper 37, pp. 499–502]., 1946.

vonNeumann:1946:SSS

[vN46b] John von Neumann. Symmetric solutions of some general N person games. Unpublished manuscript. Reviewed by D. B. Gillies. Reprinted in [Tau63c, Paper 10, pp. 98–99]., 1946.

vonNeumann:1947:AEC

[vN47a] John von Neumann. Automatic electronic computers. In ????, editor, American Institute of Electrical Engineers (AIEE) Winter Meeting, January 27–31 1947, page ?? American Institute of Electrical Engineers, ????, 1947. URL https://ethw.org/w/images/

b/b4/AR-641_-_AIEE_Winter_Meeting%2C_Jan_27-31%2C_1947_
-_Program_-_Compact.pdf.

vonNeumann:1947:FRR

[vN47b] John von Neumann. The future role of rapid computing: Meteorology. Aeronautical Engineering Review, 6(??):30–??, ???? 1947. ISSN 0095-9812.

vonNeumann:1947:FMM

[vN47c] John von Neumann. Les fondements mathématiques de la mécanique quantique. (French) [The Mathematical Foundations of Quantum Mechanics]. Felix Alcan, Paris, France, 1947. 336 pp.

vonNeumann:1947:M

[vN47d] John von Neumann. The mathematician. In Heywood [Hey47], pages 180–196. LCCN BF408 .C5. Reprinted in [Tau61a, Paper 1].

vonNeumann:1947:MP

[vN47e] John von Neumann. On a maximization problem. Manuscript, Institute for Advanced Studies, Princeton University, Princeton, NJ 08544, USA, 1947.

vonNeumann:1947:CSD

[vN47f] John von Neumann. On combination of stochastic and deterministic processes. Bulletin of the American Mathematical Society, 53 (??):1120–??, ???? 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1948:FRN

[vN48a] John von Neumann. First report on the numerical calculation of flow problems. Prepared for Standard Oil Development Company, June 22–July 6, 1948. Reprinted in [Tau63b, Paper 19, pp. 664– 712]., 1948.

vonNeumann:1948:NMD

[vN48b] John von Neumann. A numerical method for determination of the value and the best strategies of a zero-sum two-person game with large numbers of strategies. Manuscript/mimeograph. Reviewed by H. W. Kuhn and A. W. Tucker. Reprinted in [Tau63c, Paper 9, pp. 96–97]., 1948.

vonNeumann:1948:TSD

[vN48c] John von Neumann. On the theory of stationary detonation waves. File X122, Ballistic Research Laboratory, Aberdeen Proving Ground, Aberdeen, Maryland, September 20, 1948. 26 pp.

vonNeumann:1948:SRN

[vN48d] John von Neumann. Second report on the numerical calculation of flow problems. Prepared for Standard Oil Development Company, July 25–August 22, 1948. Reprinted in [Tau63b, Paper 20, pp. 713– 750]., 1948.

vonNeumann:1949:BRN

[vN49a] John von Neumann. Book review: Norbert Wiener, Cybernetics or Control and Communication in the Animal and the Machine. Physics Today, 2(5):33–34, May 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v2/i5/p33_s1.

vonNeumann:1949:FMM

[vN49b] John von Neumann. Fundamentos matemáticos de la mecánica cuántica. (Spanish) [Mathematical foundations of quantum mechanics], volume 1 of Monografias de matemátika. Manuales y Textos Universitarios Ciencas. Consejo Superior de Investigaciones Cientificas, Madrid, Spain, 1949. ???? pp. Translated from the 1932 German original by Ramón Ortiz Fornaguera.

vonNeumann:1949:MGQ

[vN49c] John von Neumann. Mathematische Grundlagen der Quantenmechanik. Aus dem Deutschen übersetzt von R. Ortiz. (Spanish) [The Mathematical Foundations of Quantum Mechanics. Translated from the German by R. Ortiz]. Consejo superior de investigaciones cientificas. Monografias de matematica. Instituto de matematicas ‘Jorge Juan’, Madrid, Spain, 1949. vi + 351 pp.

vonNeumann:1949:ROR

[vN49d] John von Neumann. On rings of operators: Reduction theory. Annals of Mathematics (2), 50(2):401–485, April 1949. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau63a, Paper 7].

vonNeumann:1949:RTT

[vN49e] John von Neumann. Recent theories of turbulence. Unpublished report to the Office of Naval Research. Reprinted in [Tau63c, Paper 33, p. 437–472]., 1949.

vonNeumann:1950:FOV

[vN51a] John von Neumann. Functional Operators. Vol. 1: Measures and integrals. Vol 2: The geometry of orthogonal spaces. Number 21, 22 in Annals of Mathematics Studies. Princeton University Press, Princeton, NJ, 1950–1951. 261 (vol. 1), 107 (vol. 2) pp. LCCN QA320 .V64.

vonNeumann:1951:VTU

[vN51b] John von Neumann. 13. Various techniques used in connection with random digits. In Householder et al. [HFG51], pages 36–

38. URL https://dornsifecms.usc.edu/assets/sites/520/
docs/VonNeumann-ams12p36-38.pdf. Summary written by G. E.

Forsythe. Reprinted in [Tau63b, Paper 23, pp. 768–770].

vonNeumann:1951:DCM

[vN51c] John von Neumann. Description of the conformal mapping method for the integration of partial differential equation systems with 1 + 2 independent variables. Unpublished manuscript, December 16, 1950–January 8, 1951. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 34, p. 473–476]., 1951.

vonNeumann:1951:DEU

[vN51d] John von Neumann. Discussion on the existence and uniqueness or multiplicity of solutions of the aerodynamical equations. In Anonymous [Ano51], pages 75–84. LCCN QB790 .S9 1949. Reprinted in [Tau63c, Paper 25].

vonNeumann:1951:SAO

[vN51e] John von Neumann. Eine Spektraltheorie für allgemeine Operatoren eines unitären Raumes. (German) [A spectral theory for general operators in unitary spaces]. Mathematische Nachrichten, 4:258–281, 1951. CODEN MTMNAQ. ISSN 0025-584X. Reprinted in [Tau62, Paper 26].

vonNeumann:1951:FHS

[vN51f] John von Neumann. The future of high-speed computing. In Proceedings of a Computation Seminar [IBM Department of Education, Endicot, NY, from December 5 to 9, 1949], page 13. IBM, New York, NY, USA, 1951. LCCN ???? Digest of an address at the IBM Seminar on Scientific Computation, November, 1949. Reprinted in [Tau63b, Paper 6].

vonNeumann:1951:GLT

[vN51g] John von Neumann. The general and logical theory of automata. In Jeffress [Jef51], pages 1–31. LCCN BF181 .J4; QP351. Reprinted in [Tau63b, Paper 9].

vonNeumann:1951:TKG

[vN51h] John von Neumann. Tribute to Kurt Gödel. New York Times, ??(??):31, March 15, 1951. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

vonNeumann:1952:DRC

[vN52a] John von Neumann. Discussion remark concerning paper of C. S. Smith, “Grain shapes and other metallurgical applications of topology”. In American Society for Metals [Ame52], pages 108– 110. LCCN QD171 .A62. Reprinted in [Tau63c, Paper 41].

vonNeumann:1952:FNC

[vN52b] John von Neumann. Five notes on continuous geometry. Prepared by W. Givens, Univ. of Tennessee., 1952.

vonNeumann:1953:CZS

[vN53a] John von Neumann. A certain zero-sum two-person game equivalent to the optimal assignment problem. Ann. Math. Studies, 28 (??):5–12, ???? 1953. Seminar talk, October 26, 1951, Notes by H. Rogers, Jr. Reprinted in [Tau63c, Paper 5].

vonNeumann:1953:CBN

[vN53b] John von Neumann. Communication on the Borel notes. Econometrica, 21:124–127, 1953. CODEN ECMTA7. ISSN 0012-9682. Reprinted in [Tau63c, Paper 2].

vonNeumann:1953:NPD

[vN53c] John von Neumann. Notes on the photon-disequilibrium-amplification scheme. Reviewed by J. Bardeen. Reprinted in [Tau63b, Paper 12, p. 420]., September 16, 1953.

vonNeumann:1953:TIP

[vN53d] John von Neumann. The Taylor instability problem. Unpublished manuscript (year uncertain). Reviewed by H. H. Goldstine. Reprinted in [Tau63c, Paper 32, pp. 435–436]., 1953.

vonNeumann:1954:EAN

[vN54a] John von Neumann. Entwicklung und Ausnutzung neuerer mathematischer Maschinen. (German) [Development and use of new mathematical machines]. Arbeitsgemeinschaft für Forschung des Landes Nordrhein-Westfalen, 45:??, September 15, 1954. Reprinted in [Tau63b, Paper 8].

vonNeumann:1954:NLC

[vN54b] John von Neumann. Non-linear capacitance or inductance switching, amplifying and memory devices. Basic paper for Patent 2,815,488, filed April 28, 1954, granted December 3, 1957, and assigned to IBM. Reprinted in [Tau63b, Paper 11, pp. 379–419].,

1954. URL http://www.google.com/patents/US2815488.

vonNeumann:1954:NPH

[vN54c] John von Neumann. The NORC and problems in high-speed computing. Speech at first public showing of IBM Naval Ordnance Research Calculator, December 2., 1954. Reprinted in [Tau63b, Paper 7] and [vN81a].

vonNeumann:1954:NMD

[vN54d] John von Neumann. A numerical method to determine optimum strategy. Naval Research Logistics Quarterly, 1(2):109–115, ???? 1954. CODEN NRLQAR. ISSN 0028-1441 (print), 1931-9193 (electronic). Reprinted in [Tau63c, Paper 7].

vonNeumann:1954:RMS

[vN54e] John von Neumann. The role of mathematics in the sciences and in society. Address at 4th Conference of Association of Princeton Graduate Alumni, June 1954., 1954. Reprinted in [Tau63c, Paper 35].

vonNeumann:1955:CWS

[vN55a] John von Neumann. Can we survive technology? Fortune, 91(6): 106–108, 151–152, June 1955. CODEN FORTAP. ISSN 0015-8259. Reprinted in [Tau63c, Paper 38].

vonNeumann:1955:DAW

[vN55b] John von Neumann. Defense in atomic war. Journal of the American Ordnance Association (Washington, DC), ??(??):21–33, May/ June 1955. ISSN 0030-4557. Paper delivered at a symposium in honor of Dr. R. H. Kent, December 7, 1955. Reprinted in [Tau63c, Paper 40, pp. 523–525].

vonNeumann:1955:IAE

[vN55c] John von Neumann. Impact of atomic energy on the physical and chemical sciences. Technology Review (M.I.T.), ??(??):15–17, November 1955. CODEN TEREAU. ISSN 0040-1692. Speech at MIT Alumni Day Symposium, June 13, 1955. Reprinted in [Tau63c, Paper 39].

vonNeumann:1955:MFQ

[vN55d] John von Neumann. Mathematical Foundations of Quantum Mechanics. Princeton landmarks in mathematics and physics series. Princeton University Press, Princeton, NJ, 1955. ISBN 0-691-08003-8; 0-691-02893-1. xii + 445 pp. LCCN QC174.3 .V613.

Translated from the German edition by Robert T. Beyer. Original first edition published in German in 1932.

vonNeumann:1955:MPS

[vN55e] John von Neumann. Method in the physical sciences. In Leary [Lea55], pages 157–164. Reprinted in [Tau63c, Paper 36].

vonNeumann:1956:IRD

[vN56a] John von Neumann. The impact of recent developments in science on the economy and on economics. Looking Ahead, 4(11):??, ???? 1956. Partial text of a talk at the National Planning Association, Washington, DC, December 12, 1955. Reprinted in [Tau63c, Paper 11].

vonNeumann:1956:PLS

[vN56b] John von Neumann. Probabilistic logics and the synthesis of reliable organisms from unreliable components. In Automata Studies [SM56], pages 43–98. CODEN ANMAAH. ISSN 0066-2313. Lecture notes taken by R. S. Pierce and revised by the author. Fourth Printing 1965. Reprinted in [Tau63b, Paper 10].

vonNeumann:1957:PSS

[vN57a] John von Neumann. The point source solution. In Blast Wave [BFH^{+}47], pages 27–55. LCCN ???? URL https://fas.org/sgp/

othergov/doe/lanl/docs1/00320773.pdf. Reprinted in [Tau63c,

Paper 21].

vonNeumann:1957:TBC

[vN57b] John von Neumann. Testimony before the Congressional Committee on Atomic Energy, January 31, 1946. Bulletin of the Atomic Scientists, 13(4):151, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonNeumann:1958:CB

[vN58a] John von Neumann. The Computer and the Brain. Yale University Press, New Haven, CT, USA, 1958. ISBN 0-300-00793-0. xiv + 82 pp. See second edition [vNCC00].

vonNeumann:1958:NIC

[vN58b] John von Neumann. The non-isomorphism of certain continuous rings. Annals of Mathematics, 67:485–496, 1958. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Introduction by I. Kaplansky. Reprinted in [Tau62, Paper 14].

vonNeumann:1959:NLC

[vN59a] John von Neumann. Non-linear capacitance or inductance switching, amplifying and memory organs. Canadian Patent 579033.,

July 7, 1959. URL http://brevets-patents.ic.gc.ca/opic-
cipo/cpd/eng/patent/579033/summary.html.

vonNeumann:1959:SNL

[vN59b] John von Neumann. Solution of a non-linear differential equation. Bull. Res. Council Israel Sect. F, 8F:133, 1959.

vonNeumann:1960:CG

[vN60a] John von Neumann. Continuous geometry. Princeton University Press, Princeton, NJ, 1960. xi + 299 pp. Introduction by I. Halperin.

vonNeumann:1960:RGG

[vN60b] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the Brain]. R. Oldenbourg Verlag, München, Germany, 1960. 80 pp. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN65, vN70, vN80, vN86, vN91b].

vonNeumann:1961:CFT

[vN61a] John von Neumann. Characterization of factors of type III. Unpublished draft manuscript. Reviewed by I. Kaplansky. Reprinted in [Tau63a, Paper 10, pp. 562–563]., 1961.

vonNeumann:1961:CC

[vN61b] John von Neumann. Comparison of cells. Unpublished manuscript. Reviewed by G. Hunt. Reprinted in [Tau61b, Paper 27, pp. 558ff]., 1961.

vonNeumann:1962:IS

[vN62] John von Neumann. Independence of F_{∞} from the sequence ν. Unpublished manuscript. Reviewed by I. Halperin. Reprinted in [Tau62, Paper 15, pp. 189–190]., 1962.

vonNeumann:1963:DMP

[vN63a] John von Neumann. Discussion of a maximum problem. In Taub [Tau63b], pages 89–95. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

vonNeumann:1963:GLT

[vN63b] John von Neumann. The general and logical theory of automata. In Taub [Tau63b], pages 288–328. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

vonNeumann:1963:PLS

[vN63c] John von Neumann. Probabilistic logics and synthesis of reliable organisms from unreliable components. In Taub [Tau63b], pages ix + 784. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

vonNeumann:1963:VTU

[vN63d] John von Neumann. Various techniques for use in connection with random digits. In Taub [Tau63b], pages 768–770. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

vonNeumann:1964:MGQ

[vN64] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (Russian) [Mathematical Foundations of Quantum Mechanics]. Verlag ‘Nauka’, Moscow, Russia, 1964. 367 pp. Übersetzung aus dem Deutschen von M. K. Pllivanov und B. M. Stepanov. Unter Redaktion von N. N. Bogoljubov.

vonNeumann:1965:RGG

[vN65] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the Brain]. R. Oldenbourg Verlag, München, Germany, second edition, 1965. 80 pp. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN60b, vN70, vN80, vN86, vN91b].

vonNeumann:1966:SSG

[vN66] John von Neumann. A system of 29 states with a general transition rule. In von Neumann and Burks [vNB66], pages 305–317. LCCN QA267 .V55. Von Neumann’s work on self-reproducing automata, completed and edited after his death by Arthur Burks. Also includes transcripts of von Neumann’s 1949 University of Illinois lectures on the “Theory and Organization of Complicated Automata”. See [Bur70].

vonNeumann:1967:MGQ

[vN67] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical Foundations of Quantum Mechanics]. Princeton landmarks in mathematics and physics series. Princeton University Press, Princeton, NJ, 1967. ISBN 0-691-08003-8; 0-691-02893-1. xii + 445 pp. LCCN ???? Second printing. Original first edition published in German in 1932.

vonNeumann:1968:MGQ

[vN68] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical Foundations of Quantum Mechanics], volume 38 of Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. ISBN 3-540-04133-8; 3-540-59207-5. 262 pp. LCCN QC174.12V65 1968. Original first edition published in German in 1932.

vonNeumann:1970:RGG

[vN70] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the brain]. Sciencia Nova and R. Oldenbourg Verlag, München, Germany and Wien, Austria, third edition, 1970. 80 pp. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN60b, vN65, vN80, vN86, vN91b].

vonNeumann:1971:MGQ

[vN71a] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical Foundations of Quantum Mechanics]. Princeton landmarks in mathematics and physics series. Princeton University Press, Princeton, NJ, 1971. ISBN 0-691-08003-8; 0-691-02893-1. xii + 445 pp. LCCN ???? Sixth printing. Original first edition published in German in 1932.

vonNeumann:1971:TSA

[vN71b] John von Neumann. Teoriya samovosproizvodyashchihsya avtomatov. (Russian) [Theory of self-reproducing automata]. Verlag “Mir”, Moscow, Russia, 1971. 382 pp. Edited and completed by Arthur W. Burks. Translated from the English by V. L. Stefanyuka. Edited by V. I. Varshavskij.

vonNeumann:1974:DMP

[vN74] John von Neumann. Discussion of a maximum problem. Unpublished typescript. Reprinted in [Tau63c, Paper 8, pp. 89–95]., November 15–16, 1974.

vonNeumann:1980:RGG

[vN80] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the brain]. Scientia nova. Sciencia Nova and R. Oldenbourg Verlag, München, Germany and Wien, Austria, fourth edition, 1980. ISBN 3-486-45224-X. 80 pp. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN60b, vN65, vN70, vN86, vN91b].

vonNeumann:1981:NPH

[vN81a] John von Neumann. The NORC and problems in high-speed computing. Annals of the History of Computing, 3(3):274– 279, July/September 1981. CODEN AHCOE5. ISSN 0164-

1239.  URL http://dlib.computer.org/an/books/an1981/
pdf/a3274.pdf;   http://www.computer.org/annals/an1981/

a3274abs.htm. Foreword by Herman H. Goldstine.

vonNeumann:1981:PLS

[vN81b] John von Neumann. The principles of large-scale computing machines. Annals of the History of Computing, 3(3): 263–273, July/September 1981. CODEN AHCOE5. ISSN

0164-1239. URL http://dlib.computer.org/an/books/an1981/
pdf/a3263.pdf;   http://www.computer.org/annals/an1981/

a3263abs.htm. Foreword by Nancy Stern.

vonNeumann:1982:FDR

[vN82] John von Neumann. First draft of a report on the EDVAC (1945). In Randell [Ran82], pages 383–392. ISBN 0-387-11319-3, 3-540-11319-3. LCCN TK7885.A5 O741 1982.

vonNeumann:1986:RGG

[vN86] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the brain]. Scientia nova. Sciencia Nova and R. Oldenbourg Verlag, München, Germany and Wien, Austria, fifth edition, 1986. ISBN 3-486-45225-8. 77 pp. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN60b, vN65, vN70, vN80, vN91b].

vonNeumann:1987:ITP

[vN87] John von Neumann. Izbrannye trudy po funktsional’nomu analizu. (Russian) [Selected works on functional analysis]. Klassiki Nauki. Izdatel’stvo “Nauka”, Moscow, Russia, 1987. 376 (vol. 1), 372 (vol.

2) pp. Two volumes.

vonNeumann:1988:PLS

[vN88] John von Neumann. The principles of large-scale computing machines. Annals of the History of Computing, 10(4):243–256, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1988/pdf/a4243.pdf;
 http://www.computer.org/annals/an1988/a4243abs.htm. In-

troduction by Michael R. Williams and foreword by Nancy Stern.

vonNeumann:1991:CB

[vN91a] John von Neumann. Computer and the brain. In Ferris and Fadiman [FF91], pages 478–491. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

vonNeumann:1991:RGG

[vN91b] John von Neumann. Die Rechenmaschine und das Gehirn. (German) [The Computer and the brain]. Scientia nova. Sciencia Nova and R. Oldenbourg Verlag, München, Germany and Wien, Austria, sixth edition, 1991. ISBN 3-486-45226-6. 77 pp. 19.80 EUR. Translated from the American original by Charlotte and Heinz Gumin. See other editions [vN60b, vN65, vN70, vN80, vN86].

vonNeumann:1991:FMM

[vN91c] John von Neumann. Fundamentos matemáticos de la mecánica cuántica. (Spanish) [Mathematical foundations of quantum mechanics], volume 9 of Textos universitarios CSIC. Manuales y Textos Universitarios Ciencas. Consejo Superior de Investigaciones Cientificas, Madrid, Spain, second edition, 1991. ISBN ???? lxii + 351 pp. LCCN ???? Translated from the 1932 German original by Ramón Ortiz Fornaguera.

vonNeumann:1993:CAB

[vN93a] J. von Neumann. Classic article. 2. the brain. M.D. Computing: Computers in Medical Practice, 10(3):173–183, ???? 1993. CODEN MDCOE7. ISSN 0724-6811. Reprinted from the Computer and the Brain, 1958 [vN58a].

vonNeumann:1993:FDR

[vN93b] John von Neumann. First draft of a report on the EDVAC. IEEE Annals of the History of Computing, 15(4):28–75, October/ December 1993. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-

1547 (electronic). URL http://dlib.computer.org/an/books/
an1993/pdf/a4027.pdf;   http://www.computer.org/annals/

an1993/a4027abs.htm. Edited and corrected by Michael D. Godfrey.

vonNeumann:1996:MGQa

[vN96a] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical foundations of quantum mechanics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 3-540-59207-5. x + 262 pp. LCCN QC174.3 .V6 1996.

vonNeumann:1996:MGQb

[vN96b] John von Neumann. Mathematische Grundlagen der Quantenmechanik. (German) [Mathematical Foundations of Quantum Mechanics]. Princeton landmarks in mathematics and physics series. Princeton University Press, Princeton, NJ, 1996. ISBN 0-691-08003-8; 0-691-02893-1. xii + 445 pp. LCCN QC174.3 .V613. Twelfth printing. Original first edition published in German in 1932.

vonNeumann:1996:PJN

[vN96c] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974– 1975. Gift, Nicholas A. Vonneuman, 1993.

vonNeumann:1996:TGL

[vN96d] John von Neumann. Théorie générale et logique des automates. (French) [The General and Logical Theory of Automata]. Éditions Champ Vallon, F 01420 Seyssel, France, 1996. ISBN 2-87673-232-7. 104 pp. LCCN TJ211 .V66 1996. URL http://www.champ-

vallon.com/Pagescollections/Seyssel.html. Translated from

the English edition by Jean-Paul Auffrand. Preceded by La pensée et les machines: le mécanisme algorithmique de John Von Neumann, by Gerard Chazal.

vonNeumann:1998:CGF

[vN98] John von Neumann. Continuous geometry. Princeton Landmarks in Mathematics. Princeton University Press, Princeton, NJ, 1998. ISBN 0-691-05893-8. xi + 299 pp. LCCN QA611 .V6 1998. US$24.95, UK£18.95. With a foreword by Israel Halperin. New edition.

vonNeumann:1999:IM

[vN99] John von Neumann. Invariant measures. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-0912-1. xiv + 134 pp. LCCN QA325 .V66 1999. US$39.00.

vonNeumann:19xx:PSM

[vNxxa] John von Neumann. The point-source model. Unpublished manuscript. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 16, p. 175]., 19xx.

vonNeumann:19xx:PSS

[vNxxb] John von Neumann. The point source solution, assuming a degeneracy of the semi-relativistic type, p = Kρ^{4/3}, over the entire star. Unpublished manuscript. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 17, p. 176]., 19xx.

vonNeumann:2003:LNJ

[vN03] J. von Neumann. Letter from von Neumann to Jack Northrop. IEEE Annals of the History of Computing, 25(1):15, ???? 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). Reprinted from Annals of the History of Computing, vol 9, pg 364– 365, 1988.

vonNeumann:2005:JNS

[vN05] John von Neumann. John von Neumann: selected letters, volume 27 of History of Mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1. xxvi + 301 pp. With a foreword by P. Lax and an introduction by Marina von Neumann Whitman, Edited and with a preface and introductory comments by Miklós Rédei.

vonNeumann:2010:DEU

[vN10] J. von Neumann. Discussion on the existence and uniqueness or multiplicity of solutions of the aerodynamical equations. Bulletin of the American Mathematical Society, 47(1):145–154, ???? 2010. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic). Reprint of [vN51d].

vonNeumann:2012:CBC

[vN12] John von Neumann. The Computer and the Brain. Yale University Press, New Haven, CT, USA, third edition, 2012. ISBN 0-300-18111-6 (paperback); 0-300-18808-0 (e-book). li + 83 pp. LCCN QA76 .V65 2012.

vonNeumann:2013:GVT

[vN13] Klara von Neumann. Grasshopper in the very tall grass. Unpublished and undated memoir in papers of Marina von Neumann Whitman, 1946–2013, Schlesinger Library, Radcliffe Institute for Advanced Study, Harvard University, Cambridge, MA, USA. Transcription by Marina von Neumann Whitman [HPR14a, ref. 31, page 60]., 2013.

vonNeumann:2018:MFQ

[vN18] John von Neumann. Mathematical Foundations of Quantum Mechanics. Princeton University Press, Princeton, NJ, 2018. ISBN 0-691-17856-9 (hardcover), 0-691-17857-7 (paperback), 1-4008-8992-8 (e-book). xviii + 304 pp. LCCN QC174.3 .V613 2018.

URL https://www.jstor.org/stable/j.ctt1wq8zhp. Trans-

lated from the German by Robert T. Beyer. Edited by Nicholas A. Wheeler.

vonNeumann:2021:FDR

[vN21] John von Neumann. First draft of a report on the EDVAC (1945). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 10, pages 89–106. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

vonNeumann:1966:TSR

[vNB66] John von Neumann and Arthur W. (Arthur Walter) Burks, editors. Theory of Self-Reproducing Automata. University of Illinois Press, Urbana, IL, USA, 1966. xix + 388 pp. LCCN QA267 .V55. Von Neumann’s work on self-reproducing automata, completed and edited after his death by Arthur Burks. Also includes transcripts of von Neumann’s 1949 University of Illinois lectures on the “Theory and Organization of Complicated Automata”. See [Bur70].

vonNeumann:1995:NC

[vNBV95] John von Neumann, F. Bródy, and Tibor Vámos, editors. The Neumann compendium, volume 1 of World Scientific series in 20th century mathematics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2201-7, 981-283-108-8 (e-book). lix + 699 pp. LCCN QA300.5

.V66 1995eb. URL http://lib.myilibrary.com?id=494829;
http://search.ebscohost.com/login.aspx?direct=true\%26scope=
site\%26db=nlebk\%26db=nlabk\%26AN=564432;   http://site.
ebrary.com/id/10700739;   http://www.worldscientific.com/
doi/pdf/10.1142/9789812831088_bmatter.

vonNeumann:2000:CB

[vNCC00] John von Neumann, Paul M. Churchland, and Patricia Smith Churchland. The Computer and the Brain. Mrs. Hepsa Ely Silliman memorial lectures; Yale Nota bene. Yale University Press, New Haven, CT, USA, second edition, 2000. ISBN 0-300-08473-0 (paperback). xxviii + 82 pp. LCCN QA76 .V6 2000. See first edition [vN58a].

vonNeumann:1946:OPD

[vNF46] John von Neumann and Klaus Fuchs. One proposed design for the ‘Super’. US Patent application, May 1946. This application is reported in [Ber10, page 43], but is not found in a Google patent search, or in the US Patent & Trademark Office database. Bernstein reports that the copy he has is “essentially totally redacted”. A Wikipedia article on Fuchs says that the patent is still classified.

vonNeumann:1946:PLS

[vNG46] John von Neumann and Herman H. Goldstine. On the principles of large scale computing machines. Unpublished lecture manuscript, prior to May 15, 1946. Reprinted in [Tau63b, Paper 1] and [vNBV95, AB87, GIS90]., 1946.

vonNeumann:1947:NIM

[vNG47] John von Neumann and Herman H. Goldstine. Numerical inverting of matrices of high order. Bulletin of the American Mathematical Society, 53(11):1021–1099, November 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://

projecteuclid.org/euclid.bams/1183511222. See [GvN51] for

Part II. Reprinted in [Tau61a, vol. 5, pp. 479–557].

vonNeumann:1948:PCPa

[vNG48a] John von Neumann and Herman H. Goldstine. Planning and coding of problems for an electronic computing instrument. Part II, vol. II. Technical report, ????, ????, 1948. 68 (or 69??) pp. Report prepared for U.S. Army Ordnance Department under contract W-36-034-ORD-7481. Reprinted in [Tau63b, Paper 4].

vonNeumann:1948:PCPb

[vNG48b] John von Neumann and Herman H. Goldstine. Planning and coding of problems for an electronic computing instrument. Part II, vol. III. Technical report, ????, ????, 1948. Report prepared for U.S. Army Ordnance Department under contract W-36-034-ORD-7481. Reprinted in [Tau63b, Paper 5].

vonNeumann:1953:NSC

[vNG53] John von Neumann and H. H. Goldstine. A numerical study of a conjecture of Kummer. Mathematical Tables and Other Aids to Computation, 7:133–134, 1953. CODEN MTTCAS. ISSN 0891-6837. Reprinted in [Tau63b, Paper 24].

vonNeumann:1937:TPM

[vNH37] John von Neumann and I. Halperin. On the transitivity of perspective mappings in complemented modular lattices. Bulletin of the American Mathematical Society, 43(??):37–??, ???? 1937. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1940:TPM

[vNH40] John von Neumann and I. Halperin. On the transitivity of perspective mappings. Annals of Mathematics, 41(??):87–93, ???? 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 13].

vonNeumann:1941:OMC

[vNH41] John von Neumann and P. R. Halmos. Operator methods in classical mechanics, II. Bulletin of the American Mathematical Society, 47(??):696, ???? 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1981:CGT

[vNH81] John von Neumann and Israel Halperin. Continuous geometries with a transition probability, volume 252 of Memoirs of the American Mathematical Society. American Mathematical Society, Providence, RI, USA, 1981. ISBN 0-8218-2252-7. ISSN 0065-9266 (print), 1947-6221 (electronic). vii + 210 pp. LCCN QA3 .A57 no. 252 QA171.5. Reproduces a manuscript (third draft) written by John von Neumann in 1937 (the original hand-written manuscript is kept in the J. von Neumann file in the Library of Congress). Volume 34.

vonNeumann:1940:EPE

[vNK40] John von Neumann and R. H. Kent. The estimation of the probable error from successive differences. Report 175, Ballistic Research Laboratory, Aberdeen, Maryland, USA, February 14, 1940. 19 pp.

vonNeumann:1941:MSS

[vNKBH41] John von Neumann, R. H. Kent, H. R. Bellinson, and B. I. Hart. The mean square successive difference. Annals of mathematical statistics, 12(2):153–162, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177731746. Reprinted in

[Tau62, Paper 33].

vonNeumann:1936:RO

[vNM36] John von Neumann and F. J. Murray. On rings of operators. Bulletin of the American Mathematical Society, 42(??):808–??, ???? 1936. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1944:TGE

[vNM44] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. Princeton University Press, Princeton, NJ, 1944. xviii + 625 pp. See also second edition [vNM47], third edition [vNM53a, vNM74, vNM80], and 60th anniversary edition [vNM04].

vonNeumann:1947:TGE

[vNM47] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. Princeton University Press, Princeton, NJ, second edition, 1947. LCCN QA269 .V65 1947. See also first edition [vNM44], third edition [vNM53a, vNM74, vNM80], and 60th anniversary edition [vNM04].

vonNeumann:1953:TGEa

[vNM53a] John von Neumann and Oskar Morgenstern. Theory of games and economic behavior. Princeton University Press, Princeton, NJ, third edition, 1953. xx + 644 pp. See also first edition [vNM44], second edition [vNM47], third edition reprints [vNM74, vNM80], and 60th anniversary edition [vNM04].

vonNeumann:1953:TGEb

[vNM53b] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, third edition, 1953. ISBN 0-471-91185-2. viii + 641 pp. LCCN QA269 .V89t 1953b.

vonNeumann:1961:SWV

[vNM61] John von Neumann and Oskar Morgenstern. Spieltheorie und wirtschaftliches Verhalten. (German) [Theory of Games and Economic Behavior]. Physika-Verlag, Würzburg, Germany, 1961. xxiv + 668 pp. See also second edition [vNM67a].

vonNeumann:1964:TGE

[vNM64] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, science edition, 1964. LCCN QA269 .V65 1953a; QA269 .V65 1964.

vonNeumann:1967:SWV

[vNM67a] John von Neumann and Oskar Morgenstern. Spieltheorie und wirtschaftliches Verhalten. (German) [Theory of Games and Economic Behavior]. Physica Verlag, Würzburg, Germany, second edition, 1967. xxiv + 668 pp. LCCN ????

vonNeumann:1967:TGE

[vNM67b] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, third edition, 1967. viii + 641 pp. LCCN QA269 .V65 1953a; QA269 .V89t 1964.

vonNeumann:1970:TGE

[vNM70] John von Neumann and Oskar Morgenstern. Theory of games and economic behavior. Verlag ‘Nauka’: Hauptredaktion für physikalisch-mathematische Literatur, Moscow, Russia, 1970.

vonNeumann:1973:SWV

[vNM73] John von Neumann and Oskar Morgenstern. Spieltheorie und wirtschaftliches Verhalten. (German) [Theory of Games and Economic Behavior]. Physica Verlag, Würzburg, Germany, third edition, 1973. ISBN 3-7908-0134-8. xxiv + 668 pp. LCCN ????

vonNeumann:1974:TGE

[vNM74] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. Princeton University Press, Princeton, NJ, third edition, 1974. ISBN ???? viii + 641 pp. See also first edition [vNM44], second edition [vNM47], third edition reprints [vNM53b, vNM80], and 60th anniversary edition [vNM04].

vonNeumann:1980:TGE

[vNM80] John von Neumann and Oskar Morgenstern. Theory of games and economic behavior. Princeton University Press, Princeton, NJ, third edition, 1980. ISBN 0-691-00362-9. xx + 641 pp. LCCN QA269 .V65 1980. See also first edition [vNM44], second edition [vNM47], third edition reprints [vNM53b, vNM74], and 60th anniversary edition [vNM04].

vonNeumann:2004:TGE

[vNM04] John von Neumann and Oskar Morgenstern. Theory of Games and Economic Behavior. Princeton University Press, Princeton, NJ, 60th anniversary edition, 2004. ISBN 0-691-11993-7. xxxii + 739 pp. LCCN QA269 .V65 2004. With an introduction by Harold W. Kuhn and an afterword by Ariel Rubinstein. See also first edition [vNM44], second edition [vNM47], and third edition reprints [vNM53b, vNM74, vNM80].

vonNeumann:2007:TGE

[vNM07] John von Neumann and Oskar Morgenstern. Theory of games and economic behavior. Princeton classic editions. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-13061-2 (paperback). xxxii + 739 pp. LCCN QA269 .V65 2007. 60th anniversary edition, with an introduction by Harold W. Kuhn, and an afterword by Ariel Rubinstein.

vonNeumann:1947:NSP

[vNR47a] John von Neumann and K. D. Richtmyer. On the numerical solution of partial differential equations of parabolic type. U.S. Government Document LA-657, ????, ????, December 25, 1947. 17 pp. Reprinted in [Tau63a, Paper 18].

vonNeumann:1947:SMN

[vNR47b] John von Neumann and K. D. Richtmyer. Statistical methods in neutron diffusion. U.S. Government Document LA-551, ????, ????, April 9, 1947. 22 pp. Reprinted in [Tau63b, Paper 21].

vonNeumann:1950:MNC

[vNR50] John von Neumann and R. D. Richtmyer. A method for the numerical calculation of hydrodynamic shocks. Journal of Applied Physics, 21:232–237, 1950. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Reprinted in [Tau63c, Paper 28].

vonNeumann:1957:MEH

[vNR57] John von Neumann and F. Reines. The Mach effect and the height of burst. In Blast Wave [BFH^{+}47], pages X–11–X–84. LCCN

???? URL https://fas.org/sgp/othergov/doe/lanl/docs1/

00320773.pdf. Reprinted in [Tau63c, Paper 24].

vonNeumann:1935:DRE

[vNS35] John von Neumann and M. H. Stone. The determination of representative elements in the residual classes of a Boolean algebra. Fundamenta mathematicae, 25(??):353–378, ???? 1935. ISSN 0016-2736 (print), 1730-6329 (electronic). Reprinted in [Tau62, Paper 5].

vonNeumann:1939:FIM

[vNS39] John von Neumann and I. J. Schoenberg. Fourier integrals and metric geometry. Bulletin of the American Mathematical Society, 45(??):79–??, ???? 1939. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1941:FIM

[vNS41] John von Neumann and I. J. Schoenberg. Fourier integrals and metric geometry. Transactions of the American Mathematical Society, 50(??):226–251, ???? 1941. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). Reprinted in [Tau62, Paper 22].

vonNeumann:1943:ORC

[vNS43] John von Neumann and R. J. Seeger. On oblique reflection and collision of shock waves. U.S. Government Document PB 31918, ????, ????, September 20, 1943. 3 pp.

vonNeumann:1945:FWT

[vNT45] John von Neumann and A. H. Taub. Flying wind tunnel experiments. U.S. Government Document PB 33263, ????, ????, November 5, 1945. 16 pp.

vonNeumann:1955:CFE

[vNT55] John von Neumann and Bryant Tuckerman. Continued fraction expansion of 2^{1/3}. Mathematical Tables and Other Aids to Computation, 9(49):23–25, January 1955. CODEN MTTCAS. ISSN 0891-6837. Reprinted in [Tau63b, Paper 25].

vonNeumann:1945:RET

[vNU45] John von Neumann and S. M. Ulam. Random ergodic theorems. Bulletin of the American Mathematical Society, 51(??):660–??, ???? 1945. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1928:EEEa

[vNW28a] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. 1. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin. 1.]. Zeitschrift für Physik, 47(??):203–220, ???? 1928. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 18].

vonNeumann:1928:EEEb

[vNW28b] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. 2. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin. 2.]. Zeitschrift für Physik, 49(??):73–94, ???? 1928. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 19].

vonNeumann:1928:EEEc

[vNW28c] John von Neumann and Eugene Wigner. Zur Erklärung einiger Eigenschaften der Spektren aus der Quantenmechanik des Drehelektrons. 3. (German) [On the explanation of some of the eigenvalue spectra from the quantum mechanics of electron spin. 3]. Zeitschrift für Physik, 51(??):845–888, ???? 1928. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 20].

vonNeumann:1929:VEA

[vNW29a] John von Neumann and Eugene Wigner. Über das Verhalten von Eigenwerten bei adiabatischen Prozessen. (German) [On the behavior of the eigenvalues of adiabatic processes]. Physikalische Zeitschrift (Leipzig), 30(??):467–470, ???? 1929. Reprinted in [Tau61a, Paper 24].

vonNeumann:1929:MDE

[vNW29b] John von Neumann and Eugene Wigner. Über merkwürdige diskrete Eigenwerte. (German) [On unusual discrete eigenvalues]. Physikalische Zeitschrift, 30(??):465–467, ???? 1929. CODEN PHZTAO. ISSN 0369-982X. Reprinted in [Tau61a, Paper 23].

vonNeumann:1940:MAP

[vNW40] John von Neumann and Eugene P. Wigner. Minimally almost periodic groups. Annals of Mathematics (2), 41(??):746–750, ???? 1940. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 21].

Whitman:1990:JNP

[vNW90] Marina v. N. Whitman. John von Neumann: a personal view. In The legacy of John von Neumann (Hempstead, NY, 1988), volume 50 of Proc. Sympos. Pure Math., pages 1–4. American Mathematical Society, Providence, RI, USA, 1990.

Whitman:2012:MDM

[vNW12] Marina von Neumann Whitman. The Martian’s daughter: a memoir. University of Michigan Press, Ann Arbor, MI, USA, 2012. ISBN 0-472-11842-0 (hardcover), 0-472-02855-3 (e-book), 1-283-60181-8. xiv + 308 pp. LCCN HB119.W4 W44 2012.

Vollmar:2006:JNS

[Vol06] Roland Vollmar. John von Neumann and self-reproducing cellular automata. J. Cell. Autom., 1(4):353–376, 2006. ISSN 1557-5969.

Vonneuman:1987:JNS

[Von87] Nicholas A. Vonneuman. John von Neumann: as seen by his brother. Nicholas A. Vonneuman, Meadowbrook, PA, USA (P.O. Box 3097, Meadowbrook 19046), revised edition, 1987. ISBN 0-9619681-0-9. iv + 71 pp. LCCN QA29.V66 V664 1987.

Vonneuman:1988:BJN

[Von88] Nicholas A. Vonneuman. Biography of John von Neumann. Historia Mathematica, 15(3):275, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900687.

Vonneuman:1989:JNF

[Von89] Nicholas Vonneuman. John von Neumann: Formative years. Annals of the History of Computing, 11(3):171–175, July/ September 1989. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1989/pdf/a3171.pdf;
 http://www.computer.org/annals/an1989/a3171abs.htm.

Vonneuman:1990:PLJ

[Von90a] Nicholas Vonneuman. The philosophical legacy of John von Neumann, in the light of its inception and evolution in his formative years. In The legacy of John von Neumann (Hempstead, NY, 1988), volume 50 of Proc. Sympos. Pure Math., pages 19–24. American Mathematical Society, Providence, RI, USA, 1990.

Vonneumann:1990:NJA

[Von90b] Vonneumann Nicholas. Neumann János általános emberi vonásai. (Hungarian) [John von Neumann’s universal human traits]. Fizikai Szemle (Budapest), 40(1):13–17, 1990. CODEN FISZA6. ISSN ????

Vonneuman:1991:JNS

[Von91] Nicholas A. Vonneuman. John von Neumann as seen by his brother. N. A. Vonneuman, Meadowbrook, PA, USA (P.O. Box 3097, Meadowbrook 19046), revised edition, 1991. iv + 89 pp. LCCN QA29.V66 V664 1987.

vonNeumann:1997:AST

[von97] J. von Neumann. An axiomatization of set theory. Aut Aut, ?? (280–281):107–123, ???? 1997. ISSN 0005-0601.

vonNeumannWhitman:2005:LF

[von05] Marina von Neumann Whitman. Life with father. SIAM News, 38(2):??, March 2005. ISSN 0036-1437. URL http://www.siam.

org/news/news.php?id=38.

vonNeumann:2010:PET

[von10] J. von Neumann. Proof of the ergodic theorem and the H-theorem in quantum mechanics. European Physical Journal H, 35(2):201– 237, ???? 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic).

Wahlster:1996:EEC

[Wah96] Wolfgang Wahlster, editor. ECAI 96: 12th European Conference on Artificial Intelligence, August 11–16, 1996, Budapest, Hungary: proceedings. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-96809-9 (paperback). LCCN Q334 .E97 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/;   http://www.loc.gov/catdir/
description/wiley0310/99205889.html;   http://www.loc.
gov/catdir/toc/onix07/99205889.html.

Wallace:1996:FPG

[Wal96] C. S. Wallace. Fast pseudorandom generators for normal and exponential variates. ACM Transactions on Mathematical Software, 22(1):119–127, March 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1996-22-1/p119-wallace/.  See

comments [Bre08].

Waldinger:2016:BBS

[Wal16] Fabian Waldinger. Bombs, brains, and science: The role of human and physical capital for the creation of scientific knowledge. Review of Economics and Statistics, 98(5):811–831, December 2016. CODEN RECSA9. ISSN 0034-6535 (print), 1530-9142

(electronic). URL https://direct.mit.edu/rest/article/98/
5/811/58619/Bombs-Brains-and-Science-The-Role-of-Human-
and.

Ware:2008:RIE

[War08] Willis H. Ware. RAND and the information evolution: a history in essays and vignettes. Rand Corporation, Santa Monica, CA, 2008. ISBN 0-8330-4513-X, 0-8330-4816-3, 1-282-45123-5. xxvi + 201 pp. LCCN QA76.27. URL http://www.jstor.org/

stable/10.7249/cp537rc;   https://www.rand.org/content/
dam/rand/pubs/corporate_pubs/2008/RAND_CP537.pdf.

Walden:2014:ED

[WE14] David Walden and Nathan Ensmenger. From the Editor’s desk. IEEE Annals of the History of Computing, 36(3):2–3, July 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1988/pdf/a4348.pdf;   http://www.computer.org/
annals/an1988/a4348abs.htm.

Weintraub:1992:THG

[Wei92] E. Roy (Eliot Roy) Weintraub, editor. Towards a history of game theory: annual supplement to volume 24, History of political economy, volume 24 of History of political economy. Duke University Press, Durham, NC, USA, 1992. ISBN 0-8223-1253-0, 99-2403131-8. ISSN 0018-2702. vi + 306 pp. LCCN HB1 .H55 suppl. vol. 24 1992; HB144 .T68 1992.

Wigner:1996:SWJ

[WG96] E. P. Wigner and H. H. Goldstine. The scientific work of John von Neumann. In Wigner [Wig96c], pages 123–126. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Whitaker:2016:JSB

[Whi16] Andrew Whitaker. John Stewart Bell and twentieth century physics: vision and integrity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-874299-1. xvi + 460 pp. LCCN QC16.B442 W45 2016.

Wigner:1957:JN

[Wig57] Eugene P. Wigner. John von Neumann, 1903–1957. In ????, editor, Yearbook of the American Philosophical Society, pages 149– 153. American Philosophical Society, Philadelphia, PA, USA, 1957. ISSN 0065-9762. LCCN ????

Wigner:1967:JN

[Wig67a] Eugene Wigner. John von Neumann, 1903–1957. In Symmetries and reflections; scientific essays of Eugene P. Wigner [Wig67b], page ?? LCCN Q171 .W65.

Wigner:1967:SRS

[Wig67b] Eugene Paul Wigner. Symmetries and reflections; scientific essays of Eugene P. Wigner. Indiana University Press, Bloomington, IN, USA, 1967. viii + 280 pp. LCCN Q171 .W65.

Wigner:1967:NJH

[Wig67c] Wigner Jenő. Neumann János, 1903–1957. (Hungarian) [John von Neumann, 1903–1957]. Fizikai Szemle (Budapest), 17(8):227–229, ???? 1967. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Wigner:1980:BJN

[Wig80] Eugene P. Wigner. Biography of John von Neumann, Dec. 28, 1903–Feb. 8, 1957. In ????, editor, Dictionary of American Biography, Supplement Six: 1956–1960, pages 655–656. Scribner’s Sons, New York, NY, USA, 1980. ISBN ???? LCCN ????

Wigner:1996:BJN

[Wig96a] E. P. Wigner. Biography of John von Neumann. In Wigner [Wig96c], pages 209–210. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:JN

[Wig96b] E. P. Wigner. John von Neumann (1903–1957). In Wigner [Wig96c], pages 127–130. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96c] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wilkinson:1971:MEA

[Wil71] J. H. Wilkinson. Modern error analysis. SIAM Review, 13(4): 548–568, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). The 1970 von Neumann lecture.

Williams:1986:WNN

[Wil86] J. Williams. When is a von Neumann not a von Neumann? IEEE Micro, 6(1):5–6, January/February 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers: personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (e-book), 1-57955-011-8. 250 (est.) pp. LCCN

Q141 .W678562 2016. URL http://www.wolfram-media.com/
products/idea-makers.html.

Wigner:1992:REP

[WS92] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Plenum Press, New York, NY, USA; London, UK, 1992. ISBN 0-306-44326-0. xxiv + 335 pp. LCCN QC16.W52 A3 1992.

Wigner:2003:REP

[WS03] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Basic Books, New York, NY, USA, 2003. ISBN 0-7382-0886-8. xxiv + 335 pp. LCCN QC16.W52.

Wigner:1954:SLT

[WvN54] Eugene P. Wigner and John von Neumann. Significance of Loewner’s theorem in the quantum theory of collisions. Annals of Mathematics, 59:418–433, 1954. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). Reprinted in [Tau62, Paper 27].

Yee:1982:LWJ

[Yee82] Ooi Poh Yee. On the life and work of John von Neumann. Menemui Mat., 4(3):114–127, 1982. ISSN 0126-9003.

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

Zador:1983:VSN

[Zád83] Zádor Erika. 80 éve született Neumann János. (Hungarian) [John von Neumann was born 80 years ago]. Magyar hírlap, 16(405):8–??, ???? 1983. CODEN ???? ISSN ????

Zalai:2003:NJK

[Zal03] Zalai Ernő. Neumann János és a közgazdaságtan. (Hungarian) [John von Neumann and economics]. Magyar Tudomány [Hungarian Science], 48(12):1533–1538, 2003. CODEN ???? ISSN ????

Zund:2002:GDB

[Zun02] Joseph D. Zund. George David Birkhoff and John von Neumann: a question of priority and the ergodic theorems, 1931– 1932. Historia Mathematica, 29(2):138–156, May 2002. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086001923389.